A Complete Bibliography of Publications in the Journal of Grid Computing

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: https://www.math.utah.edu/~beebe/

08 August 2024 Version 1.40

Title word cross-reference

3 [387]. ^{3} [209]. K [554].

-Means [554].

15 [112]. 19 [762, 789].

2 [233]. 2.0 [411]. 2015 [417]. 2D [735].

3D [750].

512 [657]. 5G [731, 761].

Abdominal [552]. Abnormal [553]. Abstract [4]. Academic [207, 528]. Accelerate [719]. Accelerated [266].

Access [68, 109, 165, 170, 181, 223, 285, 293, 294, 416, 426, 480, 494, 574, 648, 666, 679]. Accessing [367]. Accountable [566]. Accounting [166, 741, 748]. Accuracy [119]. Accurate [559]. Achieving [119, 177, 196]. Across [113, 159, 173, 591, 748, 781]. Action [683]. Active [39, 490]. Activities [206]. Actor [550]. Actor-Critic [550]. Ad [236]. Adaptable [347]. Adaptation [7, 117, 538]. Adaptive [79, 100, 154, 161, 196, 244, 276, 297, 332, 453, 521, 547, 556, 664, 682, 775, 782]. Adding [239, 326]. Addressable [34]. Addressing [533]. Admission [318]. Adopters [86]. Adrenal [419]. Advance [213]. Advanced [26, 315]. Advancing [418]. Adversarial [552, 653, 694, 706]. Against [134]. Age [551]. Agent [103, 531, 752, 774]. Agents [67, 139, 248, 747]. Aggregation [164]. Agreement [111, 168, 759]. Agreements [578]. Agricultural [627]. Ahead [122]. AI [717, 728, 749, 769, 779]. AI-Based [769]. AI-Driven [717, 749]. Air [576, 690]. AirCargoChain [576]. Algebra [300]. Algorithm [127, 259, 266, 347, 356, 357, 366, 396, 432, 451, 467, 481, 485, 489, 526, 534, 542, 550, 554– 556, 561, 562, 601, 603, 615, 626, 665, 671, 676, 679, 682, 689, 704, 708, 726, 729, 770, 785]. Algorithms [6, 197, 346, 374– 377, 410, 454, 534, 553, 557, 775]. Alleviation [691]. Allgather [676]. Allocation [67, 150, 225, 236, 334, 338, 345, 409, 453, 455, 467, 522, 542, 604, 617, 625, 637, 639, 728, 771, 787]. ALPS [119]. Amazon [223, 251, 317, 371]. AMGA [136]. among [304]. Analyses [413]. Analysing [569]. Analysis [38, 45, 62, 137, 158, 169, 187, 198, 261, 275, 286, 309, 313, 317, 332, 368, 371, 403, 418, 431, 436, 442, 474, 501, 515, 519, 528, 537, 553, 558, 596, 635, 667, 700, 715, 733]. Analytical [740]. Analytics [491, 518, 571, 597]. Anatomy [360]. Android [570]. Annotation [511]. Annual [363]. Anomaly [555, 742, 782]. Answering [363]. Ant [485, 594]. Anti [714]. Anti-Attack [714]. Apache [719]. APPA [124]. Application [39, 90, 92, 119, 183, 190, 203, 221, 282, 325, 366, 373, 389, 446, 462, 515, 533, 568, 629, 636, 682, 703, 767, 791]. Application-Aware [39]. Application-level [119, 636]. Applications [12–14, 16, 25, 29, 35, 40, 45, 60, 66, 80, 86, 88, 89, 101, 117, 128, 141, 155, 161, 193, 197, 216, 234, 240, 243, 244, 247, 270, 344, 349, 353, 357, 394, 403, 424, 447, 459, 468, 487, 524, 530, 532, 536, 543, 572, 575, 582–584, 588, 589, 592, 612, 616, 621, 630, 637, 638, 677, 702, 717, 775]. Applied [612]. Applying [82]. Approach [42, 84, 100, 168, 220, 224, 257, 265, 315, 321, 336, 337, 369, 381, 391, 398, 402, 408, 414, 428, 434, 478, 490, 506, 507, 510, 558, 565, 576, 597, 598, 605, 632, 635, 650, 653, 656, 659, 663, 681, 720, 722, 727, 745, 752, 759]. Approaches [153, 235, 274, 541, 560, 577, 639]. Approximate [687]. Approximator [550]. Architectural [220]. Architecture [19, 50, 57, 85, 92, 103, 164, 194, 311, 324, 333, 355, 398, 406, 447, 467, 492, 542, 543, 549, 585, 630, 652, 719]. Architectures [48, 462]. Area [114, 141, 199, 400]. Arenas [501]. Array [386]. Arrhythmia [743]. Art [148, 667]. Artificial [587, 733]. Assessing [283, 295]. Assessment [241, 363, 390, 736]. Assisted [316, 703, 728]. Assurance [544]. Astro [200]. Astro-WISE [200]. Astrophysics [429]. Asynchronous [101]. ATLAS [132]. Atrial [214]. Attack [714]. Attacks [706]. Attention [707, 751]. Attribute [34, 165]. Attributes [3]. Auction [247, 610, 639]. Auction-based [247]. Audio [375]. Authentication [55, 693, 759]. Author [31]. Authorization [37, 53, 55, 104, 173, 650]. Authors [9]. Auto [27, 353, 445, 508, 727]. Auto-scaled [508]. Auto-Scaling [353, 445, 727]. Auto-Sizing [27]. Autoencoder [642]. Automated [97, 460, 511, 570, 596, 756]. Automatic [7, 252, 537, 603]. Automation [497]. Autonomic [51, 191, 273, 520, 583, 702]. Autonomous [449, 656]. Autoscaling [539, 753]. Availability [17, 115, 181, 184, 291, 294, 431, 464, 734]. Average [193]. Avian [215]. Aware [39, 108, 164, 273, 277, 319, 364, 370, 394, 397, 398, 402, 408, 410, 427, 432, 434, 446, 466, 475, 485, 508, 522, 574, 605, 606, 618, 628, 632, 648, 651, 659, 676, 681, 684, 691, 699, 704, 709, 734, 767, 785, 792]. Awareness [372]. Azure [387].

Back [207]. Back-End [207]. Badger [755]. Balance [526, 594]. Balancing [147, 235, 254, 294, 534, 546, 557, 608, 628, 630, 665]. BAN [618]. BAN-Storm [618]. Bandwidth [278, 415, 618, 787]. Bandwidth-Aware [618].

Bandwidth-Limited [278]. Based [35, 37, 52, 63, 74, 84, 96, 101, 109, 155, 165, 170, 175, 184, 192, 195, 213, 214, 222, 232, 245, 263, 291, 299, 314, 334–336, 350, 375– 377, 379, 385, 391, 432, 435, 448, 449, 470, 478, 479, 481, 485, 486, 489, 490, 492, 519, 520, 529, 531, 537, 546, 550, 553, 555– 557, 559, 574, 575, 590, 592, 595, 597, 598, 600, 605, 609, 611, 633, 634, 641, 651, 665, 671, 675, 685, 687, 689, 707, 710, 712– 714, 718, 720, 724, 725, 731, 738, 740, 744, 745, 756, 760, 764, 769, 770, 772, 774, 780, 783, 789]. based [5, 87, 100, 103, 111, 150, 169, 247, 249, 276, 282, 285, 289, 320, 348, 381, 395, 442, 468, 493, 513, 522, 577, 584, 635, 638, 639, 646, 658, 659, 661, 662, 672, 680, 690, 711, 723, 754, 757, 762, 775, 782, 784, 786]. Bases [283]. BaSTI [201]. Batch [118, 244]. Batches [185]. Battle [501]. Bayesian [259, 685]. Be [593]. Behavior [222, 553]. Behavioural [660]. Belief [609]. Belle [223]. Benchmarking [460, 579]. Benefits [118]. Better [181]. between [421, 674]. Beyond [162, 217, 419, 511]. Bi [393]. Bi-Objective [393]. Bibliometric [733]. Bid [610]. Big [411, 424, 425, 469, 472, 479, 488, 490– 493, 509, 512, 513, 517–519, 530, 540, 554, 556, 557, 559, 587, 632, 680, 691, 721]. BiLSTM [760]. BiLSTM-Based [760]. Bin [785]. Bin-packing [785]. Binding [65, 444]. Bioinformatics [183, 190, 288, 309]. Bioinspired [374]. Biology [287]. Biomechanics [640]. Biomedical [286, 568]. Bitcoin [499, 573]. Black [755]. Blade [290]. Block [477]. Blockchain [572, 574–576, 590, 641, 703, 711, 717, 718, 723, 728, 733, 736, 749, 752, 769, 779, 781]. Blockchain-Based [574, 641, 711, 723]. Blockchain-Enabled [575, 717]. Blockchains [701]. BlockFaaS [717]. Blocks [429]. BOINC [176, 187, 413, 545]. Border [758]. Bottom [597]. Bottom-Up [597]. Boundary [551]. BPEL [78]. Brain [634]. Bridging [176, 421]. Broadcast [272].

Broader [279]. Broker

[170, 289, 330, 506, 698]. Brokering [95, 111, 430]. Budget [318, 357]. Budget-Deadline [318]. Buffer [27]. Building [46, 139, 429, 569, 726]. Bulk [36]. Bunching [673]. Bundle [311]. Bus [328, 673]. Business [78, 144, 336, 590]. Butterfly [601, 615].

C [506]. C-RCE [506]. CA [394]. CA-DAG [394]. Cache [412, 513, 574]. Caching [181, 444, 472, 700, 757]. Cameras [502]. Campus [89]. Can [150]. Cancer [419]. Canopy [556]. Canopy-K-Means [556]. Capability [164, 347, 492]. Capability-Aware [164]. Capacitated [674]. Capacity [255]. Capturing [660]. Carbohydrate [218]. Cargo [576]. Carlo [202]. CAS [109]. Case [194, 270, 288, 308, 364, 414, 495, 635]. Catalog [238]. Causal [666]. Causes [673]. Center [407, 668, 715]. Centers [370, 397, 406, 547, 646, 671, 778]. Centres [466]. Centric [85, 513]. Chain [752, 758]. Chaining [709]. Challenge [363]. Challenges [59, 122, 152, 303, 404, 411, 447, 581, 584, 596]. Channels [452]. Chaos [675]. Characteristics [714]. Characterizing [3]. Charité [285]. checking [189]. Checkpoint [298]. Checkpoint/Restart [298]. Checkpointing [354]. Chemistry [91, 358, 433]. Cherenkov [386]. China [47]. Chinese [528]. Chirp [160]. Chromodynamics [69]. CIA [779]. Circuits [30]. CIT [564]. Cities [712, 736, 737, 768]. City [258, 725, 726, 761, 765]. CJSpector [655]. Classification [493, 512, 557, 584, 738, 743]. Clients [222]. Climate [341]. Cloud [178, 220– 222, 247, 252, 255, 258, 260, 269, 273, 289, 290, 292, 321, 329, 335, 336, 353, 361, 367, 374, 378, 389, 391, 394, 399, 400, 405, 406, 408, 414, 424, 426, 438, 445, 449, 450, 457, 461, 463, 465, 467, 470, 478, 494, 496, 503, 505, 506, 521, 526, 529– 532, 534, 536, 543, 546, 560, 564, 565, 578, 581– 583, 591, 597, 607, 608, 610, 614, 616, 621, 628, 631, 637, 643, 645, 656, 657, 661, 663, 664, 670, 671, 689, 691, 693, 702, 703, 709, 715, 734, 742, 747, 750, 760, 777]. Cloud [223, 301, 303, 330, 338, 359, 379, 383, 387, 396, 398, 404, 407, 412, 435, 442, 443, 446, 459, 464, 482, 525, 537, 538, 542, 547, 556, 561, 564, 577, 592, 594, 604, 606, 635, 636, 649, 650, 668, 693, 722, 727, 729, 739, 745, 748, 757, 769, 772, 778, 782]. Cloud-Based [379, 478, 529, 531]. Cloud-Edge [616, 757]. Cloud-Edge-End [645]. Cloud-Manager-Based [391]. Cloud-Native [583, 702]. Cloud-Network [631]. Cloud-Oriented [465]. Cloud-to-Edge [636]. Clouds [191, 250, 257, 276, 297, 309, 324, 325, 331, 342, 346, 393, 395, 403, 410, 431, 439, 460, 462, 484, 486, 507, 520, 522, 539, 589, 599, 620, 705]. Cluster [160, 240, 317, 405, 595, 792]. Clustered [455, 520]. Clustering [554, 556, 658]. Clusters [51, 508, 591, 658, 753]. CMS [132, 198, 472]. CMSV [747]. CNN [735]. Coalition [622]. Coatings [612]. Code [66, 648, 697]. Code-Oriented [648]. Codes [278, 452]. Coding [678]. CoG [76]. Cognitive [259]. Collaboration [757]. Collaborations [162]. Collaborative [496, 590, 645, 694, 774]. Collecting [611]. Collection [613]. Colluding [188]. Colony [485, 594]. Combination [674]. Combined [759]. Combining [224, 523]. Commerce [758]. Commercial [130, 152, 158]. Common [308]. Communication [147, 355, 394, 632, 672, 792]. Communication-Aware [394]. Communities [280]. Community [52, 490, 500, 507]. Comparative [414, 442]. Comparison [110]. Compatible [610]. CompatibleOne [330]. COMPCHEM [217]. Completion [193, 653]. Complex [4, 369, 380, 462, 463, 599]. Complexity [269]. Compliant [161]. Components [420]. Composition [43, 314, 361]. Compositional [70]. Comprehensive [447, 541, 560, 736]. Compression [297, 471, 473]. COMPSs [450]. COMPSs-Mobile [450]. Computation [6, 22, 80, 150, 189, 253, 577, 648, 672, 684, 708, 724]. Computational [7, 83, 90, 91, 101, 126, 196, 210, 236, 322, 346, 359, 754]. Compute [223, 317, 496]. Computer [42, 316, 710]. Computer-Assisted [316]. Computing [2, 21, 28, 56, 77, 93, 140, 148, 156, 165, 178, 180, 185, 206, 220, 223, 237, 250, 252, 272, 279, 290, 300, 310, 319, 356, 358, 363, 365, 369, 373, 384, 389, 394, 406, 408, 409, 438, 449, 450, 454, 457, 460, 461, 494, 495, 505, 532, 541, 545, 546, 557, 560, 574, 578, 579, 581, 600, 601, 608, 612, 615, 617, 619, 630, 641, 644, 645, 648, 649, 659, 662, 663, 665, 674, 679, 681, 683, 689, 692, 697, 707, 708, 712, 715, 717, 720, 728, 730, 732, 735, 736, 739, 742, 743, 763, 773, 781, 783, 791]. Computing [5, 16, 18, 42, 47, 65, 73, 142, 147, 158, 160, 186, 188, 189, 195, 253, 259, 303, 359, 404, 432, 438, 448, 458, 538, 549, 577, 580, 585, 598, 602, 604, 605, 623, 626, 636, 651, 672, 690, 694, 699, 716, 725, 727, 740, 741, 744, 749, 754, 761, 764, 766, 768, 787]. Computing-Based [598, 720]. Concepts [76]. Concurrent [193]. Conditional [448]. Conference [417]. Configurable [718, 784, 786]. Configuration [56, 252]. Configuring [760]. Conflicts [650]. Connectivity [159]. Considering [294, 604, 654]. Consistency [437, 619]. Consolidated [436]. Consolidation [485, 646]. Constrained [318, 357, 370]. Constraint [307]. constraints [484]. Construct [565]. Constructing [12, 324, 506]. Construction [455]. Consuming [342]. Consumption [320, 481, 486, 624, 654]. Container [467, 492, 646, 753]. Container-Based [492, 646]. Containerised [592, 621]. Containerized [704]. Containers [468, 539, 620, 753, 776]. Content [100, 503]. Content-based [100]. Contents [32]. Context [105, 449, 605, 607]. Context-Aware [605]. Context-Based [449]. Contextual [709]. Continuum [621, 631, 636, 750, 791]. Contract [672]. Contract-based [672]. Contracts [578, 590]. Contrast [584]. Contribution [85]. Control [7, 109, 165, 170, 293, 318, 331, 402, 416, 574, 586]. Controlled [29, 789]. Controlling [675]. Convergence [733]. Converging [544]. Convolutional [652]. Cooperation [400, 531, 744]. Cooperation-Based [744]. Cooperative [444, 672, 708]. Coordination [153]. Correction [615, 786]. Correlation [633]. Cosmological [264]. Cost [156, 221, 251, 277, 363, 383, 455, 460, 486, 676, 691, 698, 734, 753, 776]. Cost-Availability [734]. Cost-Aware [277, 691]. Cost-Efficient [455, 698, 776]. Cost-Performance [460]. Cost-Time [383]. Coupled [240, 244]. Covid [762, 789]. Covid-19 [762, 789]. CP [299]. CP-Nets [299]. CPU [194, 434, 669, 719]. CPU-Intensive [434]. Crawling [483]. Create [272]. Criteria [266]. Criterion [658]. Critic [550]. Critical [589]. Cross [551, 758]. Cross-Age [551]. Cross-Border [758]. Crowd [659]. Crowdsensing [390, 613]. Cryptography [677]. Cryptojacking [655]. Cryptomining [558]. CT [552]. Cuckoo [512]. Cumulative [643]. Customized [735]. Cyber [685, 713]. Cyberinfrastructure [91, 231]. Cycle [96, 115, 157]. Cycle-Sharing [96].

D [387, 696]. DAG [394]. DAGMan [116]. DAI [109]. DARE [282]. DAScheduler [704]. Dashboard [206]. Data [6, 11, 17, 27, 38, 40, 48, 64, 80, 98, 113, 120, 121, 123, 124, 135, 138, 155, 162, 163, 170, 177, 181, 207, 250, 258, 264, 286, 303, 309, 333– 335, 337, 346, 348, 349, 351, 370, 382, 406, 407, 411, 418, 424, 444, 451, 465, 466, 470, 472, 478, 479, 488, 490–493, 496, 509, 512, 513, 517– 519, 524, 526– 528, 530, 554, 557, 567, 576, 587, 590, 597, 601, 613, 615, 619, 632, 633, 646, 657, 664, 667, 671, 673, 676, 683, 686, 691, 715, 721, 723, 739, 759]. Data [36, 108, 197, 259, 271, 294, 338, 397, 425, 443, 469, 473, 476, 481, 547, 555, 556, 624, 668, 680, 699, 700, 716, 752, 778]. Data-Aware [632]. Data-Driven [664]. Data-Flow [271]. Data-Grids [48]. Data-Intensive [80, 197, 250, 303, 338, 382]. Data-Oriented [492]. Databases [109, 201, 540]. Datacenters [508]. DataCloud [480, 544]. DataGrid [58–61]. Dataset [249, 472]. Datasets [559, 568]. Datastore [207, 292]. DCDedupe [471]. DCI [232, 239, 280]. Deadline [318, 606, 767]. Deadline-aware [767]. Deadlines [787]. Deal [269]. Decentralized [52, 101, 238, 500, 563, 575, 620, 701, 754, 769]. DECIDE [284]. Decision [300, 381, 519, 603, 683, 732]. Decision-Making [603]. Decisions [297]. Declarative [629]. Dedicated [183, 190]. Deduplication [471]. Deep [511, 548, 549, 551, 553, 558, 602, 609, 617, 632, 642, 645, 653, 655, 663, 690, 695, 703, 721, 726, 731, 732, 737– 739, 743, 752, 757, 760, 762, 763, 766, 770, 772– 774, 789]. DEEPBIN [738]. Defeating [188]. Defense [452]. Defined [455, 466, 513, 603, 711]. Definition [87, 173, 557]. Definitions [3]. DEISA [234]. Delay [147, 761]. Delegation [239, 580]. Delta [471]. Demand [230, 272, 607, 698, 775]. Demand-Driven [230]. Demands [599, 771]. Dengue [215]. Dense [783]. Departmentalized [149]. Dependable [291]. Dependency [704]. Dependency-Aware [704]. Dependent [600, 679, 704, 763]. Dependent-Tasks [600]. Deploy [463, 513]. Deploying [262]. Deployment [87, 172, 238, 252, 343, 348, 422, 464, 526, 588, 589, 591, 778]. Depth [696].

Describing [582]. Design [8, 49, 90, 97, 111, 142, 245, 316, 366, 442, 443, 553, 568, 682, 709, 726, 778]. Desktop [128, 180–182, 188, 195, 211, 216, 232, 296]. Detected [516]. Detecting [558]. Detection [490, 514, 553, 559, 593, 632, 655, 687, 696, 713, 714, 742, 743, 750, 751, 760, 782, 788, 789]. Detector [585]. Determining [435, 442]. Developing [91, 612, 660]. Development [129, 228, 281, 389, 422, 651, 671, 740]. Device [525]. Devices [320, 531, 604, 694, 695, 723, 741, 748, 777, 792]. DGSI [239]. DHT [100]. DHT-based [100]. Diagnosis [766]. Diagonal [657]. DIANA [108]. Differences [421]. Differencing [483]. Differential [724]. Digital [257, 488, 741]. dimensional [442, 662, 785]. Dimensioning [414]. DiPerF [97]. DIRAC [223]. Directions [148, 447, 596]. Disaster [529, 652]. DisCoP [365]. Discovering [327]. Discovery [23, 52, 125, 126, 154, 167, 208, 215, 496, 500, 503, 509, 588, 644]. Discrete [785]. Discriminative [551]. Discussion [145, 261]. Disease [419]. Disk [117, 492]. Disk-as-a-Resource [492]. Dispatching [787]. Dissemination [162, 517]. Distance [26]. Distributed [11, 18, 22, 36, 45, 97, 127, 181, 198, 234, 246, 247, 253, 270, 279, 282, 307, 310, 320, 333, 351, 354, 355, 366, 369, 409, 412, 422, 444, 456, 471, 477, 502, 576, 589, 590, 595, 596, 611, 620, 644, 685, 697, 712, 718, 770, 787]. Distribution [98, 290, 665, 769]. Distributions [255]. Divisible [197]. DNA [515]. DNN [595]. Docker [405]. Docking [569]. Does [156]. Domain [55, 333, 414, 731]. Dominating [573]. Double [247, 657, 773]. Downloads [36]. DPark [701]. Drawbacks [118]. DRIHM [423]. Driven [43, 230, 231, 322, 343, 664, 678, 684, 686, 717, 749]. Driving [653, 747]. DRL [754]. DRL-based [754]. Drug [215]. Dueling [773]. Duplication [600].

Duplication-Based [600]. During [652]. DVFS [396]. DVFS-enabled [396]. DynaGrid [161]. Dynamic [43, 62, 63, 113, 147, 163, 166, 202, 344, 362, 384, 414, 446, 493, 500, 510, 526, 558, 628, 665, 670, 671, 765, 771]. Dynamically [50]. Dynamics [44].

E-Commerce [758]. E-Healthcare [512]. e-Infrastructure [228, 233, 497]. E-Infrastructures [229, 480, 544]. e-Research [171]. e-Science [106, 229, 566]. e-Service [249]. Early [86]. Earth [11]. Earthquake [123, 649]. Eastern [228]. Easy [281, 300]. EC2 [251, 317, 371]. Economic [152, 224, 340]. Economics [144, 145, 149, 150]. Economics-based [150]. Ecosystem [339, 656]. Ecosystems [627]. Edge [533, 543, 574, 577, 578, 589, 601, 602, 604, 605, 613, 615–617, 620, 623, 624, 626, 636, 639, 645, 648, 672, 674, 679, 683, 690, 693– 696, 699, 707, 708, 710, 712, 720, 723, 725, 730, 732, 733, 735–737, 739–741, 743–745, 748– 751, 756, 757, 761, 763– 766, 768, 773, 777, 781, 783, 787, 790, 792]. Edge-Based [745]. Edge-Cloud [543, 693]. Edge-enabled [613]. Edge-Inspired [696]. Edge-to-Cloud [578]. Edge/Cloud [577]. EDGeS [176]. Editor [305, 461]. Editorial [10, 24, 33, 93, 392, 498]. Editors [1, 219, 226, 250, 325, 335, 469]. Education [777]. Educational [777]. EEG [516]. EELA [233]. EELA-2 [233]. Effective [150, 186, 471, 595, 613, 699, 790]. Effectively [233]. Effects [487]. Efficiency [119, 334, 395]. Efficient [13, 46, 208, 236, 237, 273, 354– 356, 359, 383, 396, 407, 455, 457, 476, 484, 517, 540, 561, 597, 598, 631, 653, 658, 663, 665, 666, 668, 683, 698, 705, 708, 729, 772, 776]. EGEE [155, 173, 176, 200, 202–205, 207, 215, 227]. EGI [497]. Elastic [221, 223, 353, 459, 460, 487, 508, 591, 638, 771]. Elasticity [403, 638]. ElasticSim [445]. Electrocardiogram [743]. Elements [249].

Elite [708]. Embedding [742]. EmBOINC [187]. Emerging [639]. EMI [327]. Emotion [516, 609]. EMPEROR [63]. Empirical [115, 187, 332, 673]. Employing [768]. Empowered [217, 766]. Enabled [134, 140, 194, 199, 286, 351, 396, 575, 613, 645, 717, 751]. Enabler [692]. Enabling [19, 68, 179, 249, 304, 426, 459, 784, 786]. Encoding [624]. Encryption [391, 657, 675, 703]. End [77, 207, 241, 271, 504, 645, 731]. End-to-End [77, 271, 504, 731]. End-user [241]. Energy [132, 236, 273, 319, 359, 370, 385, 395, 396, 398, 401, 407, 408, 434, 454, 476, 481, 484–486, 489, 508, 561, 594, 606, 654, 667, 668, 684, 700, 792]. Energy-Aware [319, 370, 408, 434, 485, 508, 684]. Energy-Based [385]. Energy-Efficient [236, 273, 396, 476, 484, 668]. Engineering [49, 380, 544, 590]. Engines [667]. English [518]. Enhanced [764, 792]. Enhancement [99, 702]. Enhancing [7, 126, 321, 380, 505, 516, 564, 627, 783]. Ensemble [199, 379, 658]. Ensembles [410, 622]. Ensuring [393]. Enterprise [122, 128, 220]. Enterprises [149]. ENTICE [474]. Entropic [102]. Entropy [767, 789]. Environment [15, 19, 64, 137, 282, 290, 344, 391, 396, 419, 448, 534, 577, 588, 590, 598, 604, 606, 628, 637, 657, 665, 691, 716]. Environmental [726]. Environments [4, 16, 21, 55, 146, 185, 237, 247, 297, 353, 354, 529, 536, 601, 610, 614, 615, 622, 644, 650, 662, 680, 681, 727, 729, 784, 786, 790]. EPMA [638]. Era [425, 469]. Erratum [190]. Essential [420]. Estimates [225, 267]. Estimating [552]. Estimation [665]. Estimation-Based [665]. EU [341]. EUAsiaGrid [215]. Eucalyptus [431]. Europe [228]. European [58, 61, 232, 234, 352, 363, 544]. Evaluating [482]. Evaluation [8, 26, 97, 115, 116, 142, 162, 166, 216, 238, 251, 256, 433, 435–437, 442, 529, 536, 696, 791].

Event [514, 611, 678]. Event-Driven [678]. Events [519, 613]. Evolution [544, 724]. Evolutionary [375–377, 380, 598, 656]. EvoSpace [376]. ExaFlooding [644]. Examination [756]. Examinations [510]. Example [201]. Exascale [644]. Executing [203, 383]. Execution [13, 74, 78, 107, 183, 190, 196, 332, 338, 388, 403, 445, 520, 705]. Executions [244, 311]. Expansion [433]. Expectations [303]. Experience [60, 132, 203, 302]. Experiences [172]. Experiment [206, 301]. Experimenting [39]. Experiments [106, 179, 456, 627]. Exploration [264]. Exploring [313, 781]. Expression [609]. Extending [258, 424]. Extraction [512]. Extremely [272, 458]. Eye [559]. Eyewitness [652].

Fabrics [56, 252]. Face [551, 585].

Facilitate [480]. Facility [127]. Failure [261, 711]. Fair [96, 771]. Fairness [694]. Fast [26, 70, 298, 706]. Fault [20, 88, 125, 307, 332, 377, 384, 543, 562, 593, 687]. Fault-Detection [593]. Fault-Tolerance [307, 377]. Fault-Tolerant [88, 125, 332, 384, 543, 562]. Faults [205]. Feature [380, 512, 568]. Features [209, 501, 673, 702]. Featuring [112]. Federated [207, 293, 321, 324, 326, 327, 439, 470, 496, 497, 543, 605, 694, 705, 737, 751, 757]. Federation [239, 325, 661]. Fever [215]. Few [688]. Few-Shot [688]. Fields [448]. File [30, 99, 238, 293, 320, 451, 477, 483]. filesystem [160]. Filter [559]. Filter-Based [559]. Financial [441, 730]. Fine [115, 310, 416, 566]. Fine-Grain [310]. Fine-Grained [115, 416, 566]. Firefly [408]. Firewalls [159]. Fitness [736]. Fixed [413]. Fixed-Length [413]. Flash [241]. FlexGP [379]. Flexible [20, 165, 182, 258, 260, 568]. Flood [241]. Flow [7, 271, 557]. Flow-Control [7]. Flowbster [465]. Flu [215]. Flutter [214]. FOCALB [630]. Fog [525, 541, 542, 563, 610, 619, 622, 629, 630, 637, 649, 651, 656, 662, 718, 728, 739, 767]. Fog-Based [718]. Fog-Cloud [525, 610, 637, 739]. Forecast [222, 415]. Forecasting [84, 199, 241, 547]. Formal [245, 263, 428, 650]. Formalisms [3]. Formation [622]. Forum [2]. Four [110]. Fourier [675]. FPGA [719]. Fractional [675]. Fragment [784, 786]. Fragmentation [474]. Framework [20, 81, 97, 105, 107, 146, 152, 161, 168, 208, 221, 253, 263, 280, 282, 329, 332, 359, 365, 430, 435, 443, 483, 493, 514, 519, 520, 527, 554, 555, 569, 570, 583, 637, 649, 661, 669, 680, 689, 694, 705, 711, 717, 721, 723, 728, 731, 767, 775]. Frameworks [325, 424, 567]. freeCycles [457]. Frequent [333, 540]. Friendly [285, 568]. Frontend [780]. Fulfilling [277]. Function [458, 681]. Fundamental [520]. Future [148, 447]. Futures [224]. Fuzzy [710, 746].

g [5, 695]. Gabor [512]. Galaxy [420]. Game [275, 491, 713]. Game-Theoretic [275]. Garbage [738]. GATE [202]. Gateway [280, 283, 284, 286, 386, 417, 418, 423, 424, 433, 635]. Gateways [279, 281, 282, 288, 421, 422, 426, 427, 429, 430, 569]. Gaussian [550]. GEMLCA [66]. GEMS [217]. GeneGrid [92]. General [298]. Generalized [675]. Generation [11, 178, 568, 692, 716]. Generative [552, 653]. Generic [50, 168, 280, 569]. Genetic [266, 356, 379, 467, 475, 542, 626, 770]. Genome [214]. Genomics [414]. Geo [425, 666]. Geo-Linkage [425]. Geo-Replicated [666]. Gesture [735]. gLite [135, 249]. gLite-based [249]. Global [2, 93, 160, 229, 523]. GMATE [344]. go [315]. Goal [760]. Goals [385]. Gorilla [679]. GPS [673]. GPU [157, 595, 669]. Grade [15, 75, 300]. Gradient [706]. Grain [310]. Grained [115, 416, 566]. Granularity [407]. Graph [478, 707, 742]. Graph-Based [478]. Greedy [489]. Greek [194, 262]. Green [295, 399, 400, 581, 726]. Greener [378]. Greening [600]. Grey [785]. Grid [2, 4, 13–15, 20, 28, 29, 34, 37, 45, 55– 57, 63, 64, 66–68, 70, 75, 78, 80, 82, 85– 87, 90, 91, 101–103, 129, 132, 134, 137, 145, 146, 149, 165, 169, 172–175, 177, 179, 182, 183, 190, 192, 194, 199, 200, 203, 205, 209, 214, 216, 217, 228, 242, 245, 248, 249, 263, 264, 266, 267, 275, 299, 302, 322, 332, 334, 340, 341, 344, 349, 351, 352, 354, 356, 358, 366, 367, 372– 374, 385, 437, 438, 505, 612, 654, 665, 682]. Grid [5, 12, 14, 16, 25, 35, 36, 40, 42–44, 47, 49– 53, 62, 65, 69, 72–74, 77, 108, 109, 111, 120, 121, 130, 135, 138, 140, 143, 144, 147, 153, 155, 156, 159, 160, 167, 170, 171, 178, 184, 188, 191, 193, 195, 201, 208, 213, 224, 231, 235, 241, 243, 247, 261, 269, 274, 285– 287, 291, 304, 319, 345, 348, 360, 368, 562, 580]. Grid-Based [74, 101, 285]. Grid-Controlled [29]. Grid-Enabled [134, 140, 194, 199, 286]. GridCertLib [243]. GridRPC [88]. Grids [3, 6, 7, 17, 19, 38, 39, 48, 79, 83, 89, 95, 104, 107, 122, 125, 126, 128, 133, 139, 141, 148, 152, 157, 164, 180, 196, 210– 212, 225, 230, 232, 233, 236, 240, 244, 257, 262, 270, 294–296, 346, 347, 416, 434, 451]. GridSAT [90]. Group [348, 451]. Grouping [208]. Growth [689]. GRUBER [111]. GS [209]. Guaranteed [196]. Guest [226, 250, 305, 325, 335, 461, 469]. Guidelines [350]. gUSE [280]. GVSS [215].

H [719]. H-Storm [719]. Hadoop

[350, 475, 554]. Handling [337]. Happening [514]. Hard [787]. Hardware [655, 677]. Hash [246]. HDCache [412]. Health [478, 604, 721, 725]. Healthcare [512, 703, 717, 720, 736, 737, 762, 766, 790]. Heart [739]. Heavy [255]. HECTOR [179]. Heliophysics [312]. Hellenic [179]. Help [296]. HEP [60]. Heterogeneity [350, 482]. Heterogeneous [50, 223, 290, 319, 384, 432, 454, 462, 520, 521, 527, 534, 549, 557, 585, 598– 600, 665, 674, 691, 729]. Heuristic [276, 546, 608, 659, 695]. Heuristics [141, 484]. Hierarchical [225, 278, 401, 658, 685, 718]. High [18, 21, 24, 25, 27, 29, 30, 113, 132, 215, 271, 272, 317, 346, 415, 464, 602, 667, 700]. High-Bandwidth [415]. High-Energy [132]. High-Performance [21, 27, 346]. High-Speed [30, 271]. High-throughput [272]. Highlights [18]. Highly [599]. Highway [548]. HIWDEO [724]. Hoc [236]. Holistic [443]. Homogeneous [223]. Homology [511]. Honey [755]. Honeypot [713]. Host [110, 663]. Households [738]. HPC [273, 302, 422, 537]. HPDC [112]. HPDC-15 [112]. HRNN [788]. HTTP [320]. HTTP-Based [320]. Human [640]. Hybrid [224, 324, 391, 398, 400, 439, 521, 565, 586, 591, 598, 608, 675, 705, 719, 722, 724, 726, 746, 748, 775, 785]. Hybridized [703, 755]. Hydro [423]. Hydro-meteorological [423]. Hypergraph [788].

I/O [317, 436, 477]. IaaS [325, 393, 395, 486, 561, 660]. IaaSMon [406]. iCanCloud [260]. ICSMPC [709]. Identification [642, 650, 652]. Identifying [340]. Identity [326, 496]. iHOME [479]. II [322]. IKAROS [320]. Image [474, 675, 710]. Images [61, 552]. iMapReduce [253]. Imbalance [546, 785]. Imbalance-ness [785]. Imbalanced [493]. Immune [724]. Impact [482, 573, 624]. Impacts [571]. Impersonation [510]. Implementation [5, 8, 88, 90, 92, 142, 164, 169, 173, 439, 553]. Improve [395, 737]. Improved [372, 435, 513, 554, 555, 601, 615, 626, 653, 679]. Improvement [274, 780]. Improving [181, 186, 388, 671, 709, 777]. Incentive [491, 610]. Incentive-Compatible [610]. Incomplete [653]. Incremental [526]. Index [31, 479]. Index-Based [479]. Indexing [563]. INDIA [341]. INDIGO [480, 544]. INDIGO-DataCloud [480].

Industrial [373, 575, 586, 612, 713, 769]. Industry [790]. Inference [774]. Information [34, 52, 57, 70, 139, 164, 175, 437, 513, 551, 653, 710, 741]. Informatization [748]. Informed [297]. Infrastructure [101, 111, 155, 179, 203, 204, 216, 227, 228, 230, 233, 234, 239, 249, 260, 264, 273, 288, 308, 315, 331, 343, 352, 378, 400, 418, 427, 456, 497, 522, 612]. Infrastructures [229, 272, 279, 289, 310, 341, 362, 409, 421, 422, 463, 480, 496, 508, 544, 587, 616]. Initio [387]. Innovation [230]. Innovative [386, 641, 658, 740]. Input [567]. Inspired [696]. Installation [56]. Instances [371, 698]. Institutions [777]. Instructions [9]. Instrument [44]. Instrumentation [62]. Integrated [158, 175, 321, 419, 637, 649, 663]. Integrates [207]. Integrating [140, 398, 565, 777]. Integration [51, 65, 89, 177, 178, 201, 281, 335, 339, 693, 735, 749, 779]. Intelligence [235, 733, 755, 765]. Intelligent [527, 530, 553, 669, 711]. Intensive [80, 108, 197, 250, 259, 303, 338, 382, 434, 466, 524]. Inter [482, 564]. Inter-cloud [482]. Inter-operation [564]. Interaction [349]. Interactive [19, 337]. Intercloud [381]. Intercontinental [230]. Interfaced [135]. Interfaces [325]. Interfering [673]. Internal [660]. Internet [575, 604, 711, 718, 720, 740, 746, 751, 752, 754, 764, 765, 769, 770, 773, 779]. Internet-of-Things [711]. Interoperability [171, 173, 239, 245, 304, 310, 313, 325, 328, 428]. Interoperable [175, 329]. Interoperate [233]. Interoperation [172, 174, 177]. Interplay [741]. Interprocess [355]. Introduce [129]. Introduction [18, 250, 305, 325, 335, 461, 469]. Intrusion [632, 714, 751, 760, 788]. Invalidating [255]. Invasive [689]. Investor [730]. IoMT [737]. IOT [525, 530, 531, 542, 543, 597, 619, 624, 631, 637, 649, 651, 656, 693, 703, 712, 717, 721, 724, 726, 736, 762, 781]. IoT-Based [712, 762]. IoT-Cloud [631]. IoT-Fog-Cloud [542]. IoTEF [543]. IoVs [602]. IQ [113]. IQ-Paths [113]. Irradiance [686]. Island [375, 377]. Island-Based [377]. Issue [18, 24, 93, 112, 120, 180, 250, 305, 325, 335, 417, 438, 461, 469, 488, 509, 530, 572]. Issues [13, 121, 404, 581]. Items [333]. Itemsets [333, 540]. Iterative [101, 253, 709]. Iterative-Learning [709].

Java [76]. JETS [306]. Job [140, 192, 225, 266, 275, 331, 334, 409, 453]. Jobs [116, 137, 205, 434, 489, 618, 704]. JOIN [479]. Joint [221, 712, 746, 753, 764, 770, 778, 787]. Jones [585]. Journal [438]. Journey [499].

KDBD [509]. Keeping [407]. Kernel [298, 687]. Kestrel [301]. Key [256, 759]. Keyword [22]. Kit [76]. Knowledge [283, 509, 660, 742, 784, 786]. Knowledge-based [784, 786]. Kosha [99]. Kubernetes [623].

Labs [94]. Lake [700]. Landscapes [781]. Language [78, 169, 252, 306, 311, 518, 697]. Large [17, 51, 56, 77, 123, 126, 127, 162, 183, 190, 272, 280, 300, 307, 349, 379, 422, 458, 465, 473, 515, 600, 611, 633, 667, 697, 718]. Large-Scale [56, 77, 127, 307, 515, 600, 667, 697, 718]. Late [444]. Late-binding [444]. Latency [185, 533, 589, 623, 761]. Latency-Critical [589]. Launch [141]. Launch-time [141]. Layer [207]. Layout [476]. LCG [238]. Learned [91, 422, 423]. Learning [67, 212, 348, 379, 490, 511, 538, 548–551, 558, 571, 602, 605, 607, 617, 632, 640, 645, 647, 653, 655, 663, 677, 686, 688, 690, 695, 703, 709, 710, 721, 722, 726, 731, 732, 738, 739, 742, 743, 748, 752, 757, 762, 766, 770, 777, 780, 782, 789, 790]. Learning-Based [605, 731, 757, 770]. Least [687]. Ledgers [504]. Legacy [66]. Length [413]. Lesson [423]. Lessons [91, 422]. Level [21, 111, 119, 177, 578, 636, 723]. Leveled [356]. Leveraging [470, 635, 667]. LHC [206]. Libraries [528]. Library [237]. License [646]. Lidar [750]. life [270]. Lifecycles [85]. Lightpaths [29]. Lightweight [507, 735]. Limited [278, 479]. LincoSim [537]. Linear [300]. Lines [745]. Link [711]. Linkage [425]. Linux [298]. Live [459]. LiveWN [262]. Load [147, 197, 235, 290, 294, 523, 526, 534, 546, 557, 608, 628, 630, 663, 665]. Loadable [298]. Local [117, 127]. Locality [676]. Locality-Aware [676]. Localization [755]. Location [82, 127]. Logarithmic [676]. Logarithmic-Cost [676]. Logistics [758]. Long [26, 413]. Long-Distance [26]. Long-Running [413]. Low [185, 320, 623, 697, 757]. Low-Code [697]. Low-Latency [623]. Low-power [757]. LSTM [772].

MAAN [34]. Machine [84, 254, 297, 395, 397, 408, 440, 452, 485, 486, 538, 581, 607, 640, 643, 647, 677, 686, 710, 722, 742, 748, 777, 780, 782, 790]. Machines [370, 398, 539, 594, 671, 753]. MAGDA [103]. Main [534]. Mainstream [86]. Maintain [705]. Major [340]. Makespan [385, 606, 729]. Making [297, 300, 519, 603]. Malaria [134]. Malware [558]. Manage [459, 463]. Management [51, 56, 58, 59, 64, 73, 100, 120, 121, 124, 192, 207, 224, 269, 270, 273, 321, 326, 336, 340, 346, 349, 362, 378, 382, 398, 402, 405, 433, 446, 470, 478, 487, 496, 523, 530, 541, 583, 588, 603, 619, 629, 635, 654, 660, 661, 668– 670, 722, 725, 752, 765, 781]. Manager [104, 135, 163, 391]. Managers [427, 593]. Managing [40, 109, 123, 309, 365, 506]. MANO [631]. Manufacturing [745, 756]. Many [306, 440]. Many-Objective [440]. Many-Parallel-Task [306]. Mapping [4, 635]. MapReduce [265, 388, 447, 475, 476, 481, 489, 587, 595, 658, 721]. MapReduce-based [658]. Marine [760]. Market [148, 153, 224, 661]. Market-based [661]. Market-oriented [148]. Marketplace [150]. Markets [224, 365]. Markov [732]. Markovian [683, 737]. Mashroom [337]. Mashup [337]. Mass [68, 512, 555]. Massive [553, 604, 624]. Master [459]. Matching [222]. Matchmaking [81, 95, 299]. Mathematical [81, 644]. Matrix [568, 729]. Max [767]. May [593]. Mayfly [707]. MCDM [442]. MCDM-based [442]. MDS [167]. Means [328, 372, 554, 556]. Measurements [415]. Measuring [571]. MEC [684, 724, 765, 773]. Mechanism [154, 233, 546, 610, 618, 625, 643, 651, 703, 716]. Mechanisms [55, 85, 261, 345, 491, 742]. Media [482]. Mediating [57]. Medical [61, 130, 135, 283, 720]. Mediterranean [199]. Meet [89]. Memory [117, 254, 401, 669]. Merging [200]. Merrier [593]. Mesh [83]. Message [1, 107, 219, 226, 237]. Meta [63, 239, 304, 399, 428, 430, 608]. Meta-Brokering [430]. Meta-Heuristic [608]. Meta-Scheduler [63]. Meta-Schedulers [304]. Meta-Scheduling [399]. Meta-workflows [428]. Metadata [40, 136, 433, 483, 499, 590, 611]. Metamodel [323]. Metascheduling [89]. meteorological [423]. Method [165, 435, 473, 503, 513, 555, 559, 580, 655, 677, 687, 706, 712, 750]. Methodologies [341]. Methodology [569]. Methods [164, 690, 762, 766]. Metric [551]. Metrics [678]. MiCADO [636]. MiCADO-Edge [636]. Micro [507, 780]. Micro-Clouds [507]. Microarray [179]. MicroGrid [44]. Microservice [588, 778]. Microservices [638, 670, 727, 780]. Microservices-based [638]. Microsoft [387]. Middleware [5, 12, 60, 135, 173, 257, 282, 352, 470]. Migration [408, 475, 647].

Migration-Aware [475]. MINDS [351]. Minimization [189]. Minimize [761]. Minimizing [193, 489]. Mining [40, 333, 339, 346, 351, 366, 518, 528, 540, 554– 556, 680]. Mitigate [510]. Mitigation [452, 716]. Mix [373]. MobiByte [389]. Mobile [103, 236, 355, 359, 389–391, 434, 449, 450, 494, 503, 525, 532, 577, 605, 613, 617, 626, 645, 659, 672, 699, 708, 730, 732, 740, 744, 763, 783]. MobileNetV2 [756]. Mobility [648, 684]. Mobility-Aware [648]. Mobility-Driven [684]. MOCCAA [532]. Model [166, 207, 248, 259, 315, 322, 369, 376, 389, 448, 486, 515, 644, 651, 662, 671, 709, 714, 735, 743, 761]. Model-as-you-go [315]. Modeling [17, 89, 137, 143, 205, 265, 340, 394, 415, 431, 441, 460, 550, 668, 718]. Modeling-Learning-Based [550]. Modelling [218, 363, 416, 586, 663]. Models [144, 163, 242, 255, 340, 346, 380, 409, 548, 564, 686, 748, 772]. Modern [518]. Module [298]. Modulo [481]. Molecular [217, 569]. Moment [642]. Monitoring [8, 21, 57, 62, 167, 175, 206, 248, 308, 321, 375, 406, 523, 524, 721, 725, 762, 766]. Monte [202]. Monte-Carlo [202]. MoSGrid [433]. Most [186]. Motion [640]. Movement [501]. MPI [13, 107, 676]. MPNN [714]. MR [552]. Multi [34, 55, 75, 212, 240, 244, 266, 311, 333, 334, 356, 398, 426, 442, 457, 463, 464, 467, 470, 484, 508, 531, 542, 551, 560, 563, 592, 605, 648, 650, 662, 679, 689, 690, 708, 712, 723, 731, 745, 747, 748, 752, 764, 767, 771, 774, 792]. Multi-Access [648, 679]. Multi-Agent [531, 752, 774]. Multi-Agents [747]. Multi-Attribute [34]. Multi-Cloud [398, 426, 457, 463, 464, 592, 650]. Multi-Cluster [240, 792]. Multi-constraints [484]. Multi-Criteria [266]. Multi-dimensional [442, 662]. Multi-Domain [55, 333, 731]. Multi-elastic [508]. Multi-Grid [75]. Multi-job [334].

Multi-joint [712]. Multi-Language [311]. Multi-Level [723]. Multi-Objective [212, 356, 467, 542, 560, 679, 689, 745]. Multi-Physics [244]. Multi-Population [708]. Multi-Queueing [767]. Multi-Resource [771]. Multi-Streams [563]. Multi-Task [551, 690, 764]. Multi-Tasking [748]. Multi-Tenant [470]. Multi-User [75, 605]. Multi-zone [464]. Multicast [39, 517]. Multicloud [588]. Multicore [319]. Multicriteria [659]. Multicriteria-based [659]. Multimemetic [377]. Multimodal [609]. Multiplayer [501]. Multiple [140, 210, 267, 290, 342, 454, 564, 675, 735]. Multiplexing [159]. Multithreaded [298]. Music [162]. Muti [634]. Muti-Pathway [634]. Mutual [759].

NASA [11]. National [207]. Native [583, 702]. Natural [738]. NDN [574]. Near [562]. Near-Optimal [562]. Negotiating [168]. Negotiation [130, 168, 213, 566]. Negotiation-Based [213]. Nephele [293]. ness [785]. Nets [299, 531, 768]. Network [7, 8, 23, 34, 84, 99, 108, 113, 178, 194, 195, 221, 262, 372, 400, 415, 455, 466, 473, 507, 513, 603, 609, 631, 640, 652, 685, 696, 707, 723, 737, 746, 751, 758, 761, 772–774, 788]. Network-based [513]. Network-Sensitive [23]. Networking [24, 25, 513]. Networks [22, 26, 100, 114, 271, 272, 369, 439, 455, 500, 517, 548, 552, 555, 557, 574, 586, 587, 632, 653, 695, 718, 731, 763, 773, 783]. Neural [587, 643, 652, 737, 746, 751, 763, 788]. Neuroimage [313]. News [519]. Next [178, 692]. Ninf [5]. Ninf-G [5]. Node [755]. Nodes [188]. NOMA [645, 764]. NOMA-Enabled [645]. Non [222, 242, 504]. Non-periodic [222]. Non-Permissioned [504]. Non-Stationary [242]. Novel [100, 478, 552, 655, 696, 720, 722, 743]. NSGA [322]. Numerical [49].

O [317, 436, 477]. Object [611, 696]. Objective [212, 356, 393, 440, 467, 542, 560, 679, 689, 745]. Objects [706]. Occopus [463]. Off [752]. Off-Chain [752]. Offering [520]. Offloading [359, 577, 602, 605, 610, 617, 622, 626, 645, 648, 679, 683, 684, 690, 695, 699, 708, 724, 728, 732, 740, 744, 746, 754, 764, 767, 770, 783]. OGSA [63, 85, 109]. Omni [573]. On-Chain [752]. On-demand [272, 698]. Online [225, 297, 393, 500, 501, 510, 535, 625]. Only [246]. onto [4]. Ontology [381]. Ontology-based [381]. OP [573]. Open [85, 121, 129, 195, 231, 330, 411, 651, 671]. Open-Source [651, 671]. OpenStack [326]. Operating [254]. Operation [564, 725]. Operations [87, 436]. Opportunistic [233, 295, 589]. Opportunities [178]. Opportunity [364]. Optical [30, 675]. Optimal [189, 290, 562, 608, 638, 705, 734, 783]. Optimisation [38, 49, 712]. Optimised [474]. Optimising [35]. Optimization [13, 205, 221, 271, 307, 371, 408, 449, 454, 467, 475, 479, 481, 536, 601, 604, 615, 646, 682, 689, 707, 724, 753, 755, 758, 770, 785, 789, 792]. Optimization-Based [707]. Optimized [378, 512, 686, 739]. Optimizer [760]. Optimizing [74, 309, 532, 748]. OptorSim [38]. Orbweb [195]. Orchestrating [462, 464, 616, 697]. Orchestration [78, 343, 461, 466, 468, 620]. Orchestrator [463, 592, 636, 641]. Organisational [571]. Organizations [166, 175, 206]. Organized [208]. organizing [114, 139]. Orientation [519]. Oriented [142, 148, 343, 426, 465, 492, 648]. OSG [173]. Osmotic [641]. Outlier [366]. Output [567]. Overcommitted [670]. Overlay [100, 114]. Overlays [46, 113]. Oversubscription [669]. Own [405].

P [15, 75, 300]. P-GRADE [15, 75, 300].

P2P [22, 107, 154, 294, 365, 503]. PaaS [492]. PaaSword [443]. Package [681]. Package-Aware [681]. packing [785]. Pagerank [22]. PaGrid [83]. Pallet [756]. Pandemic [613]. Papers [112]. PAPMSC [402]. Parallel [35, 101, 107, 128, 141, 155, 211, 240, 276, 300, 306, 344, 366, 370, 374, 375, 401, 414, 448, 450, 468, 495, 515, 549, 554–556, 763]. Parallelism [271]. Parameter [218]. Parameters [582, 607, 675]. Parking [701, 744]. Partial [249, 388, 687]. Particle [559, 682, 739]. Partition [716]. Partitioner [83]. Partitioning [202, 648]. Passing [107, 237]. Passive [372]. Paths [113]. Pathway [634]. Pattern [222]. Patterns [339, 584, 663, 680]. Peer [16, 82, 93, 96, 98, 99, 105, 107, 125, 126, 139, 162, 164, 195, 211, 216, 295]. Peer-to-Peer [82, 93, 96, 99, 105, 107, 125, 126, 139, 162, 164, 195, 211, 216, 295]. Performability [529, 536]. Performance [7, 18, 21, 24, 25, 27, 29, 62, 97, 111, 113, 169, 181, 187, 196, 216, 221, 238, 251, 256, 294, 317, 341, 343, 346, 348, 388, 401, 402, 460, 571, 603, 705, 715, 741, 777, 780]. Performance-cost [221]. periodic [222]. Permission [570]. Permissioned [504]. PERSIST [470]. Personalities [535]. Perspective [368, 407, 577]. Petri [768]. PGRADE/gUSE [280]. Phase [434]. Physical [508, 685, 713]. Physics [132, 244, 612, 667, 700]. Physiological [759]. Physiology [360]. Pilot [612]. Pipeline [77]. Pipelines [678]. Placement [138, 294, 378, 397, 440, 475, 481, 507, 581, 594, 601, 614, 615, 651, 656, 671, 674, 791]. Plagiarism [296]. Plaintext [675]. Plan [383, 538]. Planning [255, 314, 318, 656]. Platform [292, 313, 317, 457, 480, 524, 537, 545, 556, 570, 580, 611, 638, 756]. Platforms [188, 468, 579, 600]. Pliant [395]. Pliant-based [395]. Poisoning [574]. Policies [109, 186, 378, 416, 436]. Policy [263, 470, 650, 774]. Policy-Based [470].

PoMic [670]. Pool [311, 376]. Pool-Based [376]. Popularity [472]. Population [708]. Portal [75, 87, 285, 300, 367]. Portal-based [87]. Portals [243]. Porting [13, 203]. Post [677]. Post-Quantum [677]. Power [400, 402, 668, 670, 757]. Power-Aware [402]. Powered [765]. Powerful [232]. Practical [160, 174, 201, 203]. Practices [544]. Precedence [370]. Precedence-Constrained [370]. Predictably [113]. Predicting [612]. Prediction [115, 184, 187, 441, 472, 523, 548, 607, 643, 649, 663, 664, 686, 690, 729, 730, 762, 775, 780, 782, 790]. Prediction-Based [184]. Predictions [63, 387]. Predictive [452]. Preface [41, 54, 71, 120, 131, 180, 268]. Preference [299]. Preference-Based [299]. Preservation [257, 703, 723]. Preserving [425, 473, 503, 720]. Preventive [274]. Price [610]. Pricing [147, 639]. PRIMA [53]. Prioritisation [233]. Prioritizing [116]. Priority [709]. Priority-Aware [709]. Privacy [425, 443, 473, 535, 574, 604, 703, 720]. Privacy-Aware [574]. Privacy-Preserving [473, 720]. Private [431, 701]. Private-Parking [701]. Proactive [581]. Probabilistic [767]. Problem [328, 679, 732]. Problems [121, 145, 379, 785]. Procedure [585]. Process [78, 278, 336, 369, 465, 550, 590, 784, 786]. Processes [298]. Processing [61, 314, 411, 476, 567, 582, 590, 691, 739]. Product [373]. Product-Mix [373]. Production [26, 91, 155, 203, 205, 227, 231]. Productivity [186]. Products [11]. Profile [173, 510]. Programmable [339, 539]. Programming [5, 14, 15, 117, 325, 329, 379, 450]. Programs [107]. Progress [217]. Project [58, 544]. Projects [187]. Proof [625]. Proportional [119]. Proportional-share [119]. Proposal [91]. Protection [574, 604]. Protein [387].

Proteins [511]. Protocol [28, 517, 566]. Prove [150]. Provenance [106, 276, 590]. Provenance-based [276]. Providers [435, 442]. Providing [223]. Provision [45]. Provisioning [161, 263, 400, 407, 439, 520, 566, 567]. pseudo [642]. pseudo-Zernike [642]. PSO [555]. Public [95, 403, 406]. Public-Resource [95]. Publications [417]. PUF [759]. Pulmonary [642]. Pupil [559]. Purpose [298]. Puzzle [657].

Q [695, 772–774]. Q-Inference [774]. Q-LSTM [772]. QCDgrid [69]. QCG [422]. QoS [45, 263, 348, 390, 466, 522, 607, 651]. QoS-Aware [466, 522, 651]. QoS-based [348]. Qualities [612]. Quality [25, 130, 169, 393, 544, 582, 654, 678, 687, 690, 696]. Quality-based [169]. Quality-related [687]. Quantum [69, 358, 433, 677]. Quattor [56]. Queries [70, 256]. Query [479]. Querying [563]. Questions [510]. Queueing [767]. QVIA [522]. QVIA-SDN [522].

Racking [756]. Rafhyc [324]. Random [448]. Randomized [652]. Range [256]. Rapid [128, 129]. Rates [567]. RCE [506]. RD [644]. Re [391]. Re-Encryption [391]. Reactive [274]. Read [246]. Read-Only [246]. Real [270, 308, 412, 517, 624, 716, 735]. Real-life [270]. Real-Time [308, 412, 517, 624, 716, 735]. Recognition [218, 609, 735]. Recognizing [587]. Recommendation [633]. Recovery [261, 529, 711]. Recurrent [548, 737, 788]. Redeployment [347]. Reducing [487]. Redundancy [278, 562]. Redundant [118, 431]. Refactoring [423]. Reference [5]. Refining [594]. Regional [199, 341]. Regional-Scale [199]. Registration [87]. Registry [327, 419]. Regression [379, 380]. Reinforcement [212, 602, 617, 647, 688, 732, 752, 757, 770].

Related [675, 687]. Relating [48]. Relational [57, 70]. Relativistic [358]. Reliability [241, 432]. Reliability-aware [432]. Reliable [39, 161, 354, 602, 611, 623, 693]. Remos [8]. Remote [117, 580]. REMUS [159]. Repair [278]. Repetitive [222]. Replica [17, 38, 58, 82, 138, 475]. Replicated [36, 666]. Replication [80, 184, 432, 451, 619]. Representation [361]. Reputation [146]. Requests [118]. Requirements [43, 48, 106, 220, 303, 533]. Rerouting [159]. Rescheduling [240, 261, 274, 384]. Research [145, 171, 207, 363, 419, 421, 423, 447, 528, 544, 549, 556, 596]. Reservation [538]. Reservations [213]. Residual [548]. Resilient [324]. ResNet [634]. Resource [21, 52, 63, 67, 69, 85, 95, 105, 115, 125, 142, 150, 154, 157, 161, 163, 170, 175, 184, 208, 224, 236, 255, 263, 274, 291, 340, 345, 347, 404, 445, 453, 455, 491, 492, 520, 523, 534, 541, 566, 599, 604, 607, 617, 624, 625, 635, 638, 639, 644, 646, 659, 661, 664, 668, 682, 722, 728, 765, 771, 772, 775, 782, 790, 792]. Resource-Aware [659]. Resource-Centric [85]. Resource-oriented [142]. Resources [40, 140, 177, 223, 227, 239, 285, 299, 342, 643, 672, 709, 712, 715]. Responsive [133, 212]. Restart [298, 625]. Results [59, 405, 569]. Retrieval [186]. Returns [573]. Reuse [627, 784, 786]. Review [353, 447, 541, 560, 584, 640]. Revisiting [360]. Reward [683]. Rewarding [613]. RGB [696]. RGB-D [696]. Riding [269]. Riemann [458]. Risk [158, 730]. RNN [768]. Road [473]. ROAM [104]. Robotic [791]. Robust [107, 237, 278, 311, 559]. Robustness [694]. Role [109, 170]. Role-Based [109, 170]. Routing [154, 471, 778]. RPC [5, 140]. RPC-based [5]. RT [616]. Rules [539]. Run [225, 267]. Running [66, 413]. Runtime [117, 252, 282, 445, 453].

S [167]. S-MDS [167]. SaaS [470]. SABUL [28]. Salient [696]. SAML [173]. SAML-XACML [173]. SAT [211]. Scalability [110, 715]. Scalable [50, 161, 182, 260, 293, 311, 483, 524, 563, 576, 647]. Scale [56, 77, 123, 126, 127, 183, 190, 199, 272, 300, 307, 349, 422, 515, 600, 667, 697, 718]. scaled [508]. Scaling [44, 353, 363, 387, 445, 491, 727, 734]. Scaling-out [491]. Scavenging [194]. Scenarios [464, 688]. Scheduler [63, 79, 119, 261, 595]. Schedulers [239, 304, 350]. Scheduling [6, 30, 38, 80, 102, 108, 128, 133, 140, 141, 143, 184, 197, 210, 213, 225, 247, 259, 266, 267, 275, 276, 291, 319, 322, 347, 356, 357, 370, 372, 383, 385, 393–396, 399, 404, 410, 432, 434, 436, 444, 454, 475, 477, 484, 489, 520, 521, 533, 534, 560– 562, 598–601, 606, 615, 618, 622, 654, 659, 681, 682, 689, 691, 704, 705, 707, 722, 725, 729, 745, 749, 763, 772, 775]. Scheme [196, 236, 278, 391, 476, 617, 693, 755]. Schemes [442]. School [194, 262]. Science [11, 106, 151, 229, 279, 281– 284, 288, 338, 386, 417, 421– 424, 426, 427, 429, 430, 433, 438, 566, 569, 635]. Scientific [44, 62, 64, 117, 162, 181, 213, 247, 251, 252, 270, 276, 300, 307, 308, 310, 315, 316, 338, 367, 382, 403, 414, 427, 428, 456, 463, 465, 622, 627, 630]. Screening [134]. SDN [522, 603, 718, 725]. SDN-based [522]. Sealed [610]. Sealed-Bid [610]. Seamless [480]. Search [22, 495, 512]. Searching [204, 296, 365]. Second [610]. Second-Price [610]. Secondary [642]. Secure [96, 135, 258, 285, 637, 693, 711, 768]. Secured [723, 725]. Securing [657]. Security [65, 209, 294, 322, 351, 381, 443, 574, 575, 677, 699, 703, 709, 779]. Security-enabled [351]. Segmentation [634, 710]. Selected [112]. Selecting [350, 565]. Selection [157, 169, 570]. Selective [145, 471]. Self [114, 139, 208, 653, 747, 760]. Self-Configuring [760]. Self-Driving [653, 747]. Self-Organized [208]. Self-organizing [114, 139]. Semantic [167, 519]. Semantics [314]. Sensitive [23, 606, 723]. Sensitivity [431]. Sensor [517, 555]. Sequence [515, 548]. Series [415, 643]. Serverless [579, 621, 678, 681, 692, 717, 787]. Servers [290, 402, 477, 513, 674]. Service [23, 25, 35, 37, 70, 78, 85, 87, 111, 126, 130, 136, 151, 158, 215, 233, 238, 249, 277, 321, 335, 343, 393, 407, 420, 435, 441, 442, 449, 496, 503, 506, 507, 527, 564, 578, 582, 614, 635, 651, 654, 698, 776, 783]. Service-Based [35]. Service-Centric [85]. Service-Level [578]. Services [34, 43, 49, 66, 81, 82, 84, 85, 105, 129, 130, 169, 227, 239, 285, 293, 300, 323, 324, 327, 342, 361, 363, 367, 390, 412, 429, 438, 443, 464, 565, 656, 674, 731, 734, 736, 747]. ServiceSs [329]. Set [409, 682]. Sets [465, 633]. Setup [223]. share [119]. ShareGrid [216]. Sharing [96, 115, 162, 258, 566, 576, 672]. Shibboleth [170]. SHIWA [313]. Short [413]. Shot [688]. Side [709]. Sign [243, 288, 494, 706]. Sign-On [243, 288, 494]. Signals [516]. Signature [782]. Signature-based [782]. Similarity [503, 658]. Simple [233]. SimSim [503]. Simulating [445]. Simulation [6, 61, 153, 368, 660]. Simulations [123, 202, 264, 341]. Simulator [217, 260, 525]. Simultaneous [80, 748]. Single [243, 288, 494]. Sites [591]. Sizing [27]. Skew [691, 716]. Sky [258]. Skyline [680]. SLA [289, 361, 628, 662]. SLA-aware [628]. SLA-based [289, 662]. SLAs [168, 705]. Slave [459]. Slicing [631]. SLO [277]. SLO-Fulfilling [277]. Smart [502, 578, 590, 649, 700, 701, 712, 720, 725, 726, 730, 736, 737, 745, 756, 761, 765, 766, 768, 773, 777, 781, 784, 786, 790]. Social [339, 500]. Socket [27]. SODALITE [616]. Software [204, 252, 277, 355, 358, 455, 466, 497, 513, 544, 557, 603, 627, 646, 711]. Software-defined [711]. SOG [208]. Solar [686]. Solution [243, 395, 405, 529]. Solutions [43, 303].

Solving [211, 328, 785]. Source [129, 330, 651, 671]. Sources [527]. South [228]. South-Eastern [228]. Space [515]. Sparbit [676]. Spark [448, 493, 555, 680, 716]. Spark-based [493]. Sparse [642]. Spatial [366, 503]. Spatiotemporal [673, 714]. Special [18, 24, 93, 112, 120, 180, 250, 305, 325, 335, 417, 438, 461, 469, 488, 509, 530, 572]. Specification [320]. Speculative [388]. Speech [609]. Speed [30, 271]. Sporadic [519]. Spot [189, 224, 698]. Spot-checking [189]. Squares [687]. SRB [174]. SRM [174]. Stacked [642]. Stacks [26]. Stage [689]. Standard [12, 657]. Standard-512 [657]. Standardization [482]. Standardized [331]. Standards [2, 192, 195, 282, 323]. Standards-Based [192, 282]. STAR [301]. State [148, 667, 683]. State-of-the-Art [148, 667]. Static [163, 454, 558]. Stationary [242]. Statistical [137, 265, 334]. Stealing [157]. Stochastic [369, 373, 402, 445, 577, 599, 628]. Stochastic-based [577]. Stopping [783]. Storage [61, 68, 98, 163, 170, 174, 209, 249, 364, 410, 469–471, 473, 477, 479, 482, 597, 666, 693, 791]. Storage-aware [410]. Store [611]. Stores [256]. Storing [590]. Storm [269, 618, 719]. Straggler [587]. Strategic [741]. Strategies [38, 138, 225, 240, 267, 295]. Strategy [238, 521, 602, 604, 625, 626, 699, 744, 749, 755, 757, 773]. Strategy-Proof [625]. Stream [491, 618]. Streaming [775]. Streamlining [87]. Streams [113, 333, 563]. Strengths [340]. Strongly [129]. Structural [287, 288]. Structures [387]. Student [721]. Studies [6, 214]. Study [48, 181, 187, 194, 270, 308, 364, 414, 495, 635]. Study-based [635]. Studying [377]. Subdividing [413]. Submission [205, 331]. Subscription [100]. Substrate [195]. Successful [420]. Successive [409]. SuMo [371]. Supercomputing [234]. Superscalar [14]. Supervised [584]. Supply [758]. Support [84, 117, 252, 292, 298, 301, 306, 381, 424, 428, 433, 456, 544, 564, 613, 644]. Supporting [143, 539]. Surveillance [502, 553]. Survey [345, 382, 403–405, 514, 523, 577, 614, 639, 688]. Sustainable [726, 738]. Swarm [235, 682, 739, 755]. Sweep [218]. Symposium [18]. Synergy [781]. System [8, 21, 53, 59, 96, 98, 99, 124, 139, 140, 157, 159, 167, 175, 182, 199, 209, 270, 306, 320, 350, 351, 354, 381, 412, 454, 465, 481, 483, 485, 523, 568, 586, 594, 626, 632, 633, 660, 685, 701, 713, 721, 725, 730, 735, 767]. System-Level [21]. Systematic [585, 635]. Systems [22, 42, 73, 115, 127, 140, 192, 220, 254, 278, 319, 355, 380, 384, 401, 411, 426, 432, 437, 477, 525, 597, 617, 666, 720, 741, 774, 777, 787, 792]. SZDG [182]. SZTAKI [182].

Table [246]. Tags [339]. Tailed [255]. Take [279]. Take-up [279]. Tank [537]. Target [750]. Task [128, 184, 186, 306, 348, 396, 432, 445, 468, 489, 551, 560, 598, 645, 689, 690, 699, 707, 722, 725, 732, 744–746, 749, 754, 763, 764, 769, 770, 787]. Task-Based [468, 489]. Tasking [748]. Tasks [476, 484, 587, 600, 606, 679]. Taxonomy [73, 411]. Taylor [707]. TCP [26]. Technical [152]. Technique [575, 643, 658, 739]. Techniques [56, 82, 162, 353, 380, 514]. Technologies [200, 635, 639, 738]. Technology [126, 241, 590, 721, 741]. Telescope [386]. Tenant [470]. Testbed [61]. Testcase [264]. Text [516, 518]. Their [51, 60]. Theophys [302]. Theoretic [275]. Theoretical [332]. Theories [572]. Theory [682, 713, 783]. Thermal [273]. Thermal-Aware [273]. Things [575, 604, 711, 718, 720, 733, 751, 752, 765, 769]. Threat [452, 713]. Threats [768]. Thresholding [459]. Throughout [400]. Throughput [215, 271, 272, 307]. Tightly [240]. Tightly-Coupled [240]. Time [141, 189, 193, 225, 267, 308, 383, 401, 412, 415, 454, 517, 534, 624, 625, 643, 716, 735]. Time-Series [415]. Time-Varying [625]. Times [445, 487]. Timing [452]. Tolerance [20, 307, 377]. Tolerant [88, 125, 332, 384, 543, 562]. Tool [35, 116]. Toolkit [129, 281, 445, 564]. Tools [56, 110, 368]. Topic [490]. Topology [397, 446, 582, 588]. Topology-Aware [397]. TOPSIS [435, 603]. TORCH [592]. TOSCA [592]. TOSCA-Based [592]. Towing [537]. Trace [334, 655]. Tracing [596]. Traditional [255]. Traffic [714, 781]. Training [585, 595]. Trajectory [473, 673]. Transfer [249, 511]. Transfers [27, 30]. Transform [675]. Transformation [336, 488, 743]. Transformation-Based [336]. Translated [296]. Translation [643]. Transparent [298, 468]. Transport [28, 30, 624]. Transportation [761]. Tree [495, 597]. TRENCADIS [283]. Trends [528, 584]. Triana [16, 72]. Trolling [490]. Troops [679]. Trust [65, 95, 146, 381, 435, 442, 662]. Trusted [65]. Trustworthiness [435, 442]. Trustworthy [96]. Tuberculosis [642]. Tumor [689]. Tumour [634]. Tuned [760]. Tuning [344]. Turbine [418]. Turnaround [128]. Tweet [652]. Twelfth [18]. Twitter [514]. Two [85, 145, 207, 434, 689, 785]. Two-dimensional [785]. Two-Layer [207]. Two-Phase [434]. Two-Stage [689]. Typed [129]. Typology [535].

Ultra [623]. Ultra-Reliable [623]. umd [497]. umd-verification [497]. Uncertain [540]. Uncertainty [337, 453]. UNet [634]. Unified [669]. Unite [94]. Unmodified [254]. Unreliable [185]. Unstructured [644]. Urban [425, 781]. Usability [283]. Usage [111, 255, 264, 664, 782]. Use [60, 95, 97, 194, 288, 300]. User [75, 87, 225, 241, 267, 280, 285, 349, 398, 568, 605].

User-friendly [285]. Users [86, 391]. Using [12, 38, 78, 106, 109, 170, 186, 205, 223, 233, 248, 266, 272, 301, 308, 322, 356, 361, 378, 380, 421, 422, 502, 512, 516, 526, 531, 538, 542, 552, 555, 569, 581, 587, 602, 603, 607, 610, 617, 626, 643, 645, 647, 653, 655, 673, 683, 685, 686, 698, 701, 703, 706, 713, 721, 723, 726, 730, 738, 739, 742–744, 748, 751, 762, 763, 776, 780, 782, 790]. Utilities [466]. Utility [148, 158, 459]. Utilization [274, 772]. Utilizing [761].

v2.2 [163]. Vague [682]. Validating [44]. Validation [352, 497]. Value [256]. Values [458]. VANET [761]. Variable [413]. Variable-Length [413]. Variety [280]. Varying [625]. Vector [84, 515]. Vega [42]. Vehicles [653, 740, 746, 747, 754, 764, 770, 773, 779]. Vehicular [574]. Veriblock [573]. Verification [497, 551]. Via [117, 139, 177, 402, 642]. Video [314, 502, 553, 559, 678]. View [229, 523]. Views [368]. Vine [281]. Viola [585]. Virtual [114, 134, 166, 175, 206, 221, 254, 297, 362, 370, 395, 397, 408, 419, 440, 455, 485, 486, 508, 522, 537, 539, 581, 591, 647, 671, 753, 783]. Virtualization [110, 492]. Virtualized [239, 273, 370, 402, 484, 646]. Vision [710]. Visual [72, 375, 697, 706]. Visualization [249]. VizLitG [249]. VM [446, 474, 538, 670]. VM-Microservices [670]. VMMB [254]. VMs [487]. Vol [31, 32]. Volunteer [180, 183, 185, 186, 189, 190, 237, 495, 545]. VOs [207]. Voting [504]. Vulnerability [586].

Ways [85]. WDM [400]. Weaknesses [340]. Weather [84]. Web [43, 81, 105, 243, 323, 367, 402, 429, 537]. WebAssembly [692]. WeNMR [287]. Whale [724]. Wide [114, 141, 214, 400]. Widow [755]. Wien2k [203]. WiGriMMA [248]. Wind [418]. Wireless [248, 517].

WISE [200]. within [16, 466, 650]. WMSN [759]. Wolf [785]. Work [772]. Workers [185]. Workflow [37, 72–74, 76, 77, 155, 177, 267, 270, 291, 305, 308, 311, 313, 314, 316, 318, 356, 357, 364, 382, 410, 426, 427, 465, 531, 536, 560– 562, 622, 630, 660, 698, 729, 776]. Workflow-Aware [364]. Workflow-Based [37, 155]. Workflow-Nets [531]. Workflow-Oriented [426]. Workflows [4, 62, 75, 210, 213, 218, 251, 276, 306– 310, 312, 315, 332, 338, 383, 414, 427, 428, 445, 454, 621, 784, 786]. Workload [59, 205, 265, 542, 547, 643, 705]. Workloads [520, 521]. Workshops [417]. Workstations [50, 114, 181]. Workunits [413]. World [94, 327, 514]. WOW [114]. WS [168, 280, 343]. WS-Agreement [168]. WS-PGRADE [280]. WS-PGRADE/gUSE [280]. WSANs [478]. WSN [755]. WSRF [161]. WSRF-Compliant [161]. WSRF.NET [142].

XACML [173, 258]. XMatch [169]. XMPP [301]. XtremWeb [176].

ZENTURIO [35]. Zernike [642]. Zero [546]. Zeta [458]. zone [464].

References

Foster:2003:EM

[1] Ian Foster and Peter Kacsuk. Editors’ message. Journal of Grid Computing, 1(1):1–2, ???? 2003. CODEN ???? ISSN 1570-7873 (print), 1572-9184

(electronic). URL http://ipsapp009.
kluweronline.com/content/getfile/
6160/1/8/abstract.htm; http://ipsapp009.
kluweronline.com/content/getfile/
6160/1/8/fulltext.pdf.

Catlett:2003:SGC

[2] Charlie Catlett. Standards for grid computing: Global Grid Forum. Journal of Grid Computing, 1(1):3–7, ???? 2003. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (elec-

tronic).  URL http://ipsapp009.
kluweronline.com/content/getfile/
6160/1/7/abstract.htm; http://ipsapp009.
kluweronline.com/content/getfile/
6160/1/7/fulltext.pdf.

Nemeth:2003:CGA

[3] Zsolt Németh and Vaidy Sunderam. Characterizing Grids: Attributes, definitions, and formalisms. Journal of Grid Computing, 1(1):9–23, ???? 2003. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (elec-

tronic).  URL http://ipsapp009.
kluweronline.com/content/getfile/
6160/1/9/abstract.htm; http://ipsapp009.
kluweronline.com/content/getfile/
6160/1/9/fulltext.pdf.

Deelman:2003:MAC

[4] Ewa Deelman, James Blythe, Yolanda Gil, Carl Kesselman, Gaurang Mehta, Karan Vahi, Kent Blackburn, Albert Lazzarini, Adam Arbree, Richard Cavanaugh, and Scott Koranda. Mapping abstract complex workflows onto grid environments. Journal of Grid Computing, 1(1):25–39, ???? 2003. CODEN ???? ISSN 1570-7873 (print), 1572-9184

(electronic). URL http://ipsapp009.
kluweronline.com/content/getfile/
6160/1/2/abstract.htm; http://ipsapp009.
kluweronline.com/content/getfile/
6160/1/2/fulltext.pdf.

Tanaka:2003:NGR

[5] Y. Tanaka, H. Nakada, S. Sekiguchi, T. Suzumura, and S. Matsuoka. Ninf-G: a reference implementation of RPCbased programming middleware for grid computing. Journal of Grid Computing, 1(1):41–51, ???? 2003. CODEN ???? ISSN 1570-7873 (print), 1572-9184

(electronic). URL http://ipsapp009.
kluweronline.com/content/getfile/
6160/1/3/abstract.htm; http://ipsapp009.
kluweronline.com/content/getfile/
6160/1/3/fulltext.pdf.

Ranganathan:2003:SSC

[6] Kavitha Ranganathan and Ian Foster. Simulation studies of computation and data scheduling algorithms for data Grids. Journal of Grid Computing, 1(1):53–62, ???? 2003. CODEN ???? ISSN 1570-7873 (print), 1572-9184

(electronic). URL http://ipsapp009.
kluweronline.com/content/getfile/
6160/1/4/abstract.htm; http://ipsapp009.
kluweronline.com/content/getfile/
6160/1/4/fulltext.pdf.

Feng:2003:AFC

[7] Wu chun Feng, Mark K. Gardner, Michael E. Fisk, and Eric H. Weigle. Automatic flow-control adaptation for enhancing network performance in computational Grids. Journal of Grid Computing, 1(1):63–74, ???? 2003. CODEN ???? ISSN 1570-7873 (print), 1572-9184

(electronic). URL http://ipsapp009.
kluweronline.com/content/getfile/
6160/1/5/abstract.htm; http://ipsapp009.
kluweronline.com/content/getfile/
6160/1/5/fulltext.pdf.

Lowekamp:2003:DIE

[8] Bruce Lowekamp, Nancy Miller, Roger Karrer, Thomas Gross, and Peter Steenkiste. Design, implementation, and evaluation of the Remos Network Monitoring System. Journal of Grid Computing, 1(1):75–93, ???? 2003. CODEN ???? ISSN 1570-7873 (print), 1572-9184

                                        (electronic). URL http://ipsapp009.
                                        kluweronline.com/content/getfile/
                                        6160/1/6/abstract.htm; http://ipsapp009.
                                        kluweronline.com/content/getfile/
                                        6160/1/6/fulltext.pdf.

Anonymous:2003:IA

[9] Anonymous. Instructions for authors. Journal of Grid Computing, 1(1):95– 98, ???? 2003. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (elec-

                                        tronic).  URL http://ipsapp009.
                                        kluweronline.com/content/getfile/
                                        6160/1/1/abstract.htm; http://ipsapp009.
                                        kluweronline.com/content/getfile/
                                        6160/1/1/fulltext.pdf.

Kielmann:2003:E

[10] Thilo Kielmann. Editorial. Journal of Grid Computing, 1(2):99–100, ???? 2003. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (elec-

                                        tronic).  URL http://ipsapp008.
                                        kluweronline.com/IPS/content/ext/
                                        x/J/6160/I/4/A/1/abstract.htm.

Barkstrom:2003:DGN

[11] Bruce R. Barkstrom, Thomas H. Hinke, Shradha Gavali, Warren Smith, William J. Seufzer, Chaumin Hu, and David E. Cordner. Distributed generation of NASA earth science data products. Journal of Grid Computing, 1(2):101–116, ???? 2003. CODEN ???? ISSN 1570-7873 (print), 1572-9184

(electronic). URL http://ipsapp008.
kluweronline.com/IPS/content/ext/
x/J/6160/I/4/A/2/abstract.htm.

Takemiya:2003:CGA

[12] H. Takemiya, K. Shudo, Y. Tanaka, and S. Sekiguchi. Constructing grid applications using standard grid middleware. Journal of Grid Computing, 1(2):117–131, ???? 2003. CODEN ???? ISSN 1570-7873 (print), 1572-9184

(electronic). URL http://ipsapp008.
kluweronline.com/IPS/content/ext/
x/J/6160/I/4/A/3/abstract.htm.

Keller:2003:TEE

[13] Rainer Keller, Edgar Gabriel, Bettina Krammer, Matthias S. Müller, and Michael M. Resch. Towards efficient execution of MPI applications on the grid: Porting and optimization issues. Journal of Grid Computing, 1(2):133–149, ???? 2003. CODEN ???? ISSN 1570-7873 (print), 1572-9184

(electronic). URL http://ipsapp008.
kluweronline.com/IPS/content/ext/
x/J/6160/I/4/A/4/abstract.htm.

Badia:2003:PGA

[14] Rosa M. Badia, Jesús Labarta, Raül Sirvent, Josep M. Pérez, José M. Cela, and Rogeli Grima. Programming grid applications with GRID superscalar. Journal of Grid Computing, 1(2):151–170, ???? 2003. CODEN ???? ISSN 1570-7873 (print), 1572-9184

(electronic). URL http://ipsapp008.
kluweronline.com/IPS/content/ext/
x/J/6160/I/4/A/5/abstract.htm.

Kacsuk:2003:PGG

[15] P. Kacsuk, G. Dózsa, J. Kovács, R. Lovas, N. Podhorszki, Z. Balaton, and G. Gombás. P-GRADE: a grid programming environment. Journal of Grid Computing, 1(2):171–197, ???? 2003. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (elec-

                                        tronic).  URL http://ipsapp008.
                                        kluweronline.com/IPS/content/ext/
                                        x/J/6160/I/4/A/6/abstract.htm.

Taylor:2003:TAW

[16] Ian Taylor, Matthew Shields, Ian Wang, and Omer Rana. Triana applications within grid computing and peer to peer environments. Journal of Grid Computing, 1(2):199–217, ???? 2003. CODEN ???? ISSN 1570-7873 (print), 1572-9184

                                        (electronic). URL http://ipsapp008.
                                        kluweronline.com/IPS/content/ext/
                                        x/J/6160/I/4/A/7/abstract.htm.

Schintke:2003:MRA

[17] Florian Schintke and Alexander Reinefeld. Modeling replica availability in large data Grids. Journal of Grid Computing, 1(2):219–227, ???? 2003. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (elec-

                                        tronic).  URL http://ipsapp008.
                                        kluweronline.com/IPS/content/ext/
                                        x/J/6160/I/4/A/8/abstract.htm.

Chien:2003:ISI

[18] Andrew A. Chien. Introduction: Special issue: Highlights of the Twelfth Symposium on High Performance Distributed Computing. Journal of Grid Computing, 1(3):229, ???? 2003. CODEN ???? ISSN 1570-7873 (print), 1572-9184

                                        (electronic). URL http://ipsapp008.
                                        kluweronline.com/IPS/content/ext/
                                        x/J/6160/I/9/A/5/abstract.htm.

Talwar:2003:AEE

[19] Vanish Talwar, Sujoy Basu, and Raj Kumar. Architecture and environment for enabling interactive Grids. Journal of Grid Computing, 1(3):231–250, ???? 2003. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (elec-

tronic).  URL http://ipsapp008.
kluweronline.com/IPS/content/ext/
x/J/6160/I/9/A/3/abstract.htm.

Hwang:2003:FFF

[20] Soonwook Hwang and Carl Kesselman. A flexible framework for fault tolerance in the grid. Journal of Grid Computing, 1(3):251–272, ???? 2003. CODEN ???? ISSN 1570-7873 (print), 1572-9184

(electronic). URL http://ipsapp008.
kluweronline.com/IPS/content/ext/
x/J/6160/I/9/A/2/abstract.htm.

Agarwala:2003:SLR

[21] Sandip Agarwala, Christian Poellabauer, Jiantao Kong, Karsten Schwan, and Matthew Wolf. System-level resource monitoring in high-performance computing environments. Journal of Grid Computing, 1(3):273–289, ???? 2003. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (elec-

tronic).  URL http://ipsapp008.
kluweronline.com/IPS/content/ext/
x/J/6160/I/9/A/4/abstract.htm.

Sankaralingam:2003:PCK

[22] Karthikeyan Sankaralingam, Madhulika Yalamanchi, Simha Sethumadhavan, and James C. Browne. Pagerank computation and keyword search on distributed systems and P2P networks. Journal of Grid Computing, 1(3):291–307, ???? 2003. CODEN ???? ISSN 1570-7873 (print), 1572-9184

(electronic). URL http://ipsapp008.
kluweronline.com/IPS/content/ext/
x/J/6160/I/9/A/6/abstract.htm.

Huang:2003:NSS

[23] An-Cheng Huang and Peter Steenkiste. Network-sensitive service discovery. Journal of Grid Computing, 1(3):309– 326, ???? 2003. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (elec-

                                        tronic).  URL http://ipsapp008.
                                        kluweronline.com/IPS/content/ext/
                                        x/J/6160/I/9/A/1/abstract.htm.

Crowcroft:2003:ESI

[24] Jon Crowcroft and Saleem Bhatti. Editorial: Special issue: High performance networking. Journal of Grid Computing, 1(4):327, ???? 2003. CODEN ???? ISSN 1570-7873 (print), 1572-9184

                                        (electronic). URL http://ipsapp008.
                                        kluweronline.com/IPS/content/ext/
                                        x/J/6160/I/11/A/7/abstract.htm.

Rio:2003:QSN

[25] Miguel Rio, Andrea di Donato, Frank Saka, Nicola Pezzi, Richard Smith, Saleem Bhatti, and Peter Clarke. Quality of service networking for high performance grid applications. Journal of Grid Computing, 1(4):329–343, ???? 2003. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (elec-

                                        tronic).  URL http://ipsapp008.
                                        kluweronline.com/IPS/content/ext/
                                        x/J/6160/I/11/A/1/abstract.htm.

Bullot:2003:EAT

[26] Hadrien Bullot, R. Les Cottrell, and Richard Hughes-Jones. Evaluation of advanced TCP stacks on fast longdistance production networks. Journal of Grid Computing, 1(4):345–359, ???? 2003. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (elec-

                                        tronic).  URL http://ipsapp008.

kluweronline.com/IPS/content/ext/
x/J/6160/I/11/A/5/abstract.htm.

Prasad:2003:SBA

[27] Ravi S. Prasad, Manish Jain, and Constantinos Dovrolis. Socket buffer autosizing for high-performance data transfers. Journal of Grid Computing, 1 (4):361–376, ???? 2003. CODEN ???? ISSN 1570-7873 (print), 1572-9184

(electronic). URL http://ipsapp008.
kluweronline.com/IPS/content/ext/
x/J/6160/I/11/A/4/abstract.htm.

Gu:2003:STP

[28] Yunhong Gu and Robert Grossman. SABUL: a transport protocol for grid computing. Journal of Grid Computing, 1(4):377–386, ???? 2003. CODEN ???? ISSN 1570-7873 (print), 1572-9184

(electronic). URL http://ipsapp008.
kluweronline.com/IPS/content/ext/
x/J/6160/I/11/A/3/abstract.htm.

Boutaba:2003:GCL

[29] Raouf Boutaba, Wojciech Golab, Youssef Iraqi, Tianshu Li, and Bill St.Arnaud. Grid-controlled lightpaths for high performance grid applications. Journal of Grid Computing, 1(4):387–394, ???? 2003. CODEN ???? ISSN 1570-7873 (print), 1572-9184

(electronic). URL http://ipsapp008.
kluweronline.com/IPS/content/ext/
x/J/6160/I/11/A/2/abstract.htm.

Veeraraghavan:2003:STF

[30] Malathi Veeraraghavan, Xuan Zheng, Wu chun Feng, Hojun Lee, Edwin K. P. Chong, and Hua Li. Scheduling and transport for file transfers on high-speed optical circuits. Journal of Grid Computing, 1(4):395–405, ???? 2003. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (elec-

tronic).  URL http://ipsapp008.
kluweronline.com/IPS/content/ext/
x/J/6160/I/11/A/6/abstract.htm.

Anonymous:2003:AIV

[31] Anonymous. Author index vol. 1 (2003). Journal of Grid Computing, 1(4):407–408, ???? 2003. CODEN ???? ISSN 1570-7873 (print), 1572-9184

(electronic). URL http://ipsapp008.
kluweronline.com/IPS/content/ext/
x/J/6160/I/11/A/9/abstract.htm.

Anonymous:2003:CV

[32] Anonymous. Contents of vol. 1 (2003). Journal of Grid Computing, 1(4):409– 411, ???? 2003. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (elec-

tronic).  URL http://ipsapp008.
kluweronline.com/IPS/content/ext/
x/J/6160/I/11/A/8/abstract.htm.

Stockinger:2004:E

[33] Heinz Stockinger. Editorial. Journal of Grid Computing, 2(1):1, March 2004. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1570-7873&
volume=2&issue=1&spage=1.

Cai:2004:MMA

[34] Min Cai, Martin Frank, Jinbo Chen, and Pedro Szekely. MAAN: a multiattribute addressable network for grid information services. Journal of Grid Computing, 2(1):3–14, March 2004. CO-DEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1570-7873&
volume=2&issue=1&spage=3.

Prodan:2004:ZGS

[35] Radu Prodan and Thomas Fahringer. ZENTURIO: a grid service-based tool for optimising parallel and grid applications. Journal of Grid Computing, 2(1):15–29, March 2004. CO-DEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1570-7873&
volume=2&issue=1&spage=15.

Vazhkudai:2004:DDB

[36] Sudharshan Vazhkudai. Distributed downloads of bulk, replicated grid data. Journal of Grid Computing, 2(1):31–42, March 2004. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic).

URL http://www.springerlink.com/
openurl.asp?genre=article&issn=
1570-7873&volume=2&issue=1&spage=
31.

Kim:2004:WBA

[37] Seung-Hyun Kim, Kyong Hoon Kim, Jong Kim, Sung-Je Hong, and Sangwan Kim. Workflow-based authorization service in the grid. Journal of Grid Computing, 2(1):43–55, March 2004. CO-DEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1570-7873&
volume=2&issue=1&spage=43.

Cameron:2004:ASR

[38] D. G. Cameron, A. P. Millar, C. Nicholson, R. Carvajal-Schiaffino, K. Stockinger, and F. Zini. Analysis of scheduling and replica optimisation strategies for data grids using OptorSim. Journal of Grid Computing, 2(1):57–69, March 2004. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=1570-7873&
                                        volume=2&issue=1&spage=57.

Maimour:2004:EAR

[39] M. Maimour and C. Pham. Experimenting active reliable multicast on application-aware grids. Journal of Grid Computing, 2(1):71–83, March 2004. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=1570-7873&
                                        volume=2&issue=1&spage=71.

Mastroianni:2004:MMG

[40] Carlo Mastroianni, Domenico Talia, and Paolo Trunfio. Metadata for managing grid resources in data mining applications. Journal of Grid Computing, 2(1):85–102, March 2004. CO-DEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=1570-7873&
                                        volume=2&issue=1&spage=85.

Sun:2004:P

[41] Xian-He Sun and Minglu Li. Preface. Journal of Grid Computing, 2 (2):107–108, June 2004. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=1570-7873&
                                        volume=2&issue=2&spage=107.

Xu:2004:VCS

[42] Zhiwei Xu, Wei Li, Li Zha, Haiyan Yu, and Donghua Liu. Vega: a computer systems approach to grid computing. Journal of Grid Computing, 2(2):109–120, June 2004. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic).

URL http://www.springerlink.com/
openurl.asp?genre=article&issn=
1570-7873&volume=2&issue=2&spage=
109.

Zhang:2004:RDD

[43] Liang-Jie Zhang and Bing Li. Requirements driven dynamic services composition for Web services and grid solutions. Journal of Grid Computing, 2(2):121–140, June 2004. CO-DEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1570-7873&
volume=2&issue=2&spage=121.

Liu:2004:VSM

[44] Xin Liu, Huaxia Xia, and Andrew A. Chien. Validating and scaling the MicroGrid: a scientific instrument for grid dynamics. Journal of Grid Computing, 2(2):141–161, June 2004. CO-DEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1570-7873&
volume=2&issue=2&spage=141.

Al-Ali:2004:APQ

[45] Rashid J. Al-Ali, Kaizar Amin, Gregor von Laszewski, Omer F. Rana, David W. Walker, Mihael Hategan, and Nestor Zaluzec. Analysis and provision of QoS for distributed grid applications. Journal of Grid Computing, 2(2):163–182, June 2004. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic).

URL http://www.springerlink.com/
openurl.asp?genre=article&issn=
1570-7873&volume=2&issue=2&spage=
163.

Liu:2004:BEO

[46] Yunhao Liu, Li Xiao, Lionel M. Ni, and Yunhuai Liu. Building efficient overlays. Journal of Grid Computing, 2(2):183–192, June 2004. CO-DEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=1570-7873&
                                        volume=2&issue=2&spage=183.

Yang:2004:GCC

[47] Guangwen Yang, Hai Jin, Minglu Li, Nong Xiao, Wei Li, Zhaohui Wu, Yongwei Wu, and Feilong Tang. Grid computing in China. Journal of Grid Computing, 2(2):193–206, June 2004. CO-DEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=1570-7873&
                                        volume=2&issue=2&spage=193.

Finkelstein:2004:RRA

[48] Anthony Finkelstein, Clare Gryce, and Joe Lewis-Bowen. Relating requirements and architectures: a study of data-grids. Journal of Grid Computing, 2(3):207–222, September 2004. CO-DEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=1570-7873&
                                        volume=2&issue=3&spage=207.

Xue:2004:NOG

[49] Gang Xue, Wenbin Song, Simon J. Cox, and Andy Keane. Numerical optimisation as grid services for engineering design. Journal of Grid Computing, 2(3):223–238, September 2004. CO-DEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1570-7873&
volume=2&issue=3&spage=223.

Purusothaman:2004:DSH

[50] T. Purusothaman, S. Annadurai, H. Vijay Ganesh, C. T. Chockalingam, and B. Uthra Kumar. Dynamically scalable, heterogeneous and generic architecture for a grid of workstations. Journal of Grid Computing, 2(3):239–246, September 2004. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic).

URL http://www.springerlink.com/
openurl.asp?genre=article&issn=
1570-7873&volume=2&issue=3&spage=
239.

Roblitz:2004:AML

[51] Thomas Röblitz, Florian Schintke, Alexander Reinefeld, Olof Bärring, Maite Barroso Lopez, German Cancio, Sylvain Chapeland, Karim Chouikh, Lionel Cons, Piotr Poznański, Philippe Defert, Jan Iven, Thorsten Kleinwort, Bernd Panzer-Steindel, Jaroslaw Polok, Catherine Rafflin, Alan Silverman, Tim Smith, Jan Eldik, David Front, Massimo Biasotto, Cristina Aiftimiei, Enrico Ferro, Gaetano Maron, Andrea Chierici, Luca dell’Agnello, Marco Serra, Michele Michelotto, Lord Hess, Volker Lindenstruth, Frank Pister, Timm Morten Steinbeck, David Groep, Martijn Steenbakkers, Oscar Koeroo, Wim Som de Cerff, Gerben Venekamp, Paul Anderson, Tim Colles, Alexander Holt, Alastair Scobie, Michael George, Andrew Washbrook, and Rafael A. García Leiva. Autonomic management of large clusters and their integration into the grid. Journal of Grid Computing, 2 (3):247–260, September 2004. CO-DEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=1570-7873&
                                        volume=2&issue=3&spage=247.

Zhu:2004:DGR

[52] Cheng Zhu, Zhong Liu, Weiming Zhang, Weidong Xiao, Zhenning Xu, and Dongsheng Yang. Decentralized grid resource discovery based on resource information community. Journal of Grid Computing, 2(3):261–277, September 2004. CO-DEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=1570-7873&
                                        volume=2&issue=3&spage=261.

Lorch:2004:PGA

[53] Markus Lorch and Dennis Kafura. The PRIMA grid authorization system. Journal of Grid Computing, 2 (3):279–298, September 2004. CO-DEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=1570-7873&
                                        volume=2&issue=3&spage=279.

Laure:2004:P

[54] Erwin Laure. Preface. Journal of Grid Computing, 2(4):299–300, December 2004. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic).

                                        URL http://www.springerlink.com/
                                        openurl.asp?genre=article&issn=
                                        1570-7873&volume=2&issue=4&spage=
                                        299.

Cornwall:2004:AAM

[55] Linda A. Cornwall, Jens Jensen, David P. Kelsey, Ákos Frohner, Daniel Kouřil, Franck Bonnassieux, Sophie Nicoud, Károly Lõrentey, Joni Hahkala, Mika Silander, Roberto Cecchini, Vincenzo Ciaschini, Luca dell’Agnello, Fabio Spataro, David O’Callaghan, Olle Mulmo, Gian Luca Volpato, David Groep, Martijn Steenbakkers, and Andrew McNab. Authentication and authorization mechanisms for multidomain grid environments. Journal of Grid Computing, 2(4):301–311, December 2004. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic).

URL http://www.springerlink.com/
openurl.asp?genre=article&issn=
1570-7873&volume=2&issue=4&spage=
301.

Leiva:2004:QTT

[56] R. García Leiva, M. Barroso López, G. Cancio Meliá, B. Chardi Marco, L. Cons, P. Poznański, A. Washbrook, E. Ferro, and A. Holt. Quattor: Tools and techniques for the configuration, installation and management of largescale grid computing fabrics. Journal of Grid Computing, 2(4):313–322, December 2004. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic).

URL http://www.springerlink.com/
openurl.asp?genre=article&issn=
1570-7873&volume=2&issue=4&spage=
313.

Cooke:2004:RGM

[57] A. W. Cooke, A. J. G. Gray, W. Nutt, J. Magowan, M. Oevers, P. Taylor, R. Cordenonsi, R. Byrom, L. Cornwall, A. Djaoui, L. Field, S. M. Fisher, S. Hicks, J. Leake, R. Middleton, A. Wilson, X. Zhu, N. Podhorszki, B. Coghlan, S. Kenny, D. O’Callaghan, and J. Ryan. The relational grid monitoring architecture: Mediating information about the grid. Journal of Grid Computing, 2(4):323–339, December 2004. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=1570-7873&
                                        volume=2&issue=4&spage=323.

Cameron:2004:RME

[58] David Cameron, James Casey, Leanne Guy, Peter Kunszt, Sophie Lemaitre, Gavin McCance, Heinz Stockinger, Kurt Stockinger, Giuseppe Andronico, William Bell, Itzhak Ben-Akiva, Diana Bosio, Radovan Chytracek, Andrea Domenici, Flavia Donno, Wolfgang Hoschek, Erwin Laure, Levi Lucio, Paul Millar, Livio Salconi, Ben Segal, and Mika Silander. Replica management in the European DataGrid Project. Journal of Grid Computing, 2(4):341–351, December 2004. CO-DEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=1570-7873&
                                        volume=2&issue=4&spage=341.

Avellino:2004:DWM

[59] G. Avellino, S. Beco, B. Cantalupo, A. Maraschini, F. Pacini, M. Sottilaro, A. Terracina, D. Colling, F. Giacomini, E. Ronchieri, A. Gianelle, M. Mazzucato, R. Peluso, M. Sgaravatto, A. Guarise, R. Piro, A. Werbrouck, D. Kouřil, A. Křenek, L. Matyska, M. Mulač, J. Pospišil, M. Ruda, Z. Salvet, J. Sitera, J. Škrabal, M. Vocuring, M. Mezzadri, F. Prelz, S. Monforte, and M. Pappalardo. The DataGrid Workload Management System: Challenges and results. Journal of Grid Computing, 2(4):353–367, December 2004. CO-DEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1570-7873&
volume=2&issue=4&spage=353.

Burke:2004:HAT

[60] S. Burke, F. Harris, I. Stokes-Rees, I. Augustin, F. Carminati, J. Closier, E. van Herwijnen, A. Sciaba, D. Boutigny, J. J. Blaising, V. Garonne, A. Tsaregorodtsev, P. Capiluppi, A. Fanfani, C. Grandi, R. Barbera, E. Luppi, G. Negri, L. Perini, S. Resconi, M. Reale, A. De Salvo, S. Bagnasco, P. Cerello, K. Bos, D. Groep, W. van Leeuwen, J. Templon, O. Smirnova, O. Maroney, F. Brochu, and D. Colling. HEP applications and their experience with the use of DataGrid middleware. Journal of Grid Computing, 2(4):369–386, December 2004. CO-DEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1570-7873&
volume=2&issue=4&spage=369.

Montagnat:2004:MIS

[61] J. Montagnat, F. Bellet, H. Benoit-Cattin, V. Breton, L. Brunie, H. Duque, Y. Legré, I. E. Magnin, L. Maigne, S. Miguet, J.-M. Pierson, L. Seitz, and T. Tweed. Medical images simulation, storage, and processing on the European DataGrid Testbed. Journal of Grid Computing, 2(4):387–400, December 2004. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic).

URL http://www.springerlink.com/
openurl.asp?genre=article&issn=
1570-7873&volume=2&issue=4&spage=
387.

Truong:2005:DIP

[62] Hong-Linh Truong, Thomas Fahringer, and Schahram Dustdar. Dynamic instrumentation, performance monitoring and analysis of grid scientific workflows. Journal of Grid Computing, 3(1–2):1– 18, June 2005. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic).

                                        URL http://www.springerlink.com/
                                        openurl.asp?genre=article&issn=
                                        1570-7873&volume=3&issue=1&spage=
                                        1.

Adzigogov:2005:EOG

[63] Lazar Adzigogov, John Soldatos, and Lazaros Polymenakos. EMPEROR: An OGSA grid meta-scheduler based on dynamic resource predictions. Journal of Grid Computing, 3(1–2):19–37, June 2005. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic).

                                        URL http://www.springerlink.com/
                                        openurl.asp?genre=article&issn=
                                        1570-7873&volume=3&issue=1&spage=
                                        19.

James:2005:SDM

[64] H. A. James and K. A. Hawick. Scientific data management in a grid environment. Journal of Grid Computing, 3(1–2):39– 51, June 2005. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic).

                                        URL http://www.springerlink.com/
                                        openurl.asp?genre=article&issn=
                                        1570-7873&volume=3&issue=1&spage=
                                        39.

Song:2005:TGC

[65] Shanshan Song, Kai Hwang, and Yu-Kwong Kwok. Trusted grid computing with security binding and trust integration. Journal of Grid Computing, 3(1–2):53–73, June 2005. CO-DEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1570-7873&
volume=3&issue=1&spage=53.

Delaitre:2005:GRL

[66] Thierry Delaitre, Tamas Kiss, Ariel Goyeneche, Gabor Terstyanszky, Stephen Winter, and Peter Kacsuk. GEMLCA: Running legacy code applications as grid services. Journal of Grid Computing, 3(1–2):75–90, June 2005. CO-DEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1570-7873&
volume=3&issue=1&spage=75.

Galstyan:2005:RAG

[67] Aram Galstyan, Karl Czajkowski, and Kristina Lerman. Resource allocation in the grid with learning agents. Journal of Grid Computing, 3(1–2):91–100, June 2005. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic).

URL http://www.springerlink.com/
openurl.asp?genre=article&issn=
1570-7873&volume=3&issue=1&spage=
91.

Jensen:2005:EGA

[68] Jens G. Jensen, Tara Shah, Owen Synge, John Gordon, Glen Johnson, and Regina Tam. Enabling grid access to mass storage. Journal of Grid Computing, 3(1–2):101–112, June 2005. CO-DEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1570-7873&
volume=3&issue=1&spage=101.

Perry:2005:QGR

[69] J. Perry, L. Smith, A. N. Jackson, R. D. Kenway, B. Joo, C. M. Maynard, A. Trew, D. Byrne, G. Beckett, C. T. H. Davies, S. Downing, A. C. Irving, C. McNeile, Z. Sroczynski, C. R. Allton, W. Armour, and J. M. Flynn. QCDgrid: a grid resource for quantum chromodynamics. Journal of Grid Computing, 3(1–2):113–130, June 2005. CO-DEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=1570-7873&
                                        volume=3&issue=1&spage=113.

Dinda:2005:FCQ

[70] Peter Dinda and Dong Lu. Fast compositional queries in a relational grid information service. Journal of Grid Computing, 3(1–2):131–150, June 2005. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=1570-7873&
                                        volume=3&issue=1&spage=131.

Deelman:2005:P

[71] Ewa Deelman and Ian Taylor. Preface. Journal of Grid Computing, 3(3–4):151, September 2005. CO-DEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=1570-7873&
                                        volume=3&issue=3&spage=151.

Taylor:2005:VGW

[72] Ian Taylor, Matthew Shields, Ian Wang, and Andrew Harrison. Visual grid workflow in Triana. Journal of Grid Computing, 3(3–4):153–169, September 2005.

CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1570-7873&
volume=3&issue=3&spage=153.

Yu:2005:TWM

[73] Jia Yu and Rajkumar Buyya. A taxonomy of workflow management systems for grid computing. Journal of Grid Computing, 3(3–4):171–200, September 2005. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic).

URL http://www.springerlink.com/
openurl.asp?genre=article&issn=
1570-7873&volume=3&issue=3&spage=
171.

Singh:2005:OGB

[74] Gurmeet Singh, Carl Kesselman, and Ewa Deelman. Optimizing grid-based workflow execution. Journal of Grid Computing, 3(3–4):201–219, September 2005. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic).

URL http://www.springerlink.com/
openurl.asp?genre=article&issn=
1570-7873&volume=3&issue=3&spage=
201.

Kacsuk:2005:MGM

[75] Péter Kacsuk and Gergely Sipos. Multigrid, multi-user workflows in the P-GRADE Grid Portal. Journal of Grid Computing, 3(3–4):221–238, September 2005. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic).

URL http://www.springerlink.com/
openurl.asp?genre=article&issn=
1570-7873&volume=3&issue=3&spage=

vonLaszewski:2005:WCJ

[76] Gregor von Laszewski and Mike Hategan. Workflow concepts of the Java CoG Kit. Journal of Grid Computing, 3(3–4):239–258, September 2005. CO-DEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=1570-7873&
                                        volume=3&issue=3&spage=239.

McGough:2005:EEW

[77] A. Stephen McGough, Jeremy Cohen, John Darlington, Eleftheria Katsiri, William Lee, Sofia Panagiotidi, and Yash Patel. An end-to-end workflow pipeline for large-scale grid computing. Journal of Grid Computing, 3 (3–4):259–281, September 2005. CO-DEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=1570-7873&
                                        volume=3&issue=3&spage=259.

Emmerich:2005:GSO

[78] Wolfgang Emmerich, Ben Butchart, Liang Chen, Bruno Wassermann, and Sarah L. Price. Grid service orchestration using the Business Process Execution Language (BPEL). Journal of Grid Computing, 3(3–4):283–304, September 2005. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic).

                                        URL http://www.springerlink.com/
                                        openurl.asp?genre=article&issn=
                                        1570-7873&volume=3&issue=3&spage=
                                        283.

Lucchese:2006:ASG

[79] Fabiano de O. Lucchese, Eduardo J. Huerta Yero, Francisco S. Sambatti, and Marco A. A. Henriques. An adaptive scheduler for Grids. Journal of Grid Computing, 4(1):1–17, March 2006. CO-DEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1570-7873&
volume=4&issue=1&spage=1.

Desprez:2006:SSR

[80] Frédéric Desprez and Antoine Vernois. Simultaneous scheduling of replication and computation for data-intensive applications on the grid. Journal of Grid Computing, 4(1):19–31, March 2006. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1570-7873&
volume=4&issue=1&spage=19.

Ludwig:2006:MFM

[81] Simone A. Ludwig, Omer F. Rana, Julian Padget, and William Naylor. Matchmaking framework for mathematical Web services. Journal of Grid Computing, 4(1):33–48, March 2006. CO-DEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1570-7873&
volume=4&issue=1&spage=33.

Chervenak:2006:APP

[82] Ann L. Chervenak and Min Cai. Applying peer-to-peer techniques to grid replica location services. Journal of Grid Computing, 4(1):49–69, March 2006. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1570-7873&
volume=4&issue=1&spage=49.

Huang:2006:PMP

[83] Sili Huang, Eric Aubanel, and Virendrakumar C. Bhavsar. PaGrid: a mesh partitioner for computational Grids. Journal of Grid Computing, 4(1):71–88, March 2006. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic).

                                        URL http://www.springerlink.com/
                                        openurl.asp?genre=article&issn=
                                        1570-7873&volume=4&issue=1&spage=
                                        71.

Prem:2006:SVM

[84] Hema Prem and N. R. Srinivasa Raghavan. A support vector machine based approach for forecasting of network weather services. Journal of Grid Computing, 4(1):89–114, March 2006. CO-DEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=1570-7873&
                                        volume=4&issue=1&spage=89.

Brebner:2006:TWG

[85] Paul Brebner and Wolfgang Emmerich. Two ways to grid: The contribution of Open Grid Services Architecture (OGSA) mechanisms to service-centric and resource-centric lifecycles. Journal of Grid Computing, 4(1):115–131, March 2006. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic).

                                        URL http://www.springerlink.com/
                                        openurl.asp?genre=article&issn=
                                        1570-7873&volume=4&issue=1&spage=
                                        115.

Kielmann:2006:GAE

[86] Thilo Kielmann. Grid applications: From early adopters to mainstream users. Journal of Grid Computing, 4(2):133–134, June 2006. CO-DEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1570-7873&
volume=4&issue=2&spage=133.

Foster:2006:SGO

[87] Ian Foster, Veronika Nefedova, Mehran Ahsant, Rachana Ananthakrishnan, Lee Liming, Ravi Madduri, Olle Mulmo, Laura Pearlman, and Frank Siebenlist. Streamlining grid operations: Definition and deployment of a portal-based user registration service. Journal of Grid Computing, 4(2):135–144, June 2006. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1570-7873&
volume=4&issue=2&spage=135.

Tanimura:2006:IFT

[88] Yusuke Tanimura, Tsutomu Ikegami, Hidemoto Nakada, Yoshio Tanaka, and Satoshi Sekiguchi. Implementation of fault-tolerant GridRPC applications. Journal of Grid Computing, 4(2):145–157, June 2006. CO-DEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1570-7873&
volume=4&issue=2&spage=145.

Yan:2006:CGM

[89] Yonghong Yan and Barbara M. Chapman. Campus Grids meet applications: Modeling, metascheduling and integration. Journal of Grid Computing, 4(2):159–175, June 2006. CO-DEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL http:

//www.springerlink.com/openurl.

                                        asp?genre=article&issn=1570-7873&
                                        volume=4&issue=2&spage=159.

Chrabakh:2006:GDI

[90] Wahid Chrabakh and Rich Wolski. Grid-SAT: Design and implementation of a computational grid application. Journal of Grid Computing, 4(2):177–193, June 2006. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic).

                                        URL http://www.springerlink.com/
                                        openurl.asp?genre=article&issn=
                                        1570-7873&volume=4&issue=2&spage=
                                        177.

Dooley:2006:PPL

[91] Rion Dooley, Kent Milfeld, Chona Guiang, Sudhakar Pamidighantam, and Gabrielle Allen. From proposal to production: Lessons learned developing the computational chemistry grid cyberinfrastructure. Journal of Grid Computing, 4(2):195–208, June 2006. CO-DEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=1570-7873&
                                        volume=4&issue=2&spage=195.

Jithesh:2006:GAI

[92] P. V. Jithesh, P. Donachy, T. Harmer, N. Kelly, R. Perrott, S. Wasnik, J. Johnston, M. McCurley, M. Townsley, and S. McKee. GeneGrid: Architecture, implementation and application. Journal of Grid Computing, 4(2):209–222, June 2006. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic).

                                        URL http://www.springerlink.com/
                                        openurl.asp?genre=article&issn=
                                        1570-7873&volume=4&issue=2&spage=
                                        209.

Cappello:2006:ESI

[93] Franck Cappello, Adriana Iamnitchi, and Mitsuhisa Sato. Editorial: Special issue on global and peer-to-peer computing. Journal of Grid Computing, 4(3):223–224, September 2006. CO-DEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1570-7873&
volume=4&issue=3&spage=223.

Cirne:2006:LWU

[94] Walfredo Cirne, Francisco Brasileiro, Nazareno Andrade, Lauro B. Costa, Alisson Andrade, Reynaldo Novaes, and Miranda Mowbray. Labs of the world, unite!!! Journal of Grid Computing, 4(3):225–246, September 2006. CO-DEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1570-7873&
volume=4&issue=3&spage=225.

Azzedin:2006:TBU

[95] Farag Azzedin, Muthucumaru Maheswaran, and Arindam Mitra. Trust brokering and its use for resource matchmaking in public-resource Grids. Journal of Grid Computing, 4(3): 247–263, September 2006. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1570-7873&
volume=4&issue=3&spage=247.

Yang:2006:FST

[96] Shuo Yang, Ali R. Butt, Xing Fang, Y. Charlie Hu, and Samuel P. Midkiff. A fair, secure and trustworthy peer-to-peer based cycle-sharing system. Journal of Grid Computing, 4(3):265–286, September 2006. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic).

URL http://www.springerlink.com/
openurl.asp?genre=article&issn=
1570-7873&volume=4&issue=3&spage=
265.

Raicu:2006:DPU

[97] Ioan Raicu, Catalin Dumitrescu, Matei Ripeanu, and Ian Foster. The design, performance, and use of DiPerF: An automated DIstributed PERformance evaluation Framework. Journal of Grid Computing, 4(3):287–309, September 2006. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic).

URL http://www.springerlink.com/
openurl.asp?genre=article&issn=
1570-7873&volume=4&issue=3&spage=
287.

Randriamaro:2006:DDP

[98] Cyril Randriamaro, Olivier Soyez, Gil Utard, and Francis Wlazinski. Data distribution in a peer to peer storage system. Journal of Grid Computing, 4(3):311–321, September 2006. CO-DEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1570-7873&
volume=4&issue=3&spage=311.

Butt:2006:KPP

[99] Ali R. Butt, Troy A. Johnson, Yili Zheng, and Y. Charlie Hu. Kosha: a peer-to-peer enhancement for the Network File System. Journal of Grid Computing, 4(3):323–341, September 2006. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL http:

//www.springerlink.com/openurl.

asp?genre=article&issn=1570-7873&
volume=4&issue=3&spage=323.

Casalicchio:2006:NAA

[100] Emiliano Casalicchio, Federico Morabito, Giovanni Cortese, and Fabrizio Davide. A novel approach to adaptive content-based subscription management in DHT-based overlay networks. Journal of Grid Computing, 4(3):343–353, September 2006. CO-DEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1570-7873&
volume=4&issue=3&spage=343.

Li:2006:DCI

[101] Zhen Li and Manish Parashar. A decentralized computational infrastructure for grid-based parallel asynchronous iterative applications. Journal of Grid Computing, 4(4):355–372, December 2006. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1570-7873&
volume=4&issue=4&spage=355.

Derbal:2006:EGS

[102] Youcef Derbal. Entropic grid scheduling. Journal of Grid Computing, 4 (4):373–394, December 2006. CO-DEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1570-7873&
volume=4&issue=4&spage=373.

Aversa:2006:MMA

[103] Rocco Aversa, Beniamino Di Martino, Nicola Mazzocca, and Salvatore Venticinque. MAGDA: a mobile agent based grid architecture. Journal of Grid Computing, 4(4):395–412, December 2006. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1570-7873&
volume=4&issue=4&spage=395.

Burruss:2006:RAM

[104] J. R. Burruss, T. W. Fredian, and M. R. Thompson. ROAM: An authorization manager for Grids. Journal of Grid Computing, 4(4):413–423, December 2006. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic).

URL http://www.springerlink.com/
openurl.asp?genre=article&issn=
1570-7873&volume=4&issue=4&spage=
413.

Harrison:2006:WSR

[105] Andrew Harrison and Ian Taylor. The Web services resource framework in a peer-to-peer context. Journal of Grid Computing, 4(4):425–445, December 2006. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic).

URL http://www.springerlink.com/
openurl.asp?genre=article&issn=
1570-7873&volume=4&issue=4&spage=
425.

Miles:2007:RUP

[106] Simon Miles, Paul Groth, Miguel Branco, and Luc Moreau. The requirements of using provenance in e-science experiments. Journal of Grid Computing, 5(1):1–25, March 2007. CO-DEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1570-7873&
volume=5&issue=1&spage=1.

Genaud:2007:PMP

[107] Stéphane Genaud and Choopan Rattanapoka. P2P–MPI: a peer-to-peer framework for robust execution of message passing parallel programs on Grids. Journal of Grid Computing, 5(1):27–42, March 2007. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic).

URL http://www.springerlink.com/
openurl.asp?genre=article&issn=
1570-7873&volume=5&issue=1&spage=
27.

McClatchey:2007:DIN

[108] Richard McClatchey, Ashiq Anjum, Heinz Stockinger, Arshad Ali, Ian Willers, and Michael Thomas. Data Intensive and Network Aware (DIANA) Grid scheduling. Journal of Grid Computing, 5(1):43–64, March 2007. CO-DEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1570-7873&
volume=5&issue=1&spage=43.

Pereira:2007:MRB

[109] Anil L. Pereira, Vineela Muppavarapu, and Soon M. Chung. Managing rolebased access control policies for grid databases in OGSA–DAI using CAS. Journal of Grid Computing, 5(1):65–81, March 2007. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic).

URL http://www.springerlink.com/
openurl.asp?genre=article&issn=
1570-7873&volume=5&issue=1&spage=
65.

Quetier:2007:SCF

[110] Benjamin Quétier, Vincent Neri, and Franck Cappello. Scalability comparison of four host virtualization tools.

Journal of Grid Computing, 5(1):83–98, March 2007. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic).

URL http://www.springerlink.com/
openurl.asp?genre=article&issn=
1570-7873&volume=5&issue=1&spage=
83.

Dumitrescu:2007:DUP

[111] Catalin L. Dumitrescu, Ioan Raicu, and Ian Foster. The design, usage, and performance of GRUBER: a Grid Usage Service Level Agreement based BrokERing infrastructure. Journal of Grid Computing, 5(1):99–126, March 2007. CO-DEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1570-7873&
volume=5&issue=1&spage=99.

Wolski:2007:SIF

[112] Rich Wolski and Henri Bal. Special issue featuring selected papers from HPDC-15. Journal of Grid Computing, 5(2):127–128, June 2007. CO-DEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1570-7873&
volume=5&issue=2&spage=127.

Cai:2007:IPP

[113] Zhongtang Cai, Vibhore Kumar, and Karsten Schwan. IQ-Paths: Predictably high performance data streams across dynamic network overlays. Journal of Grid Computing, 5(2):129–150, June 2007. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic).

URL http://www.springerlink.com/
openurl.asp?genre=article&issn=
1570-7873&volume=5&issue=2&spage=
129.

Ganguly:2007:WSO

[114] A. Ganguly, A. Agrawal, P. O. Boykin, and R. J. Figueiredo. WOW: Selforganizing wide area overlay networks of virtual workstations. Journal of Grid Computing, 5(2):151–172, June 2007. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1570-7873&
volume=5&issue=2&spage=151.

Ren:2007:PRA

[115] Xiaojuan Ren, Seyong Lee, Rudolf Eigenmann, and Saurabh Bagchi. Prediction of resource availability in finegrained cycle sharing systems empirical evaluation. Journal of Grid Computing, 5(2):173–195, June 2007. CO-DEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1570-7873&
volume=5&issue=2&spage=173.

Malewicz:2007:TPD

[116] Grzegorz Malewicz, Ian Foster, Arnold L.

Rosenberg, and Michael Wilde. A tool for prioritizing DAGMan jobs and its evaluation. Journal of Grid Computing, 5(2):197–212, June 2007. CO-DEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1570-7873&
volume=5&issue=2&spage=197.

Mills:2007:RPS

[117] Richard T. Mills, Chuan Yue, Andreas Stathopoulos, and Dimitrios S. Nikolopoulos. Runtime and programming support for memory adaptation in scientific applications via local disk and remote memory. Journal of Grid Computing, 5(2):213–234, June 2007. CO-DEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1570-7873&
volume=5&issue=2&spage=213.

Casanova:2007:BDR

[118] Henri Casanova. Benefits and drawbacks of redundant batch requests. Journal of Grid Computing, 5(2):235–250, June 2007. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic).

URL http://www.springerlink.com/
openurl.asp?genre=article&issn=
1570-7873&volume=5&issue=2&spage=
235.

Newhouse:2007:AEA

[119] T. Newhouse and J. Pasquale. Achieving efficiency and accuracy in the ALPS application-level proportionalshare scheduler. Journal of Grid Computing, 5(2):251–270, June 2007. CO-DEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1570-7873&
volume=5&issue=2&spage=251.

Pacitti:2007:PSI

[120] Esther Pacitti, Marta Mattoso, and Patrick Valduriez. Preface to the special issue on grid data management. Journal of Grid Computing, 5(3):271–272, September 2007. CO-DEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1570-7873&
volume=5&issue=3&spage=271.

Pacitti:2007:GDM

[121] Esther Pacitti, Patrick Valduriez, and Marta Mattoso. Grid data management: Open problems and new issues. Journal of Grid Computing, 5 (3):273–281, September 2007. CO-DEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1570-7873&
volume=5&issue=3&spage=273.

Jimenez-Peris:2007:EGC

[122] R. Jiménez-Peris, M. Patiño-Martínez, and B. Kemme. Enterprise Grids: Challenges ahead. Journal of Grid Computing, 5(3):283–294, September 2007. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1570-7873&
volume=5&issue=3&spage=283.

Faerman:2007:MLS

[123] Marcio Faerman, Reagan Moore, Yifeng Cui, Yuanfang Hu, Jing Zhu, Bernard Minster, and Philip Maechling. Managing large scale data for earthquake simulations. Journal of Grid Computing, 5(3):295–302, September 2007. CO-DEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1570-7873&
volume=5&issue=3&spage=295.

Akbarinia:2007:DMA

[124] Reza Akbarinia and Vidal Martins.

Data management in the APPA system. Journal of Grid Computing, 5 (3):303–317, September 2007. CO-DEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=1570-7873&
                                        volume=5&issue=3&spage=303.

Merz:2007:FTR

[125] Peter Merz and Katja Gorunova. Faulttolerant resource discovery in peer-to-peer Grids. Journal of Grid Computing, 5(3):319–335, September 2007. CO-DEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=1570-7873&
                                        volume=5&issue=3&spage=319.

Caron:2007:ECG

[126] Eddy Caron, Frédéric Desprez, and Cédric Tedeschi. Enhancing computational Grids with peer-to-peer technology for large scale service discovery. Journal of Grid Computing, 5 (3):337–360, September 2007. CO-DEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=1570-7873&
                                        volume=5&issue=3&spage=337.

Krivitski:2007:LFL

[127] Denis Krivitski, Assaf Schuster, and Ran Wolff. A local facility location algorithm for large-scale distributed systems. Journal of Grid Computing, 5(4):361–378, December 2007. CO-DEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=1570-7873&
                                        volume=5&issue=4&spage=361.

Kondo:2007:STP

[128] Derrick Kondo, Andrew A. Chien, and Henri Casanova. Scheduling task parallel applications for rapid turnaround on enterprise desktop Grids. Journal of Grid Computing, 5(4):379–405, December 2007. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic).

URL http://www.springerlink.com/
openurl.asp?genre=article&issn=
1570-7873&volume=5&issue=4&spage=
379.

Hastings:2007:IOS

[129] Shannon Hastings, Scott Oster, Stephen Langella, David Ervin, Tahsin Kurc, and Joel Saltz. Introduce: An open source toolkit for rapid development of strongly typed grid services. Journal of Grid Computing, 5(4):407–427, December 2007. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic).

URL http://www.springerlink.com/
openurl.asp?genre=article&issn=
1570-7873&volume=5&issue=4&spage=
407.

Middleton:2007:QSN

[130] S. E. Middleton, M. Surridge, S. Benkner, and G. Engelbrecht. Quality of service negotiation for commercial medical grid services. Journal of Grid Computing, 5(4):429–447, December 2007. CO-DEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1570-7873&
volume=5&issue=4&spage=429.

Lamanna:2008:P

[131] Massimo Lamanna and Erwin Laure.

Preface. Journal of Grid Computing, 6(1):1–2, March 2008. CO-DEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1570-7873&
volume=6&issue=1&spage=1.

Andreeva:2008:HEP

[132] Julia Andreeva, Simone Campana, Federica Fanzago, and Juha Herrala. Highenergy physics on the grid: the ATLAS and CMS experience. Journal of Grid Computing, 6(1):3–13, March 2008. CO-DEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1570-7873&
volume=6&issue=1&spage=3.

Germain-Renaud:2008:SRG

[133] Cécile Germain-Renaud, Charles Loomis, Jakub T. Mościcki, and Romain Texier. Scheduling for responsive Grids. Journal of Grid Computing, 6(1):15–27, March 2008. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic).

URL http://www.springerlink.com/
openurl.asp?genre=article&issn=
1570-7873&volume=6&issue=1&spage=
15.

Jacq:2008:GEV

[134] N. Jacq, J. Salzemann, F. Jacq, Y. Legré, E. Medernach, J. Montagnat, A. Maaß, M. Reichstadt, H. Schwichtenberg, M. Sridhar, V. Kasam, M. Zimmermann, M. Hofmann, and V. Breton. Grid-enabled virtual screening against malaria. Journal of Grid Computing, 6(1):29–43, March 2008. CO-DEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1570-7873&
volume=6&issue=1&spage=29.

Montagnat:2008:SGM

[135] Johan Montagnat, Ákos Frohner, Daniel Jouvenot, Christophe Pera, Peter Kunszt, Birger Koblitz, Nuno Santos, Charles Loomis, Romain Texier, Diane Lingrand, Patrick Guio, Ricardo Brito Da Rocha, Antonio Sobreira de Almeida, and Zoltán Farkas. A secure grid medical data manager interfaced to the gLite middleware. Journal of Grid Computing, 6(1):45–59, March 2008. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1570-7873&
volume=6&issue=1&spage=45.

Koblitz:2008:AMS

[136] B. Koblitz, N. Santos, and V. Pose.

The AMGA metadata service. Journal of Grid Computing, 6(1):61–76, March 2008. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic).

URL http://www.springerlink.com/
openurl.asp?genre=article&issn=
1570-7873&volume=6&issue=1&spage=
61.

Christodoulopoulos:2008:SAM

[137] Konstantinos Christodoulopoulos, Vasileios Gkamas, and Emmanouel A. Varvarigos. Statistical analysis and modeling of jobs in a grid environment. Journal of Grid Computing, 6(1):77–101, March 2008. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1570-7873&
volume=6&issue=1&spage=77.

Rahman:2008:RPS

[138] Rashedur M. Rahman, Ken Barker, and Reda Alhajj. Replica placement strategies in data grid. Journal of Grid Computing, 6(1):103–123, March 2008. CO-DEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1570-7873&
volume=6&issue=1&spage=103.

Forestiero:2008:BPP

[139] Agostino Forestiero, Carlo Mastroianni, and Giandomenico Spezzano. Building a peer-to-peer information system in Grids via self-organizing agents. Journal of Grid Computing, 6(2):125–140, June 2008. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic).

URL http://www.springerlink.com/
openurl.asp?genre=article&issn=
1570-7873&volume=6&issue=2&spage=
125.

Nakajima:2008:ICR

[140] Yoshihiro Nakajima, Mitsuhisa Sato, Yoshiaki Aida, Taisuke Boku, and Franck Cappello. Integrating computing resources on multiple grid-enabled job scheduling systems through a Grid RPC system. Journal of Grid Computing, 6(2):141–157, June 2008. CO-DEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1570-7873&
volume=6&issue=2&spage=141.

Baraglia:2008:LTS

[141] Ranieri Baraglia, Renato Ferrini, Nicola Tonellotto, Laura Ricci, and Ramin Yahyapour. A launch-time scheduling heuristics for parallel applications on wide area Grids. Journal of Grid Computing, 6(2):159–175, June 2008. CO-DEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1570-7873&
volume=6&issue=2&spage=159.

Wasson:2008:ROC

[142] Glenn Wasson, Norm Beekwilder, David Del Vecchio, Mark Morgan, and Marty Humphrey. Resource-oriented computing: Design, implementation, and evaluation of WSRF.NET. Journal of Grid Computing, 6(2):177–194, June 2008. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1570-7873&
volume=6&issue=2&spage=177.

Pugliese:2008:MSG

[143] Andrea Pugliese, Domenico Talia, and Ramin Yahyapour. Modeling and supporting grid scheduling. Journal of Grid Computing, 6(2):195–213, June 2008. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1570-7873&
volume=6&issue=2&spage=195.

Veit:2008:GEB

[144] Daniel J. Veit and Wolfgang Gentzsch.

Grid economics and business models. Journal of Grid Computing, 6 (3):215–217, September 2008. CO-DEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1570-7873&
volume=6&issue=3&spage=215.

Krishnan:2008:GES

[145] Ramayya Krishnan. Grid economics:

a selective discussion of two research problems. Journal of Grid Computing, 6(3):219–224, September 2008. CO-DEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1570-7873&
volume=6&issue=3&spage=219.

Eymann:2008:FTR

[146] Torsten Eymann, Stefan König, and Raimund Matros. A framework for trust and reputation in grid environments. Journal of Grid Computing, 6(3):225–237, September 2008. CO-DEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1570-7873&
volume=6&issue=3&spage=225.

Zheng:2008:DLB

[147] Qin Zheng, Chen-Khong Tham, and Bharadwaj Veeravalli. Dynamic load balancing and pricing in grid computing with communication delay. Journal of Grid Computing, 6(3):239–253, September 2008. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic).

URL http://www.springerlink.com/
openurl.asp?genre=article&issn=
1570-7873&volume=6&issue=3&spage=
239.

Broberg:2008:MOG

[148] James Broberg, Srikumar Venugopal, and Rajkumar Buyya. Marketoriented Grids and utility computing: The state-of-the-art and future directions. Journal of Grid Computing, 6 (3):255–276, September 2008. CO-DEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL http:

//www.springerlink.com/openurl.

asp?genre=article&issn=1570-7873&
volume=6&issue=3&spage=255.

Beck:2008:GED

[149] Roman Beck, Michael Schwind, and Oliver Hinz. Grid economics in departmentalized enterprises. Journal of Grid Computing, 6(3):277–290, September 2008. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic).

URL http://www.springerlink.com/
openurl.asp?genre=article&issn=
1570-7873&volume=6&issue=3&spage=
277.

Mills:2008:CEB

[150] Kevin L. Mills and Christopher Dabrowski. Can economics-based resource allocation prove effective in a computation marketplace? Journal of Grid Computing, 6(3):291–311, September 2008. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic).

URL http://www.springerlink.com/
openurl.asp?genre=article&issn=
1570-7873&volume=6&issue=3&spage=
291.

Spohrer:2008:SS

[151] Jim Spohrer, Laura C. Anderson, Norman J. Pass, Tryg Ager, and Daniel Gruhl. Service science. Journal of Grid Computing, 6(3):313–324, September 2008. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic).

URL http://www.springerlink.com/
openurl.asp?genre=article&issn=
1570-7873&volume=6&issue=3&spage=
313.

Neumann:2008:FCG

[152] Dirk Neumann, Jochen Stößer, Christof Weinhardt, and Jens Nimis. A framework for commercial Grids — economic and technical challenges. Journal of Grid Computing, 6(3):325–347, September 2008. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic).

URL http://www.springerlink.com/
openurl.asp?genre=article&issn=
1570-7873&volume=6&issue=3&spage=
325.

Streitberger:2008:SGM

[153] Werner Streitberger, Sebastian Hudert, Torsten Eymann, Bjoern Schnizler, Floriano Zini, et al. On the simulation of grid market coordination approaches. Journal of Grid Computing, 6(3):349–366, September 2008. CO-DEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1570-7873&
volume=6&issue=3&spage=349.

Gatani:2008:ARM

[154] Luca Gatani, Alessandra De Paola, Giuseppe Lo Re, and Salvatore Gaglio. An adaptive routing mechanism for P2P resource discovery. Journal of Grid Computing, 6(4):367, December 2008. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1570-7873&
volume=6&issue=4&spage=367.

Montagnat:2008:WBD

[155] Johan Montagnat, Tristan Glatard, Isabel Campos Plasencia, Francisco Castejón, Xavier Pennec, et al. Workflow-based data parallel applications on the EGEE production grid infrastructure. Journal of Grid Computing, 6(4):369–383, December 2008. CO-DEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1570-7873&
volume=6&issue=4&spage=369.

Opitz:2008:WDG

[156] Alek Opitz, Hartmut König, and Sebastian Szamlewska. What does grid computing cost? Journal of Grid Computing, 6(4):385–397, December 2008. CO-DEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1570-7873&
volume=6&issue=4&spage=385.

Kotani:2008:RSS

[157] Y. Kotani, F. Ino, and K. Hagihara. A resource selection system for cycle stealing in GPU Grids. Journal of Grid Computing, 6(4):399–416, December 2008. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1570-7873&
volume=6&issue=4&spage=399.

Yeo:2009:IRA

[158] Chee Shin Yeo and Rajkumar Buyya. Integrated risk analysis for a commercial computing service in utility computing. Journal of Grid Computing, 7(1):1–24, March 2009. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic).

URL http://www.springerlink.com/
openurl.asp?genre=article&issn=
1570-7873&volume=7&issue=1&spage=
1.

Tan:2009:RRM

[159] Jefferson Tan, David Abramson, and Colin Enticott. REMUS: a rerouting and multiplexing system for grid connectivity across firewalls. Journal of Grid Computing, 7(1):25–50, March 2009. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1570-7873&
volume=7&issue=1&spage=25.

Thain:2009:CPG

[160] Douglas Thain, Christopher Moretti, and Jeffrey Hemmes. Chirp: a practical global filesystem for cluster and grid computing. Journal of Grid Computing, 7(1):51–72, March 2009. CO-DEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1570-7873&
volume=7&issue=1&spage=51.

Byun:2009:DAS

[161] Eun-Kyu Byun and Jin-Soo Kim. DynaGrid: An adaptive, scalable, and reliable resource provisioning framework for WSRF-compliant applications. Journal of Grid Computing, 7(1):73–89, March 2009. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic).

URL http://www.springerlink.com/
openurl.asp?genre=article&issn=
1570-7873&volume=7&issue=1&spage=
73.

Al-Kiswany:2009:BMS

[162] Samer Al-Kiswany, Matei Ripeanu, Adriana Iamnitchi, and Sudharshan Vazhkudai. Beyond music sharing: An evaluation of peer-to-peer data dissemination techniques in large scientific collaborations. Journal of Grid Computing, 7(1):91–114, March 2009. CO-DEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL http:

//www.springerlink.com/openurl.

asp?genre=article&issn=1570-7873&
volume=7&issue=1&spage=91.

Domenici:2009:SDD

[163] Andrea Domenici and Flavia Donno.

Static and dynamic data models for the storage resource manager v2.2. Journal of Grid Computing, 7(1):115–133, March 2009. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic).

URL http://www.springerlink.com/
openurl.asp?genre=article&issn=
1570-7873&volume=7&issue=1&spage=
115.

Schulz:2009:CAI

[164] Sven Schulz, Wolfgang Blochinger, and Hannes Hannak. Capability-aware information aggregation in peer-to-peer Grids methods, architecture, and implementation. Journal of Grid Computing, 7(2):135–167, June 2009. CO-DEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1570-7873&
volume=7&issue=2&spage=135.

Lang:2009:FAB

[165] Bo Lang, Ian Foster, Frank Siebenlist, Rachana Ananthakrishnan, and Tim Freeman. A flexible attribute based access control method for grid computing. Journal of Grid Computing, 7(2):169–180, June 2009. CO-DEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1570-7873&
volume=7&issue=2&spage=169.

Waldburger:2009:EAM

[166] Martin Waldburger, Matthias Göhner, Helmut Reiser, Gabi Dreo Rodosek, and Burkhard Stiller. Evaluation of an accounting model for dynamic virtual organizations. Journal of Grid Computing, 7(2):181–204, June 2009. CO-DEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1570-7873&
volume=7&issue=2&spage=181.

Said:2009:MSM

[167] Mirza Pahlevi Said and Isao Kojima. S-MDS: Semantic monitoring and discovery system for the grid. Journal of Grid Computing, 7(2):205–224, June 2009. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1570-7873&
volume=7&issue=2&spage=205.

Hudert:2009:NSA

[168] Sebastian Hudert, Heiko Ludwig, and Guido Wirtz. Negotiating SLAs — an approach for a generic negotiation framework for WS-agreement. Journal of Grid Computing, 7(2):225–246, June 2009. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic).

URL http://www.springerlink.com/
openurl.asp?genre=article&issn=
1570-7873&volume=7&issue=2&spage=
225.

Andreozzi:2009:IPA

[169] S. Andreozzi, P. Ciancarini, D. Montesi, R. Moretti, and S. Pardi. Implementation and performance analysis of XMatch: a language for qualitybased selection of grid services. Journal of Grid Computing, 7(2):247–264, June 2009. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic).

URL http://www.springerlink.com/
openurl.asp?genre=article&issn=
1570-7873&volume=7&issue=2&spage=
247.

Muppavarapu:2009:RBA

[170] Vineela Muppavarapu and Soon M.

Chung. Role-based access control in a data grid using the storage resource broker and Shibboleth. Journal of Grid Computing, 7(2):265–283, June 2009. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1570-7873&
volume=7&issue=2&spage=265.

Riedel:2009:GIR

[171] Morris Riedel and Gabor Terstyanszky. Grid interoperability for eresearch. Journal of Grid Computing, 7(3):285–286, September 2009. CO-DEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1570-7873&
volume=7&issue=3&spage=285.

Field:2009:GDE

[172] Laurence Field, Erwin Laure, and Markus W. Schulz. Grid deployment experiences: Grid interoperation. Journal of Grid Computing, 7 (3):287–296, September 2009. CO-DEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1570-7873&
volume=7&issue=3&spage=287.

Garzoglio:2009:DIS

[173] Gabriele Garzoglio, Ian Alderman, Mine Altunay, Rachana Ananthakrishnan, Joe Bester, et al. Definition and implementation of a SAML-XACML profile for authorization interoperability across grid middleware in OSG and EGEE. Journal of Grid Computing, 7(3):297–307, September 2009. CO-DEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1570-7873&
volume=7&issue=3&spage=297.

Jensen:2009:PGS

[174] Jens Jensen, Roger Downing, Derek Ross, Matt Hodges, and Alex Sim. Practical grid storage interoperation: Interoperation of SRM and SRB now. Journal of Grid Computing, 7 (3):309–317, September 2009. CO-DEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1570-7873&
volume=7&issue=3&spage=309.

Baur:2009:IGI

[175] Timo Baur, Rebecca Breu, Tibor Kálmán, Tobias Lindinger, Anne Milbert, et al. An interoperable grid information system for integrated resource monitoring based on virtual organizations. Journal of Grid Computing, 7 (3):319–333, September 2009. CO-DEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1570-7873&
volume=7&issue=3&spage=319.

Urbah:2009:EBE

[176] Etienne Urbah, Peter Kacsuk, Zoltan Farkas, Gilles Fedak, Gabor Kecskemeti, et al. EDGeS: Bridging EGEE to BOINC and XtremWeb. Journal of Grid Computing, 7(3):335–354, September 2009. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic).

URL http://www.springerlink.com/
openurl.asp?genre=article&issn=
1570-7873&volume=7&issue=3&spage=
335.

Kiss:2009:AIG

[177] Tamas Kiss and Tamas Kukla. Achieving interoperation of grid data resources via workflow level integration. Journal of Grid Computing, 7 (3):355–374, September 2009. CO-DEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1570-7873&
volume=7&issue=3&spage=355.

Rings:2009:GCC

[178] Thomas Rings, Geoff Caryer, Julian Gallop, Jens Grabowski, Tatiana Kovacikova, et al. Grid and cloud computing: Opportunities for integration with the next generation network. Journal of Grid Computing, 7(3):375–393, September 2009. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic).

URL http://www.springerlink.com/
openurl.asp?genre=article&issn=
1570-7873&volume=7&issue=3&spage=
375.

Kanaris:2009:HEM

[179] Ioannis Kanaris, Vasileios Mylonakis, Aristotelis Chatziioannou, Ilias Maglogiannis, and John Soldatos. HEC-TOR: Enabling microarray experiments over the Hellenic Grid Infrastructure. Journal of Grid Computing, 7(3):395–416, September 2009. CO-DEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=1570-7873&
                                        volume=7&issue=3&spage=395.

Kondo:2009:PSI

[180] Derrick Kondo. Preface to the special issue on volunteer computing and desktop Grids. Journal of Grid Computing, 7(4):417–418, December 2009. CO-DEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=1570-7873&
                                        volume=7&issue=4&spage=417.

Ma:2009:IDA

[181] Xiaosong Ma, Sudharshan S. Vazhkudai, and Zhe Zhang. Improving data availability for better access performance: a study on caching scientific data on distributed desktop workstations. Journal of Grid Computing, 7(4):419–438, December 2009. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic).

                                        URL http://www.springerlink.com/
                                        openurl.asp?genre=article&issn=
                                        1570-7873&volume=7&issue=4&spage=
                                        419.

Kacsuk:2009:SDG

[182] Peter Kacsuk, Jozsef Kovacs, Zoltan Farkas, Attila Csaba Marosi, Gabor Gombas, et al. SZTAKI Desktop Grid (SZDG): a flexible and scalable desktop grid system. Journal of Grid Computing, 7(4):439–461, December 2009. CO-DEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=1570-7873&
                                        volume=7&issue=4&spage=439.

Bertis:2009:DGV

[183] Viktors Bertis, Raphaël Bolze, Frédéric Desprez, and Kevin Reed. From dedicated grid to volunteer grid: Large scale execution of a bioinformatics application. Journal of Grid Computing, 7(4):463–478, December 2009. CO-DEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1570-7873&
volume=7&issue=4&spage=463. See er-

ratum [190].

Rood:2009:GRA

[184] Brent Rood and Michael J. Lewis. Grid resource availability prediction-based scheduling and task replication. Journal of Grid Computing, 7(4):479–500, December 2009. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic).

URL http://www.springerlink.com/
openurl.asp?genre=article&issn=
1570-7873&volume=7&issue=4&spage=
479.

Heien:2009:CLL

[185] Eric Martin Heien, David P. Anderson, and Kenichi Hagihara. Computing low latency batches with unreliable workers in volunteer computing environments. Journal of Grid Computing, 7(4):501–518, December 2009. CO-DEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1570-7873&
volume=7&issue=4&spage=501.

Toth:2009:IPV

[186] David Toth and David Finkel. Improving the productivity of volunteer computing by using the most effective task retrieval policies. Journal of Grid Computing, 7(4):519–535, December 2009. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1570-7873&
volume=7&issue=4&spage=519.

Estrada:2009:PPA

[187] Trilce Estrada, Michela Taufer, and David P. Anderson. Performance prediction and analysis of BOINC projects: An empirical study with Em-BOINC. Journal of Grid Computing, 7(4):537–554, December 2009. CO-DEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1570-7873&
volume=7&issue=4&spage=537.

Silaghi:2009:DCN

[188] Gheorghe Cosmin Silaghi, Filipe Araujo, Luis Moura Silva, Patricio Domingues, and Alvaro E. Arenas. Defeating colluding nodes in desktop grid computing platforms. Journal of Grid Computing, 7(4):555–573, December 2009. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1570-7873&
volume=7&issue=4&spage=555.

Watanabe:2009:OSC

[189] Kan Watanabe, Masaru Fukushi, and Susumu Horiguchi. Optimal spotchecking for computation time minimization in volunteer computing. Journal of Grid Computing, 7(4): 575–600, December 2009. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1570-7873&
volume=7&issue=4&spage=575.

Berstis:2009:EDG

[190] Viktors Berstis, Raphaël Bolze, Frédéric Desprez, and Kevin Reed. Erratum to: From dedicated grid to volunteer grid: Large scale execution of a bioinformatics application. Journal of Grid Computing, 7(4):601, December 2009. CO-DEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1570-7873&
volume=7&issue=4&spage=601.  See

[183].

Murphy:2010:ACG

[191] Michael A. Murphy, Linton Abraham, Michael Fenn, and Sebastien Goasguen. Autonomic clouds on the grid. Journal of Grid Computing, 8(1):1–18, March 2010. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic).

URL http://www.springerlink.com/
openurl.asp?genre=article&issn=
1570-7873&volume=8&issue=1&spage=
1.

Andreetto:2010:SBJ

[192] Paolo Andreetto, Sergio Andreozzi, Antonia Ghiselli, Moreno Marzolla, Valerio Venturi, et al. Standards-based job management in grid systems. Journal of Grid Computing, 8(1):19–45, March 2010. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic).

URL http://www.springerlink.com/
openurl.asp?genre=article&issn=
1570-7873&volume=8&issue=1&spage=
19.

Villela:2010:MAC

[193] Daniel Villela. Minimizing the average completion time for concurrent grid applications. Journal of Grid Computing, 8(1):47–59, March 2010. CO-DEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=1570-7873&
                                        volume=8&issue=1&spage=47.

Georgatos:2010:GEC

[194] Fotis Georgatos, Vasileios Gkamas, Aristeidis Ilias, Giannis Kouretis, and Emmanouel Varvarigos. A grid-enabled CPU scavenging architecture and a case study of its use in the Greek school network. Journal of Grid Computing, 8(1):61–75, March 2010. CO-DEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=1570-7873&
                                        volume=8&issue=1&spage=61.

Schulz:2010:ONS

[195] Sven Schulz, Wolfgang Blochinger, and Mathias Poths. Orbweb — a network substrate for peer-to-peer desktop grid computing based on open standards. Journal of Grid Computing, 8(1):77–107, March 2010. CO-DEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=1570-7873&
                                        volume=8&issue=1&spage=77.

Sarkar:2010:AES

[196] Ajanta De Sarkar, Sarbani Roy, Dibyajyoti Ghosh, Rupam Mukhopadhyay, and Nandini Mukherjee. An adaptive execution scheme for achieving guaranteed performance in computational Grids. Journal of Grid Computing, 8(1):109–131, March 2010. CO-DEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1570-7873&
volume=8&issue=1&spage=109.

Yu:2010:ADL

[197] Chen Yu and Dan C. Marinescu. Algorithms for divisible load scheduling of data-intensive applications. Journal of Grid Computing, 8(1):133–155, March 2010. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic).

URL http://www.springerlink.com/
openurl.asp?genre=article&issn=
1570-7873&volume=8&issue=1&spage=
133.

Fanfani:2010:DAC

[198] Alessandra Fanfani, Anzar Afaq, Jose Afonso Sanches, Julia Andreeva, Giusepppe Bagliesi, et al. Distributed analysis in CMS. Journal of Grid Computing, 8(2):159–179, June 2010. CO-DEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1570-7873&
volume=8&issue=2&spage=159.

Lagouvardos:2010:GER

[199] Kostas Lagouvardos, Evangelos Floros, and Vassiliki Kotroni. A grid-enabled regional-scale ensemble forecasting system in the Mediterranean area. Journal of Grid Computing, 8(2):181–197, June 2010. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic).

URL http://www.springerlink.com/
openurl.asp?genre=article&issn=

                                        1570-7873&volume=8&issue=2&spage=
                                        181.

Begeman:2010:MGT

[200] K. G. Begeman, A. N. Belikov, D. R. Boxhoorn, F. Dijkstra, E. A. Valentijn, et al. Merging grid technologies Astro-WISE and EGEE. Journal of Grid Computing, 8(2):199–221, June 2010. CO-DEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=1570-7873&
                                        volume=8&issue=2&spage=199.

Taffoni:2010:GDB

[201] Giuliano Taffoni, Santi Cassisi, Patrizia Manzato, Marco Molinaro, Fabio Pasian, et al. Grid and databases: BaSTI as a practical integration example. Journal of Grid Computing, 8(2):223–240, June 2010. CO-DEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=1570-7873&
                                        volume=8&issue=2&spage=223.

Camarasu-Pop:2010:DPG

[202] Sorina Camarasu-Pop, Tristan Glatard, Jakub T. Mościcki, Hugues Benoit-Cattin, and David Sarrut. Dynamic partitioning of GATE Monte-Carlo simulations on EGEE. Journal of Grid Computing, 8(2):241–259, June 2010. CO-DEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=1570-7873&
                                        volume=8&issue=2&spage=241.

Berger:2010:PEP

[203] Maximilian Berger and Thomas Fahringer. Practical experience from porting and executing the Wien2k application on the EGEE production grid infrastructure. Journal of Grid Computing, 8(2):261–279, June 2010. CO-DEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1570-7873&
volume=8&issue=2&spage=261.

Pallis:2010:SSE

[204] George Pallis, Asterios Katsifodimos, and Marios D. Dikaiakos. Searching for software on the EGEE infrastructure. Journal of Grid Computing, 8(2):281–304, June 2010. CO-DEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1570-7873&
volume=8&issue=2&spage=281.

Lingrand:2010:OJS

[205] Diane Lingrand, Johan Montagnat, Janusz Martyniak, and David Colling. Optimization of jobs submission on the EGEE production grid: Modeling faults using workload. Journal of Grid Computing, 8(2):305–321, June 2010. CO-DEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1570-7873&
volume=8&issue=2&spage=305.

Andreeva:2010:EDM

[206] Julia Andreeva, Max Boehm, Benjamin Gaidioz, Edward Karavakis, Lukasz Kokoszkiewicz, et al. Experiment dashboard for monitoring computing activities of the LHC virtual organizations. Journal of Grid Computing, 8(2):323–339, June 2010. CO-DEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=1570-7873&
                                        volume=8&issue=2&spage=323.

Coghlan:2010:BET

[207] Brian Coghlan, John Walsh, Stephen Childs, Geoff Quigley, David O’Callaghan, et al. The back-end of a two-layer model for a federated national datastore for academic research VOs that integrates EGEE data management. Journal of Grid Computing, 8(2):341–364, June 2010. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic).

                                        URL http://www.springerlink.com/
                                        openurl.asp?genre=article&issn=
                                        1570-7873&volume=8&issue=2&spage=
                                        341.

Padmanabhan:2010:SOG

[208] Anand Padmanabhan, Sukumar Ghosh, and Shaowen Wang. A Self-Organized Grouping (SOG) framework for efficient grid resource discovery. Journal of Grid Computing, 8(3):365–389, September 2010. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic).

                                        URL http://www.springerlink.com/
                                        openurl.asp?genre=article&issn=
                                        1570-7873&volume=8&issue=3&spage=
                                        365.

Cunsolo:2010:GGS

[209] Vincenzo D. Cunsolo, Salvatore Distefano, Antonio Puliafito, and Marco L. Scarpa. GS^{3}: a grid storage system with security features. Journal of Grid Computing, 8(3):391–418, September 2010. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL http:

                                        //www.springerlink.com/openurl.

asp?genre=article&issn=1570-7873&
volume=8&issue=3&spage=391.

Bittencourt:2010:TSM

[210] Luiz Fernando Bittencourt and Edmundo R. M. Madeira. Towards the scheduling of multiple workflows on computational Grids. Journal of Grid Computing, 8(3):419–441, September 2010. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic).

URL http://www.springerlink.com/
openurl.asp?genre=article&issn=
1570-7873&volume=8&issue=3&spage=
419.

Schulz:2010:PSS

[211] Sven Schulz and Wolfgang Blochinger.

Parallel SAT solving on peer-to-peer desktop Grids. Journal of Grid Computing, 8(3):443–471, September 2010. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1570-7873&
volume=8&issue=3&spage=443.

Perez:2010:MOR

[212] Julien Perez, Cécile Germain-Renaud, Balazs Kégl, and Charles Loomis. Multiobjective reinforcement learning for responsive Grids. Journal of Grid Computing, 8(3):473–492, September 2010. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1570-7873&
volume=8&issue=3&spage=473.

Prodan:2010:NBS

[213] Radu Prodan and Marek Wieczorek.

Negotiation-based scheduling of scientific grid workflows through advance reservations. Journal of Grid Computing, 8(4):493–510, December 2010. CO-DEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1570-7873&
volume=8&issue=4&spage=493.

Calabria:2010:GBG

[214] Andrea Calabria, Davide Di Pasquale, Matteo Gnocchi, Paolo Alessandro Cozzi, Alessandro Orro, Gabriele Antonio Trombetti, and Luciano Milanesi. Grid based genome wide studies on atrial flutter. Journal of Grid Computing, 8(4):511–527, December 2010. CO-DEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1570-7873&
volume=8&issue=4&spage=511.

Chen:2010:GHT

[215] Hsin-Yen Chen, Mason Hsiung, Hurng-Chun Lee, Eric Yen, Simon C. Lin, et al. GVSS: a high throughput drug discovery service of avian flu and dengue fever for EGEE and EUAsiaGrid. Journal of Grid Computing, 8(4):529–541, December 2010. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic).

URL http://www.springerlink.com/
openurl.asp?genre=article&issn=
1570-7873&volume=8&issue=4&spage=
529.

Anglano:2010:SPP

[216] Cosimo Anglano, Massimo Canonico, and Marco Guazzone. The Share-Grid peer-to-peer desktop grid: Infrastructure, applications, and performance evaluation. Journal of Grid Computing, 8(4):543–570, December 2010. CO-DEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1570-7873&
volume=8&issue=4&spage=543.

Lagana:2010:CPT

[217] Antonio Laganà, Alessandro Costantini, Osvaldo Gervasi, Noelia Faginas Lago, Carlo Manuali, et al. COM-PCHEM: Progress towards GEMS a grid empowered molecular simulator and beyond. Journal of Grid Computing, 8(4):571–586, December 2010. CO-DEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1570-7873&
volume=8&issue=4&spage=571.

Kiss:2010:PSW

[218] Tamas Kiss, Pamela Greenwell, Hans Heindl, Gabor Terstyanszky, and Noam Weingarten. Parameter sweep workflows for modelling carbohydrate recognition. Journal of Grid Computing, 8(4):587–601, December 2010. CO-DEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1570-7873&
volume=8&issue=4&spage=587.

Foster:2011:EM

[219] Ian Foster and Peter Kacsuk. Editors’ message. Journal of Grid Computing, 9(1):1–2, March 2011. CO-DEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1570-7873&
volume=9&issue=1&spage=1.

Rimal:2011:ARC

[220] Bhaskar Prasad Rimal, Admela Jukan, Dimitrios Katsaros, and Yves Goeleven. Architectural requirements for cloud computing systems: An enterprise cloud approach. Journal of Grid Computing, 9(1):3–26, March 2011. CO-DEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=1570-7873&
                                        volume=9&issue=1&spage=3.

Huu:2011:JEC

[221] Tram Truong Huu, Guilherme Koslovski, Fabienne Anhalt, Johan Montagnat, and Pascale Vicat-Blanc Primet. Joint elastic cloud and virtual network framework for application performance-cost optimization. Journal of Grid Computing, 9(1):27–47, March 2011. CO-DEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=1570-7873&
                                        volume=9&issue=1&spage=27.

Caron:2011:PMB

[222] Eddy Caron, Frédéric Desprez, and Adrian Muresan. Pattern matching based forecast of non-periodic repetitive behavior for cloud clients. Journal of Grid Computing, 9(1):49–64, March 2011. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic).

                                        URL http://www.springerlink.com/
                                        openurl.asp?genre=article&issn=
                                        1570-7873&volume=9&issue=1&spage=
                                        49.

Diaz:2011:BDS

[223] Ricardo Graciani Diaz, Adria Casajus Ramo, Ana Carmona Agüero, Thomas Fifield, and Martin Sevior. Belle– DIRAC setup for using Amazon Elastic Compute Cloud providing homogeneous access to heterogeneous computing resources. Journal of Grid Computing, 9(1):65–79, March 2011. CO-DEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1570-7873&
volume=9&issue=1&spage=65.

Vanmechelen:2011:CFS

[224] Kurt Vanmechelen, Wim Depoorter, and Jan Broeckhove. Combining futures and spot markets: a hybrid market approach to economic grid resource management. Journal of Grid Computing, 9(1):81–94, March 2011. CO-DEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1570-7873&
volume=9&issue=1&spage=81.

Ramirez-Alcaraz:2011:JAS

[225] Juan Manuel Ramírez-Alcaraz, Andrei Tchernykh, Ramin Yahyapour, Uwe Schwiegelshohn, Ariel Quezada-Pina, José Luis González-García, and Adán Hirales-Carbajal. Job allocation strategies with user run time estimates for online scheduling in hierarchical grids. Journal of Grid Computing, 9(1):95–116, March 2011. CO-DEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1570-7873&
volume=9&issue=1&spage=95.

Prnjat:2011:GEM

[226] Ognjen Prnjat. Guest Editors’ message. Journal of Grid Computing, 9(2):117–118, June 2011. CO-DEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=1570-7873&
                                        volume=9&issue=2&spage=117.

Ferrari:2011:RSE

[227] Tiziana Ferrari and Luciano Gaido. Resources and services of the EGEE production infrastructure. Journal of Grid Computing, 9(2):119–133, June 2011. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=1570-7873&
                                        volume=9&issue=2&spage=119.

Balaz:2011:DGI

[228] Antun Balaz, Ognjen Prnjat, Dusan Vudragović, Vladimir Slavnić, Ioannis Liabotis, et al. Development of grid e-infrastructure in South-Eastern Europe. Journal of Grid Computing, 9(2):135–154, June 2011. CO-DEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=1570-7873&
                                        volume=9&issue=2&spage=135.

Andronico:2011:ISG

[229] Giuseppe Andronico, Valeria Ardizzone, Roberto Barbera, Bruce Becker, Riccardo Bruno, Antonio Calanducci, Diego Carvalho, Leandro Ciuffo, Marco Fargetta, Emidio Giorgio, Guiseppe La Rocca, Albert Masoni, Marco Paganoni, Federico Ruggieri, and Diego Scardaci. e-infrastructures for e-science: a global view. Journal of Grid Computing, 9(2):155–184, June 2011. CO-DEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1570-7873&
volume=9&issue=2&spage=155.

Lloyd:2011:IGI

[230] Ashley D. Lloyd and Terence M.

Sloan. Intercontinental grids: An infrastructure for demand-driven innovation. Journal of Grid Computing, 9(2):185–200, June 2011. CO-DEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1570-7873&
volume=9&issue=2&spage=185.

Altunay:2011:SDP

[231] Mine Altunay, Paul Avery, Kent Blackburn, Brian Bockelman, Michael Ernst, Dan Fraser, Robert Quick, Robert Gardner, Sebastien Goasguen, Tanya Levshina, Miron Livny, John McGee, Doug Olson, Ruth Pordes, Maxim Potekhin, Abhisek Rana, Alain Roy, Chander Sehgal, Igor Sfiligoi, Frank Wuerthwein, and The Open Science Grid Executive Board. A science driven production cyberinfrastructure — the Open Science Grid. Journal of Grid Computing, 9(2):201–218, June 2011. CO-DEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1570-7873&
volume=9&issue=2&spage=201.

Kacsuk:2011:TPE

[232] P. Kacsuk, J. Kovacs, Z. Farkas, A. Cs.

Marosi, and Z. Balaton. Towards a powerful European DCI based on desktop grids. Journal of Grid Computing, 9(2):219–239, June 2011. CO-DEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1570-7873&
volume=9&issue=2&spage=219.

Brasileiro:2011:USP

[233] Francisco Brasileiro, Matheus Gaudencio, Rafael Silva, Alexandre Duarte, Diego Carvalho, Diego Scardaci, Leandro Ciuffo, Rafael Mayo, Herbert Hoeger, Michael Stanton, Raul Ramos, Roberto Barbera, Bernal Marechal, and Philippe Gavillet. Using a simple prioritisation mechanism to effectively interoperate service and opportunistic grids in the EELA-2 e-infrastructure. Journal of Grid Computing, 9(2):241–257, June 2011. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic).

URL http://www.springerlink.com/
openurl.asp?genre=article&issn=
1570-7873&volume=9&issue=2&spage=
241.

Gentzsch:2011:DDE

[234] Wolfgang Gentzsch, Denis Girou, Alison Kennedy, Hermann Lederer, Johannes Reetz, Morris Riedel, Andreas Schott, Andrea Vanni, Mariano Vazquez, and Jules Wolfrat. DEISA — Distributed European Infrastructure for Supercomputing Applications. Journal of Grid Computing, 9(2):259–277, June 2011. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1570-7873&
volume=9&issue=2&spage=259.

Ludwig:2011:SIA

[235] Simone A. Ludwig and Azin Moallem. Swarm intelligence approaches for grid load balancing. Journal of Grid Computing, 9(3):279–301, September 2011. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1570-7873&
volume=9&issue=3&spage=279.

Shah:2011:EER

[236] Sayed Chhattan Shah and Myong Soon Park. An energy-efficient resource allocation scheme for mobile ad hoc computational grids. Journal of Grid Computing, 9(3):303–323, September 2011. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1570-7873&
volume=9&issue=3&spage=303.

Anand:2011:REM

[237] Rakhi Anand, Troy LeBlanc, Edgar Gabriel, and Jaspal Subhlok. A robust and efficient message passing library for volunteer computing environments. Journal of Grid Computing, 9(3):325–344, September 2011. CO-DEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1570-7873&
volume=9&issue=3&spage=325.

Munoz:2011:DDS

[238] Víctor Méndez Muñoz, Gabriel Amorós Vicente, and Mohammed Kaci. A decentralized deployment strategy and performance evaluation of LCG file catalog service. Journal of Grid Computing, 9(3):345–354, September 2011. CO-DEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL http:

//www.springerlink.com/openurl.

                                        asp?genre=article&issn=1570-7873&
                                        volume=9&issue=3&spage=345.

Birkenheuer:2011:IFT

[239] Georg Birkenheuer, André Brinkmann, Mikael Högqvist, Alexander Papaspyrou, Bernhard Schott, Dietmar Sommerfeld, and Wolfgang Ziegler. Infrastructure federation through virtualized delegation of resources and services DGSI: Adding interoperability to DCI meta schedulers. Journal of Grid Computing, 9(3):355–377, September 2011. CO-DEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=1570-7873&
                                        volume=9&issue=3&spage=355.

Sanjay:2011:SRT

[240] H. A. Sanjay and Sathish S. Vadhiyar. Strategies for rescheduling tightlycoupled parallel applications in multicluster grids. Journal of Grid Computing, 9(3):379–403, September 2011. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=1570-7873&
                                        volume=9&issue=3&spage=379.

Thierion:2011:GTR

[241] Vincent Thierion, Pierre-Alain Ayral, Geisel Jacob, Sauvagnargues-Lesage Sophie, and Payrastre Olivier. Grid technology reliability for flash flood forecasting: End-user assessment. Journal of Grid Computing, 9(3):405–422, September 2011. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic).

                                        URL http://www.springerlink.com/
                                        openurl.asp?genre=article&issn=

1570-7873&volume=9&issue=3&spage=
405.

Elteto:2011:TNS

[242] Tamás Éltetö, Cécile Germain-Renaud, Pascal Bondon, and Michèle Sebag. Towards non-stationary grid models. Journal of Grid Computing, 9 (4):423–440, December 2011. CO-DEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1570-7873&
volume=9&issue=4&spage=423.

Murri:2011:GSS

[243] Riccardo Murri, Peter Z. Kunszt, Sergio Maffioletti, and Valery Tschopp. GridCertLib: a single sign-on solution for Grid Web applications and portals. Journal of Grid Computing, 9(4):441–453, December 2011. CO-DEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1570-7873&
volume=9&issue=4&spage=441.

Murugavel:2011:AEM

[244] Sivagama Sundari Murugavel, Sathish S.

Vadhiyar, and Ravi S. Nanjundiah. Adaptive executions of multi-physics coupled applications on batch Grids. Journal of Grid Computing, 9(4): 455–478, December 2011. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1570-7873&
volume=9&issue=4&spage=455.

Farkas:2011:GIB

[245] Zoltán Farkas. Grid interoperability based on a formal design. Journal of Grid Computing, 9(4):479–499, December 2011. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic).

URL http://www.springerlink.com/
openurl.asp?genre=article&issn=
1570-7873&volume=9&issue=4&spage=
479.

March:2011:ROD

[246] Verdi March and Yong Meng Teo. A read-only distributed hash table. Journal of Grid Computing, 9 (4):501–529, December 2011. CO-DEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1570-7873&
volume=9&issue=4&spage=501.

Prodan:2011:DAB

[247] Radu Prodan, Marek Wieczorek, and Hamid Mohammadi Fard. Double auction-based scheduling of scientific applications in distributed grid and cloud environments. Journal of Grid Computing, 9(4):531–548, December 2011. CO-DEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1570-7873&
volume=9&issue=4&spage=531.

Birje:2011:WWG

[248] Mahantesh N. Birje and Sunilkumar S. Manvi. WiGriMMA: a wireless grid monitoring model using agents. Journal of Grid Computing, 9(4):549–572, December 2011. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic).

URL http://www.springerlink.com/
openurl.asp?genre=article&issn=
1570-7873&volume=9&issue=4&spage=

Kaceniauskas:2011:VGV

[249] Arnas Kaceniauskas and Ruslan Pacevic. VizLitG: Grid visualization eservice enabling partial dataset transfer from storage elements of gLitebased grid infrastructure. Journal of Grid Computing, 9(4):573–589, December 2011. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic).

URL http://www.springerlink.com/
openurl.asp?genre=article&issn=
1570-7873&volume=9&issue=4&spage=
573.

Kosar:2012:GEI

[250] Tevfik Kosar and Ioan Raicu. Guest Editors’ introduction: Special issue on data-intensive computing in the clouds. Journal of Grid Computing, 10(1):1–4, March 2012. CO-DEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1570-7873&
volume=10&issue=1&spage=1.

Juve:2012:ECP

[251] Gideon Juve, Ewa Deelman, G. Bruce Berriman, Benjamin P. Berman, and Philip Maechling. An evaluation of the cost and performance of scientific workflows on Amazon EC2. Journal of Grid Computing, 10(1):5–21, March 2012. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1570-7873&
volume=10&issue=1&spage=5.

Bunch:2012:LRS

[252] Chris Bunch, Brian Drawert, Navraj Chohan, Chandra Krintz, Linda Petzold, and Khawaja Shams. Language and runtime support for automatic configuration and deployment of scientific computing software over cloud fabrics. Journal of Grid Computing, 10(1):23–46, March 2012. CO-DEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1570-7873&
volume=10&issue=1&spage=23.

Zhang:2012:IDC

[253] Yanfeng Zhang, Qixin Gao, Lixin Gao, and Cuirong Wang. iMapReduce: a distributed computing framework for iterative computation. Journal of Grid Computing, 10(1):47–68, March 2012. CO-DEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1570-7873&
volume=10&issue=1&spage=47.

Min:2012:VVM

[254] Changwoo Min, Inhyeok Kim, Taehyoung Kim, and Young Ik Eom. VMMB: Virtual machine memory balancing for unmodified operating systems. Journal of Grid Computing, 10(1):69–84, March 2012. CO-DEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1570-7873&
volume=10&issue=1&spage=69.

Loboz:2012:CRU

[255] Charles Loboz. Cloud resource usage — heavy tailed distributions invalidating traditional capacity planning models. Journal of Grid Computing, 10(1):85–108, March 2012. CO-DEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1570-7873&
volume=10&issue=1&spage=85.

Pirzadeh:2012:PER

[256] Pouria Pirzadeh, Junichi Tatemura, Oliver Po, and Hakan Hacigümüs. Performance evaluation of range queries in key value stores. Journal of Grid Computing, 10(1):109–132, March 2012. CO-DEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1570-7873&
volume=10&issue=1&spage=109.

Wittek:2012:DPG

[257] Peter Wittek and Sándor Darányi. Digital preservation in grids and clouds: a middleware approach. Journal of Grid Computing, 10(1):133–149, March 2012. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1570-7873&
volume=10&issue=1&spage=133.

Dinh:2012:CSE

[258] Tien Tuan Anh Dinh, Wang Wenqiang, and Anwitaman Datta. City on the sky: Extending XACML for flexible, secure data sharing on the cloud. Journal of Grid Computing, 10(1):151–172, March 2012. CO-DEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1570-7873&
volume=10&issue=1&spage=151.

Wang:2012:BCM

[259] Wei Wang and Guosun Zeng. Bayesian cognitive model in scheduling algorithm for data intensive computing. Journal of Grid Computing, 10(1):173–184, March 2012. CO-DEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=1570-7873&
                                        volume=10&issue=1&spage=173.

Nunez:2012:IFS

[260] Alberto Núñez, Jose L. Vázquez-Poletti, Agustin C. Caminero, Gabriel G. Castañé, Jesus Carretero, and Ignacio M. Llorente. iCanCloud: a flexible and scalable cloud infrastructure simulator. Journal of Grid Computing, 10(1):185–209, March 2012. CO-DEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=1570-7873&
                                        volume=10&issue=1&spage=185.

Saleh:2012:NGS

[261] Ahmed Ibrahim Saleh, Amany M. Sarhan, and Amr M. Hamed. A new grid scheduler with failure recovery and rescheduling mechanisms: Discussion and analysis. Journal of Grid Computing, 10(2):211–235, June 2012. CO-DEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=1570-7873&
                                        volume=10&issue=2&spage=211.

Kalochristianakis:2012:DLG

[262] Michael N. Kalochristianakis, Fotis Georgatos, Vasilis Gkamas, Giannis Kouretis, and Emmanouel Varvarigos. Deploying LiveWN grids in the Greek school network. Journal of Grid Computing, 10(2):237–248, June 2012. CO-DEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1570-7873&
volume=10&issue=2&spage=237.

Aron:2012:FQP

[263] Rajni Aron and Inderveer Chana. Formal QoS policy based grid resource provisioning framework. Journal of Grid Computing, 10(2):249–264, June 2012. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1570-7873&
volume=10&issue=2&spage=249.

Becciani:2012:CSD

[264] Ugo Becciani, Vincenzo Antonuccio-Delogu, Alessandro Costa, and Catia Petta. Cosmological simulations and data exploration: a testcase on the usage of grid infrastructure. Journal of Grid Computing, 10(2):265–277, June 2012. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1570-7873&
volume=10&issue=2&spage=265.

Yang:2012:MWM

[265] Hailong Yang, Zhongzhi Luan, Wenjun Li, and Depei Qian. MapReduce workload modeling with statistical approach. Journal of Grid Computing, 10(2):279–310, June 2012. CO-DEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL http:

//www.springerlink.com/openurl.

                                        asp?genre=article&issn=1570-7873&
                                        volume=10&issue=2&spage=279.

Gkoutioudi:2012:MCJ

[266] Kyriaki Z. Gkoutioudi and Helen D. Karatza. Multi-criteria job scheduling in grid using an accelerated genetic algorithm. Journal of Grid Computing, 10(2):311–323, June 2012. CO-DEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=1570-7873&
                                        volume=10&issue=2&spage=311.

Hirales-Carbajal:2012:MWS

[267] Adán Hirales-Carbajal, Andrei Tchernykh, Ramin Yahyapour, José Luis González-García, Thomas Röblitz, and Juan Manuel Ramírez-Alcaraz. Multiple workflow scheduling strategies with user run time estimates on a grid. Journal of Grid Computing, 10(2):325–346, June 2012. CO-DEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=1570-7873&
                                        volume=10&issue=2&spage=325.

Aubanel:2012:P

[268] Eric Aubanel, Virendra C. Bhavsar, and Michael Alex Frumkin. Preface. Journal of Grid Computing, 10 (3):347–348, September 2012. CO-DEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=article&issn=1570-7873&
                                        volume=10&issue=3&spage=347.

Montes:2012:RSH

[269] Jesús Montes, Alberto Sánchez, and María S. Pérez. Riding out the storm: How to deal with the complexity of grid and cloud management. Journal of Grid Computing, 10(3):349–366, September 2012. CO-DEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1570-7873&
volume=10&issue=3&spage=349.

Wu:2012:DWM

[270] Qishi Wu, Mengxia Zhu, Yi Gu, Patrick Brown, Xukang Lu, Wuyin Lin, and Yangang Liu. A distributed workflow management system with case study of real-life scientific applications on grids. Journal of Grid Computing, 10(3):367–393, September 2012. CO-DEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1570-7873&
volume=10&issue=3&spage=367.

Yildirim:2012:EED

[271] Esma Yildirim and Tevfik Kosar. Endto-end data-flow parallelism for throughput optimization in high-speed networks. Journal of Grid Computing, 10(3):395–418, September 2012. CO-DEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1570-7873&
volume=10&issue=3&spage=395.

Costa:2012:UBN

[272] Rostand Costa, Francisco Brasileiro, Guido Lemos Filho, and Dênio Sousa.

Using broadcast networks to create on-demand extremely large scale high-throughput computing infrastructures. Journal of Grid Computing, 10(3):419–445, September 2012. CO-DEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1570-7873&
volume=10&issue=3&spage=419.

Rodero:2012:EET

[273] Ivan Rodero, Hariharasudhan Viswanathan, Eun Kyung Lee, Marc Gamell, Dario Pompili, and Manish Parashar. Energyefficient thermal-aware autonomic management of virtualized HPC cloud infrastructure. Journal of Grid Computing, 10(3):447–473, September 2012. CO-DEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1570-7873&
volume=10&issue=3&spage=447.

Tomas:2012:IGR

[274] Luis Tomás, Blanca Caminero, Carmen Carrión, and Agustín C. Caminero. On the improvement of grid resource utilization: Preventive and reactive rescheduling approaches. Journal of Grid Computing, 10(3):475–499, September 2012. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1570-7873&
volume=10&issue=3&spage=475.

Buscemi:2012:GTA

[275] Maria Grazia Buscemi, Ugo Montanari, and Sonia Taneja. A gametheoretic analysis of grid job scheduling. Journal of Grid Computing, 10 (3):501–519, September 2012. CO-DEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1570-7873&
volume=10&issue=3&spage=501.

deOliveira:2012:PBA

[276] Daniel de Oliveira, Kary A. C. S. Ocaña, Fernanda Baião, and Marta Mattoso. A provenance-based adaptive scheduling heuristic for parallel scientific workflows in clouds. Journal of Grid Computing, 10(3):521–552, September 2012. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1570-7873&
volume=10&issue=3&spage=521.

Niehorster:2012:CAS

[277] Oliver Niehörster, André Brinkmann, Axel Keller, Christoph Kleineweber, Jens Krüger, and Jens Simon. Costaware and SLO-fulfilling software as a service. Journal of Grid Computing, 10(3):553–577, September 2012. CO-DEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=1570-7873&
volume=10&issue=3&spage=553.

Huang:2012:RRS

[278] Zhen Huang, Yisong Lin, and Yuxing Peng. Robust redundancy scheme for the repair process: Hierarchical codes in the bandwidth-limited systems. Journal of Grid Computing, 10(3):579–597, September 2012. CO-DEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL http:

//www.springerlink.com/openurl.

                                        asp?genre=article&issn=1570-7873&
                                        volume=10&issue=3&spage=579.

Kiss:2012:SGB

[279] Tamas Kiss. Science gateways for the broader take-up of distributed computing infrastructures. Journal of Grid Computing, 10(4):599– 600, December 2012. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s10723-012-9245-0;   http://
                                        link.springer.com/content/pdf/10.
                                        1007/s10723-012-9245-0.pdf.

Kacsuk:2012:WPG

[280] Peter Kacsuk, Zoltan Farkas, Miklos Kozlovszky, Gabor Hermann, Akos Balasko, Krisztian Karoczkai, and Istvan Marton. WS-PGRADE/gUSE generic DCI gateway framework for a large variety of user communities. Journal of Grid Computing, 10(4):601–630, December 2012. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL

                                        http://link.springer.com/article/
                                        10.1007/s10723-012-9240-5.

Dziubecki:2012:EDI

[281] Piotr Dziubecki, Piotr Grabowski, Michal Krysiński, Tomasz Kuczyński, Krzysztof Kurowski, and Dawid Szejnfeld. Easy development and integration of science gateways with Vine Toolkit. Journal of Grid Computing, 10(4):631–645, December 2012. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL

                                        http://link.springer.com/article/
                                        10.1007/s10723-012-9236-1.

Maddineni:2012:DAR

[282] Sharath Maddineni, Joohyun Kim, Yaakoub El-Khamra, and Shantenu Jha. Distributed Application Runtime Environment (DARE): a standardsbased middleware framework for science gateways. Journal of Grid Computing, 10(4):647–664, December 2012. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL

http://link.springer.com/article/
10.1007/s10723-012-9244-1.

Maestre:2012:AUS

[283] Cristina Maestre, J. Damiàn Segrelles Quilis, Erik Torres, Ignacio Blanquer, Rosana Medina, Vicente Hernández, and Luis Martí. Assessing the usability of a science gateway for medical knowledge bases with TREN-CADIS. Journal of Grid Computing, 10(4):665–688, December 2012. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL

http://link.springer.com/article/
10.1007/s10723-012-9243-2.

Ardizzone:2012:DSG

[284] V. Ardizzone, R. Barbera, A. Calanducci, M. Fargetta, E. Ingrà, I. Porro, G. La Rocca, S. Monforte, R. Ricceri, R. Rotondo, D. Scardaci, and A. Schenone. The DECIDE science gateway. Journal of Grid Computing, 10(4):689–707, December 2012. CO-DEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL http:

//link.springer.com/article/10.
1007/s10723-012-9242-3;   http://
link.springer.com/content/pdf/10.
1007/s10723-012-9242-3.pdf.

Wu:2012:CGP

[285] Jie Wu, René Siewert, Andreas Hoheisel, Jürgen Falkner, Oliver Strauß, Dinko Berberovic, and Dagmar Krefting. The Charité Grid Portal: Userfriendly and secure access to grid-based resources and services. Journal of Grid Computing, 10(4):709–724, December 2012. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL

http://link.springer.com/article/
10.1007/s10723-012-9234-3.

Shahand:2012:GEG

[286] Shayan Shahand, Mark Santcroos, Antoine H. C. van Kampen, and Sílvia Delgado Olabarriaga. A gridenabled gateway for biomedical data analysis. Journal of Grid Computing, 10(4):725–742, December 2012. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL

http://link.springer.com/article/
10.1007/s10723-012-9233-4.

Wassenaar:2012:WSB

[287] Tsjerk A. Wassenaar, Marc van Dijk, Nuno Loureiro-Ferreira, Gijs van der Schot, Sjoerd J. de Vries, Christophe Schmitz, Johan van der Zwan, Rolf Boelens, Andrea Giachetti, Lucio Ferella, Antonio Rosato, Ivano Bertini, Torsten Herrmann, Hendrik R. A. Jonker, Anurag Bagaria, and et al. WeNMR: Structural biology on the grid. Journal of Grid Computing, 10 (4):743–767, December 2012. CO-DEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL http:

//link.springer.com/article/10.
1007/s10723-012-9246-z;   http://
link.springer.com/content/pdf/10.
1007/s10723-012-9246-z.pdf.

Gesing:2012:SSI

[288] Sandra Gesing, Richard Grunzke, Jens Krüger, Georg Birkenheuer, Martin Wewior, Patrick Schäfer, Bernd Schuller, Johannes Schuster, Sonja Herres-Pawlis, Sebastian Breuers, Ákos Balaskó, Miklos Kozlovszky, Anna Szikszay Fabri, Lars Packschies, Peter Kacsuk, and et al. A single signon infrastructure for science gateways on a use case for structural bioinformatics. Journal of Grid Computing, 10(4):769–790, December 2012. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL

http://link.springer.com/article/
10.1007/s10723-012-9247-y.

Cuomo:2013:SBB

[289] Antonio Cuomo, Giuseppe Di Modica, Salvatore Distefano, Antonio Puliafito, Massimiliano Rak, Orazio Tomarchio, Salvatore Venticinque, and Umberto Villano. An SLA-based broker for cloud infrastructures. Journal of Grid Computing, 11(1):1–25, March 2013. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL

http://link.springer.com/article/
10.1007/s10723-012-9241-4.

Li:2013:OLD

[290] Keqin Li. Optimal load distribution for multiple heterogeneous blade servers in a cloud computing environment. Journal of Grid Computing, 11(1):27–46, March 2013. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL

http://link.springer.com/article/
10.1007/s10723-012-9239-y.

Tao:2013:DGW

[291] Yongcai Tao, Hai Jin, Song Wu, Xuanhua Shi, and Lei Shi. Dependable grid workflow scheduling based on resource availability. Journal of Grid Computing, 11(1):47–61, March 2013. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL

http://link.springer.com/article/
10.1007/s10723-012-9237-0.

Chohan:2013:CPD

[292] Navraj Chohan, Chris Bunch, Chandra Krintz, and Navyasri Canumalla. Cloud platform datastore support. Journal of Grid Computing, 11(1):63–81, March 2013. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL

http://link.springer.com/article/
10.1007/s10723-012-9238-z.

Margaritis:2013:NSA

[293] Giorgos Margaritis, Andromachi Hatzieleftheriou, and Stergios V. Anastasiadis. Nephele: Scalable access control for federated file services. Journal of Grid Computing, 11(1):83–102, March 2013. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL

http://link.springer.com/article/
10.1007/s10723-012-9217-4.

Tu:2013:DPP

[294] Manghui Tu, Hui Ma, Liangliang Xiao, I.-Ling Yen, Farokh Bastani, and Dianxiang Xu. Data placement in P2P data grids considering the availability, security, access performance and load balancing. Journal of Grid Computing, 11(1):103–127, March 2013. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL

http://link.springer.com/article/
10.1007/s10723-012-9232-5.

Ponciano:2013:AGS

[295] Lesandro Ponciano and Francisco Brasileiro. Assessing green strategies in peer-to-peer opportunistic grids. Journal of Grid Computing, 11(1): 129–148, March 2013. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10723-012-
9218-3.

Pataki:2013:STP

[296] Máté Pataki and Attila Csaba Marosi.

Searching for translated plagiarism with the help of desktop grids. Journal of Grid Computing, 11(1):149–166, March 2013. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL

http://link.springer.com/article/
10.1007/s10723-012-9224-5.

Hovestadt:2013:AOC

[297] Matthias Hovestadt, Odej Kao, Andreas Kliem, and Daniel Warneke. Adaptive online compression in clouds — making informed decisions in virtual machine environments. Journal of Grid Computing, 11(2):167–186, June 2013. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL

http://link.springer.com/article/
10.1007/s10723-013-9249-4.

Zarrabi:2013:LSF

[298] Amirreza Zarrabi, Khairulmizam Samsudin, and Wan Azizun Wan Adnan. Linux support for fast transparent general purpose checkpoint/ restart of multithreaded processes in loadable kernel module. Journal of Grid Computing, 11(2):187–210, June 2013. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL

                                        http://link.springer.com/article/
                                        10.1007/s10723-013-9248-5.

Cafaro:2013:PBM

[299] Massimo Cafaro, Maria Mirto, and Giovanni Aloisio. Preference-based matchmaking of grid resources with CP-Nets. Journal of Grid Computing, 11(2):211–237, June 2013. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL

                                        http://link.springer.com/article/
                                        10.1007/s10723-012-9235-2.

Astsatryan:2013:EUS

[300] Hrachya Astsatryan, Vladimir Sahakyan, Yuri Shoukouryan, Michel Daydé, Aurelie Hurault, Ronan Guivarch, Harutyun Terzyan, and Levon Hovhannisyan. On the easy use of scientific computing services for large scale linear algebra and parallel decision making with the P-grade portal. Journal of Grid Computing, 11(2):239–248, June 2013. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL

                                        http://link.springer.com/article/
                                        10.1007/s10723-013-9254-7.

Stout:2013:UKX

[301] Lance Stout, Matthew Walker, Jérôme Lauret, Sebastien Goasguen, and Michael A. Murphy. Using Kestrel and XMPP to support the STAR experiment in the cloud. Journal of Grid Computing, 11(2):249–264, June 2013. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL

                                        http://link.springer.com/article/
                                        10.1007/s10723-013-9253-8.

Alfieri:2013:HGT

[302] Roberto Alfieri, Silvia Arezzini, Alberto Ciampa, Roberto De Pietri, and Enrico Mazzoni. HPC on the grid: The theophys experience. Journal of Grid Computing, 11(2):265–280, June 2013. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL

http://link.springer.com/article/
10.1007/s10723-012-9223-6.

Shamsi:2013:DIC

[303] Jawwad Shamsi, Muhammad Ali Khojaye, and Mohammad Ali Qasmi. Dataintensive cloud computing: Requirements, expectations, challenges, and solutions. Journal of Grid Computing, 11(2):281–310, June 2013. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL

http://link.springer.com/article/
10.1007/s10723-013-9255-6.

Rodero:2013:EIA

[304] Ivan Rodero, David Villegas, Norman Bobroff, Yanbin Liu, Liana Fong, and S. Masoud Sadjadi. Enabling interoperability among grid metaschedulers. Journal of Grid Computing, 11(2):311–336, June 2013. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL

http://link.springer.com/article/
10.1007/s10723-013-9252-9.

Montagnat:2013:GEI

[305] Johan Montagnat and Ian J. Taylor.

Guest Editor’s introduction: Special issue on workflow. Journal of Grid Computing, 11(3):337–339, September 2013. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL http:

//link.springer.com/article/10.
1007/s10723-013-9270-7;   http://
link.springer.com/content/pdf/10.
1007/s10723-013-9270-7.pdf.

Wozniak:2013:JLS

[306] Justin M. Wozniak, Michael Wilde, and Daniel S. Katz. JETS: Language and system support for manyparallel-task workflows. Journal of Grid Computing, 11(3):341–360, September 2013. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL

                                        http://link.springer.com/article/
                                        10.1007/s10723-013-9259-2.

Gu:2013:DTO

[307] Yi Gu, Chase Qishi Wu, Xin Liu, and Dantong Yu. Distributed throughput optimization for large-scale scientific workflows under fault-tolerance constraint. Journal of Grid Computing, 11(3):361–379, September 2013. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL

                                        http://link.springer.com/article/
                                        10.1007/s10723-013-9266-3.

Vahi:2013:CSU

[308] Karan Vahi, Ian Harvey, Taghrid Samak, Daniel Gunter, Kieran Evans, David Rogers, Ian Taylor, Monte Goode, Fabio Silva, Eddie Al-Shakarchi, Gaurang Mehta, Ewa Deelman, and Andrew Jones. A case study into using common real-time workflow monitoring infrastructure for scientific workflows. Journal of Grid Computing, 11(3):381–406, September 2013. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL

                                        http://link.springer.com/article/
                                        10.1007/s10723-013-9265-4.

Emeakaroha:2013:MOB

[309] Vincent C. Emeakaroha, Michael Maurer, Patrick Stern, Paweł P. Labaj, Ivona Brandic, and David P. Kreil.

Managing and optimizing bioinformatics workflows for data analysis in clouds. Journal of Grid Computing, 11(3):407–428, September 2013. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL

http://link.springer.com/article/
10.1007/s10723-013-9260-9.

Plankensteiner:2013:FGI

[310] Kassian Plankensteiner, Radu Prodan, Matthias Janetschek, Thomas Fahringer, Johan Montagnat, David Rogers, Ian Harvey, Ian Taylor, Ákos Balaskó, and Péter Kacsuk. Finegrain interoperability of scientific workflows in distributed computing infrastructures. Journal of Grid Computing, 11(3):429–455, September 2013. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL

http://link.springer.com/article/
10.1007/s10723-013-9261-8.

Rogers:2013:BPA

[311] David Rogers, Ian Harvey, Tram Truong Huu, Kieran Evans, Tristan Glatard, Ibrahim Kallel, Ian Taylor, Johan Montagnat, Andrew Jones, and Andrew Harrison. Bundle and pool architecture for multi-language, robust, scalable workflow executions. Journal of Grid Computing, 11(3):457–480, September 2013. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL

http://link.springer.com/article/
10.1007/s10723-013-9267-2.

Blanc:2013:WH

[312] Anja Le Blanc, John Brooke, Donal Fellows, Marco Soldati, David Pérez-Suárez, Alessandro Marassi, and Andrej Santin. Workflows for heliophysics. Journal of Grid Computing, 11(3):481–503, September 2013. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL

                                        http://link.springer.com/article/
                                        10.1007/s10723-013-9256-5.

Korkhov:2013:EWI

[313] Vladimir Korkhov, Dagmar Krefting, Tamas Kukla, Gabor Z. Terstyanszky, Matthan W. A. Caan, and Silvia D. Olabarriaga. Exploring workflow interoperability for neuroimage analysis on the SHIWA Platform. Journal of Grid Computing, 11(3):505–522, September 2013. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL

                                        http://link.springer.com/article/
                                        10.1007/s10723-013-9262-7.

Nadarajan:2013:SPB

[314] Gayathri Nadarajan, Yun-Heh Chen-Burger, and Robert B. Fisher. Semantics and planning based workflow composition for video processing. Journal of Grid Computing, 11(3):523–551, September 2013. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL

                                        http://link.springer.com/article/
                                        10.1007/s10723-013-9263-6.

Sonntag:2013:MYG

[315] Mirko Sonntag and Dimka Karastoyanova. Model-as-you-go: An approach for an advanced infrastructure for scientific workflows. Journal of Grid Computing, 11(3):553–583, September 2013. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL

                                        http://link.springer.com/article/
                                        10.1007/s10723-013-9268-1.

Cerezo:2013:CAS

[316] Nadia Cerezo, Johan Montagnat, and Mireille Blay-Fornarino. Computerassisted scientific workflow design. Journal of Grid Computing, 11(3):585– 612, September 2013. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10723-013-
9264-5.

Exposito:2013:APA

[317] Roberto R. Expósito, Guillermo L.

Taboada, Sabela Ramos, Jorge González-Domínguez, Juan Touriño, and Ramón Doallo. Analysis of I/O performance on an Amazon EC2 cluster compute and high I/O platform. Journal of Grid Computing, 11(4):613–631, December 2013. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL

http://link.springer.com/article/
10.1007/s10723-013-9250-y.

Zheng:2013:BDC

[318] Wei Zheng and Rizos Sakellariou.

Budget-deadline constrained workflow planning for admission control. Journal of Grid Computing, 11(4):633– 651, December 2013. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10723-013-
9257-4.

Nesmachnow:2013:EAS

[319] Sergio Nesmachnow, Bernabé Dorronsoro, Johnatan E. Pecero, and Pascal Bouvry. Energy-aware scheduling on multicore heterogeneous grid computing systems. Journal of Grid Computing, 11(4):653–680, December 2013. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL

                                        http://link.springer.com/article/
                                        10.1007/s10723-013-9258-3.

Filippidis:2013:IHB

[320] Christos Filippidis, Yiannis Cotronis, and Christos Markou. IKAROS: An HTTP-Based distributed file system, for low consumption & low specification devices. Journal of Grid Computing, 11(4):681–698, December 2013. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL

                                        http://link.springer.com/article/
                                        10.1007/s10723-013-9271-6.

Kertesz:2013:EFC

[321] A. Kertesz, G. Kecskemeti, M. Oriol, P. Kotcauer, S. Acs, M. Rodríguez, O. Mercè, A. Cs. Marosi, J. Marco, and X. Franch. Enhancing federated cloud management with an integrated service monitoring approach. Journal of Grid Computing, 11(4):699–720, December 2013. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL

                                        http://link.springer.com/article/
                                        10.1007/s10723-013-9269-0.

Kashyap:2013:SDS

[322] Rekha Kashyap and Deo Prakash Vidyarthi. Security driven scheduling model for computational grid using NSGA–II. Journal of Grid Computing, 11(4):721–734, December 2013. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL

                                        http://link.springer.com/article/
                                        10.1007/s10723-013-9251-x.

Simon:2013:MWS

[323] Balazs Simon, Balazs Goldschmidt, and Karoly Kondorosi. A metamodel for the Web services standards. Journal of Grid Computing, 11(4):735–752, December 2013. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL

http://link.springer.com/article/
10.1007/s10723-013-9273-4.

Munoz:2013:RAC

[324] Víctor Méndez Muñoz, Adrian Casajús Ramo, Víctor Fernández Albor, Ricardo Graciani Diaz, and Gonzalo Merino Arévalo. Rafhyc: an architecture for constructing resilient services on federated hybrid clouds. Journal of Grid Computing, 11(4):753–770, December 2013. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL

http://link.springer.com/article/
10.1007/s10723-013-9279-y.

Sill:2014:GEI

[325] Alan Sill and Gabor Kecskemeti. Guest Editors’ introduction: Special issue on interoperability, federation frameworks and application programming interfaces for IaaS clouds. Journal of Grid Computing, 12(1):1–2, March 2014. CO-DEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL http:

//link.springer.com/article/10.
1007/s10723-014-9291-x;   http://
link.springer.com/content/pdf/10.
1007/s10723-014-9291-x.pdf.

Chadwick:2014:AFI

[326] David W. Chadwick, Kristy Siu, Craig Lee, Yann Fouillat, and Damien Germonville. Adding federated identity management to OpenStack. Journal of Grid Computing, 12(1):3–27, March 2014. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (elec-

tronic). URL http://link.springer.

                                        com/article/10.1007/s10723-013-
                                        9283-2;   http://link.springer.
                                        com/content/pdf/10.1007/s10723-
                                        013-9283-2.pdf.

Field:2014:ERD

[327] Laurence Field, Shiraz Memon, Iván Márton, and Gábor Szigeti. The EMI registry: Discovering services in a federated world. Journal of Grid Computing, 12(1):29–40, March 2014. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL

                                        http://link.springer.com/article/
                                        10.1007/s10723-013-9284-1.

Fabra:2014:SIP

[328] Javier Fabra, Sergio Hernández, Joaquín Ezpeleta, and Pedro Álvarez. Solving the interoperability problem by means of a bus. Journal of Grid Computing, 12(1):41–65, March 2014. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL

                                        http://link.springer.com/article/
                                        10.1007/s10723-013-9276-1.

Lordan:2014:SIP

[329] Francesc Lordan, Enric Tejedor, Jorge Ejarque, Roger Rafanell, Javier Álvarez, Fabrizio Marozzo, Daniele Lezzi, Raül Sirvent, Domenico Talia, and Rosa M. Badia. ServiceSs: An interoperable programming framework for the cloud. Journal of Grid Computing, 12(1):67–91, March 2014. CO-DEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL

                                        http://link.springer.com/article/
                                        10.1007/s10723-013-9272-5.

Yangui:2014:COS

[330] Sami Yangui, Iain-James Marshall, Jean-Pierre Laisne, and Samir Tata.

CompatibleOne: The open source cloud broker. Journal of Grid Computing, 12(1):93–109, March 2014. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL

http://link.springer.com/article/
10.1007/s10723-013-9285-0.

Troger:2014:TSJ

[331] Peter Tröger and Andre Merzky. Towards standardized job submission and control in infrastructure clouds. Journal of Grid Computing, 12(1):111–125, March 2014. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10723-013-
9275-2.

Guimaraes:2014:FAF

[332] Felipe Pontes Guimaraes, Pedro Célestin, Daniel Macedo Batista, Genaína Nunes Rodrigues, and Alba Cristina Magalhaes Alves de Melo. A framework for adaptive fault-tolerant execution of workflows in the grid: Empirical and theoretical analysis. Journal of Grid Computing, 12(1):127–151, March 2014. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL

http://link.springer.com/article/
10.1007/s10723-013-9281-4.

Cesario:2014:MDA

[333] Eugenio Cesario, Carlo Mastroianni, and Domenico Talia. A multidomain architecture for mining frequent items and itemsets from distributed data streams. Journal of Grid Computing, 12(1):153–168, March 2014. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL

http://link.springer.com/article/
10.1007/s10723-013-9277-0.

Bacso:2014:EMJ

[334] Gábor Bacsó, Ádám Visegrádi, Attila Kertesz, and Zsolt Németh. On efficiency of multi-job grid allocation based on statistical trace data. Journal of Grid Computing, 12(1):169–186, March 2014. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL

                                        http://link.springer.com/article/
                                        10.1007/s10723-013-9274-3.

Han:2014:GEI

[335] Yanbo Han and Jianwu Wang. Guest Editors’ introduction: Special issue on service and cloud based data integration. Journal of Grid Computing, 12(2):187–189, June 2014. CO-DEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s10723-014-9305-8;   http://
                                        link.springer.com/content/pdf/10.
                                        1007/s10723-014-9305-8.pdf.

Duipmans:2014:TBA

[336] Evert Ferdinand Duipmans, Luís Ferreira Pires, and Luiz Olavo Bonino da Silva Santos. A transformationbased approach to business process management in the cloud. Journal of Grid Computing, 12(2):191–219, June 2014. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL

                                        http://link.springer.com/article/
                                        10.1007/s10723-013-9278-z.

Liu:2014:MID

[337] Chen Liu, Jianwu Wang, and Yanbo Han. Mashroom+: An interactive data mashup approach with uncertainty handling. Journal of Grid Computing, 12(2):221–244, June 2014.

CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL

http://link.springer.com/article/
10.1007/s10723-013-9280-5.

Szabo:2014:SCA

[338] Claudia Szabo, Quan Z. Sheng, Trent Kroeger, Yihong Zhang, and Jian Yu. Science in the cloud: Allocation and execution of data-intensive scientific workflows. Journal of Grid Computing, 12(2):245–264, June 2014. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL

http://link.springer.com/article/
10.1007/s10723-013-9282-3.

Han:2014:MIP

[339] Yuanbin Han, Shizhan Chen, and Zhiyong Feng. Mining integration patterns of programmable ecosystem with social tags. Journal of Grid Computing, 12(2):265–283, June 2014. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL

http://link.springer.com/article/
10.1007/s10723-013-9288-x.

Haque:2014:IMS

[340] Aminul Haque, Saadat M. Alhashmi, and Rajendran Parthiban. Identifying and modeling the strengths and weaknesses of major economic models in grid resource management. Journal of Grid Computing, 12(2):285–302, June 2014. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL

http://link.springer.com/article/
10.1007/s10723-013-9289-9.

Cozzini:2014:RCS

[341] Stefano Cozzini, Deepika Vaddi, Savita Goel, Francesco De Giorgi, and S. K. Dash. Regional climate simulations on EU–INDIA grid infrastructures: Methodologies and performance. Journal of Grid Computing, 12(2):303–320, June 2014. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL

http://link.springer.com/article/
10.1007/s10723-013-9286-z.

Petcu:2014:CRS

[342] Dana Petcu. Consuming resources and services from multiple clouds. Journal of Grid Computing, 12(2):321–345, June 2014. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL

http://link.springer.com/article/
10.1007/s10723-013-9290-3.

Distefano:2014:PDW

[343] Salvatore Distefano and Giuseppe Serazzi. Performance driven WS orchestration and deployment in service oriented infrastructure. Journal of Grid Computing, 12(2):347–369, June 2014. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL

http://link.springer.com/article/
10.1007/s10723-014-9293-8.

Costa:2014:GDT

[344] Genaro Costa, Anna Sikora, Josep Jorba, and Tomàs Margalef. GMATE: Dynamic tuning of parallel applications in grid environment. Journal of Grid Computing, 12(2):371–398, June 2014. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL

http://link.springer.com/article/
10.1007/s10723-013-9287-y.

Qureshi:2014:SGR

[345] Muhammad Bilal Qureshi, Maryam Mehri Dehnavi, Nasro Min-Allah, Muhammad Shuaib Qureshi, Hameed Hussain, Ilias Rentifis, Nikos Tziritas, Thanasis Loukopoulos, Samee U. Khan, Cheng-Zhong Xu, and Albert Y. Zomaya. Survey on grid resource allocation mechanisms. Journal of Grid Computing, 12(2):399–441, June 2014. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL

http://link.springer.com/article/
10.1007/s10723-014-9292-9.

Cuzzocrea:2014:MAH

[346] Alfredo Cuzzocrea. Models and algorithms for high-performance data management and mining on computational grids and clouds. Journal of Grid Computing, 12(3):443–445, September 2014. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL http:

//link.springer.com/article/10.
1007/s10723-014-9316-5;   http://
link.springer.com/content/pdf/10.
1007/s10723-014-9316-5.pdf.

Hsu:2014:ASA

[347] Chih-Hsuan Hsu, Cho-Chin Lin, and Tsan sheng Hsu. Adaptable scheduling algorithm for grids with resource redeployment capability. Journal of Grid Computing, 12(3):447–463, September 2014. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL

http://link.springer.com/article/
10.1007/s10723-014-9298-3.

Muthuvelu:2014:QBT

[348] Nithiapidary Muthuvelu, Ian Chai, Eswaran Chikkannan, and Rajkumar Buyya. QoS-based task group deployment on grid by learning the performance data. Journal of Grid Computing, 12(3):465–483, September 2014. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL

                                        http://link.springer.com/article/
                                        10.1007/s10723-014-9308-5.

Costantini:2014:UID

[349] Alessandro Costantini, Osvaldo Gervasi, Fabiana Zollo, and Luca Caprini. User interaction and data management for large scale grid applications. Journal of Grid Computing, 12(3):485–497, September 2014. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL

                                        http://link.springer.com/article/
                                        10.1007/s10723-014-9300-0.

Rasooli:2014:GSH

[350] Aysan Rasooli and Douglas G. Down. Guidelines for selecting Hadoop schedulers based on system heterogeneity. Journal of Grid Computing, 12(3):499–519, September 2014. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL

                                        http://link.springer.com/article/
                                        10.1007/s10723-014-9299-2.

Kim:2014:SEG

[351] Seonho Kim, Jinoh Kim, and Jon B. Weissman. A security-enabled grid system for MINDS distributed data mining. Journal of Grid Computing, 12(3):521–542, September 2014. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL

                                        http://link.springer.com/article/
                                        10.1007/s10723-014-9303-x.

David:2014:VGM

[352] Mário David, Gonçalo Borges, Jorge Gomes, João Pina, Isabel Campos Plasencia, Enol Fernández del Castillo, Iván Díaz, Carlos Fernandez, Esteban Freire, Álvaro Simón, Kostas Koumantaros, Michel Dreschner, Tiziana Ferrari, and Peter Solagna. Validation of grid middleware for the European Grid Infrastructure. Journal of Grid Computing, 12(3):543–558, September 2014. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL

http://link.springer.com/article/
10.1007/s10723-014-9301-z.

Lorido-Botran:2014:RAS

[353] Tania Lorido-Botran, Jose Miguel-Alonso, and Jose A. Lozano. A review of auto-scaling techniques for elastic applications in cloud environments. Journal of Grid Computing, 12(4):559–592, December 2014. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL

http://link.springer.com/article/
10.1007/s10723-014-9314-7.

Islam:2014:RED

[354] Tanzima Zerin Islam, Saurabh Bagchi, and Rudolf Eigenmann. Reliable and efficient distributed checkpointing system for grid environments. Journal of Grid Computing, 12(4):593–613, December 2014. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL

http://link.springer.com/article/
10.1007/s10723-014-9297-4.

Bagchi:2014:SAE

[355] Susmit Bagchi. The software architecture for efficient distributed interprocess communication in mobile distributed systems. Journal of Grid Computing, 12(4):615–635, December 2014. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL

                                        http://link.springer.com/article/
                                        10.1007/s10723-014-9304-9.

Khajemohammadi:2014:EWS

[356] Hassan Khajemohammadi, Ali Fanian, and T. Aaron Gulliver. Efficient workflow scheduling for grid computing using a leveled multi-objective genetic algorithm. Journal of Grid Computing, 12(4):637–663, December 2014. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL

                                        http://link.springer.com/article/
                                        10.1007/s10723-014-9306-7.

Arabnejad:2014:BCS

[357] Hamid Arabnejad and Jorge G. Barbosa. A budget constrained scheduling algorithm for workflow applications. Journal of Grid Computing, 12(4):665–679, December 2014. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL

                                        http://link.springer.com/article/
                                        10.1007/s10723-014-9294-7.

Ilias:2014:GCR

[358] Miroslav Ilias and Miroslav Dobrucký. Grid computing with relativistic quantum chemistry software. Journal of Grid Computing, 12(4):681–690, December 2014. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL

                                        http://link.springer.com/article/
                                        10.1007/s10723-014-9309-4.

Shiraz:2015:EEC

[359] Muhammad Shiraz, Abdullah Gani, Azra Shamim, Suleman Khan, and Raja Wasim Ahmad. Energy efficient computational offloading framework for mobile cloud computing. Journal of Grid Computing, 13(1):1–18, March 2015. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL

http://link.springer.com/article/
10.1007/s10723-014-9323-6.

Mattmann:2015:RAP

[360] Chris A. Mattmann, Joshua Garcia, Ivo Krka, Daniel Popescu, and Nenad Medvidovic. Revisiting the anatomy and physiology of the grid. Journal of Grid Computing, 13(1):19–34, March 2015. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL http:

//link.springer.com/article/10.
1007/s10723-015-9324-0;   http://
link.springer.com/content/pdf/10.
1007/s10723-015-9324-0.pdf.

Garcia:2015:CSR

[361] Andrés García García and Ignacio Blanquer. Cloud services representation using SLA composition. Journal of Grid Computing, 13(1):35–51, March 2015. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL

http://link.springer.com/article/
10.1007/s10723-014-9295-6.

Caballer:2015:DMV

[362] Miguel Caballer, Ignacio Blanquer, Germán Moltó, and Carlos de Alfonso. Dynamic management of virtual infrastructures. Journal of Grid Computing, 13(1):53–70, March 2015. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL

http://link.springer.com/article/
10.1007/s10723-014-9296-5.

Heikkurinen:2015:ACA

[363] Matti Heikkurinen, Sandra Cohen, Fotis Karagiannis, Kashif Iqbal, Sergio Andreozzi, and Michele Michelotto. Answering the cost assessment scaling challenge: Modelling the annual cost of European computing services for research. Journal of Grid Computing, 13(1):71–94, March 2015. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL

http://link.springer.com/article/
10.1007/s10723-014-9302-y.

Costa:2015:CWA

[364] L. B. Costa, H. Yang, E. Vairavanathan, A. Barros, K. Maheshwari, G. Fedak, D. Katz, M. Wilde, M. Ripeanu, and S. Al-Kiswany. The case for workflow-aware storage: An opportunity study. Journal of Grid Computing, 13(1):95–113, March 2015. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL

http://link.springer.com/article/
10.1007/s10723-014-9307-6.

Castella:2015:DPF

[365] Damià Castellà, Francesc Solsona, and Francesc Giné. DisCoP: A P2P framework for managing and searching computing markets. Journal of Grid Computing, 13(1):115–137, March 2015. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL

http://link.springer.com/article/
10.1007/s10723-014-9318-3.

Chen:2015:PDS

[366] Chongcheng Chen, Jiaxiang Lin, Xiaozhu Wu, and Jianwei Wu. Parallel and distributed spatial outlier mining in grid: Algorithm, design and application. Journal of Grid Computing, 13(2):139–157, June 2015. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL

http://link.springer.com/article/
10.1007/s10723-015-9326-y.

Bencivenni:2015:AGC

[367] Marco Bencivenni, Diego Michelotto, Roberto Alfieri, Riccardo Brunetti, Andrea Ceccanti, Daniele Cesini, Alessandro Costantini, Enrico Fattibene, Luciano Gaido, Giuseppe Misurelli, Elisabetta Ronchieri, Davide Salomoni, Paolo Veronesi, Valerio Venturi, and Maria Cristina Vistoli. Accessing grid and cloud services through a scientific Web portal. Journal of Grid Computing, 13(2):159–175, June 2015. CO-DEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL http:

//link.springer.com/article/10.
1007/s10723-014-9310-y;   http://
link.springer.com/content/pdf/10.
1007/s10723-014-9310-y.pdf.

Prajapati:2015:APV

[368] Harshadkumar B. Prajapati and Vipul A.

Shah. Analysis perspective views of grid simulation tools. Journal of Grid Computing, 13(2):177–213, June 2015. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL

http://link.springer.com/article/
10.1007/s10723-015-9328-9.

Prieto-Castrillo:2015:SPA

[369] Francisco Prieto-Castrillo, Antonio Astillero, and María Botón-Fernández. A stochastic process approach to model distributed computing on complex networks. Journal of Grid Computing, 13(2):215–232, June 2015. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL

http://link.springer.com/article/
10.1007/s10723-014-9317-4.

Ebrahimirad:2015:EAS

[370] Vahid Ebrahimirad, Maziar Goudarzi, and Aboozar Rajabi. Energy-aware scheduling for precedence-constrained parallel virtual machines in virtualized data centers. Journal of Grid Computing, 13(2):233–253, June 2015. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL

http://link.springer.com/article/
10.1007/s10723-015-9327-x.

Kokkinos:2015:SAO

[371] P. Kokkinos, T. A. Varvarigou, A. Kretsis, P. Soumplis, and E. A. Varvarigos. SuMo: Analysis and optimization of Amazon EC2 instances. Journal of Grid Computing, 13(2):255–274, June 2015. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL

http://link.springer.com/article/
10.1007/s10723-014-9311-x.

Elkhatib:2015:PNA

[372] Yehia Elkhatib and Chris Edwards. Passive network awareness as a means for improved grid scheduling. Journal of Grid Computing, 13(2):275–291, June 2015. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL

http://link.springer.com/article/
10.1007/s10723-015-9332-0.

Carvalho:2015:SPM

[373] Diego Carvalho, Luiz Rossi de Souza, Rafael G. Barbastefano, and Felipe M. G. França. Stochastic productmix: A grid computing industrial application. Journal of Grid Computing, 13(2):293–304, June 2015. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL

http://link.springer.com/article/
10.1007/s10723-015-9325-z.

Hidalgo:2015:PBA

[374] J. Ignacio Hidalgo and Francisco Fernández de Vega. Parallel bioinspired algorithms on the grid and cloud. Journal of Grid Computing, 13(3):305–308, September 2015. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL

http://link.springer.com/article/
10.1007/s10723-014-9322-7.

Lutton:2015:VAM

[375] Evelyne Lutton, Hugo Gilbert, Waldo Cancino, Benjamin Bach, Joseph Pallamidessi, Pierre Parrend, and Pierre Collet. Visual and audio monitoring of island based parallel evolutionary algorithms. Journal of Grid Computing, 13(3):309–327, September 2015. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL

http://link.springer.com/article/
10.1007/s10723-014-9321-8.

Garcia-Valdez:2015:EMP

[376] Mario García-Valdez, Leonardo Trujillo, Juan-J Merelo, Francisco Fernández de Vega, and Gustavo Olague. The EvoSpace model for pool-based evolutionary algorithms. Journal of Grid Computing, 13(3):329–349, September 2015. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL

http://link.springer.com/article/
10.1007/s10723-014-9319-2.

Nogueras:2015:SFT

[377] Rafael Nogueras and Carlos Cotta.

Studying fault-tolerance in island-based evolutionary and multimemetic algorithms. Journal of Grid Computing, 13(3):351–374, September 2015. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL

                                        http://link.springer.com/article/
                                        10.1007/s10723-014-9315-6.

Pascual:2015:TGC

[378] J. A. Pascual, T. Lorido-Botrán, J. Miguel-Alonso, and J. A. Lozano. Towards a greener cloud infrastructure management using optimized placement policies. Journal of Grid Computing, 13(3):375–389, September 2015. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL

                                        http://link.springer.com/article/
                                        10.1007/s10723-014-9312-9.

Veeramachaneni:2015:FCB

[379] Kalyan Veeramachaneni, Ignacio Arnaldo, Owen Derby, and Una-May O’Reilly. FlexGP: Cloud-based ensemble learning with genetic programming for large regression problems. Journal of Grid Computing, 13(3):391–407, September 2015. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL

                                        http://link.springer.com/article/
                                        10.1007/s10723-014-9320-9.

Arroba:2015:ERM

[380] Patricia Arroba, José L. Risco-Martín, Marina Zapater, José M. Moya, and José L. Ayala. Enhancing regression models for complex systems using evolutionary techniques for feature engineering. Journal of Grid Computing, 13(3):409–423, September 2015. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL

                                        http://link.springer.com/article/
                                        10.1007/s10723-014-9313-8.

Bernabe:2015:ITS

[381] Jorge Bernal Bernabe, Gregorio Martinez Perez, and Antonio F. Skarmeta Gomez. Intercloud trust and security decision support system: an ontologybased approach. Journal of Grid Computing, 13(3):425–456, September 2015. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL

http://link.springer.com/article/
10.1007/s10723-015-9346-7.

Liu:2015:SDI

[382] Ji Liu, Esther Pacitti, Patrick Valduriez, and Marta Mattoso. A survey of data-intensive scientific workflow management. Journal of Grid Computing, 13(4):457–493, December 2015. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL

http://link.springer.com/article/
10.1007/s10723-015-9329-8.

Verma:2015:CTE

[383] Amandeep Verma and Sakshi Kaushal.

Cost-time efficient scheduling plan for executing workflows in the cloud. Journal of Grid Computing, 13(4):495– 506, December 2015. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s10723-015-
9344-9.

Mei:2015:FTD

[384] Jing Mei, Kenli Li, Xu Zhou, and Keqin Li. Fault-tolerant dynamic rescheduling for heterogeneous computing systems. Journal of Grid Computing, 13(4):507–525, December 2015. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL

http://link.springer.com/article/
10.1007/s10723-015-9331-1.

Lent:2015:GSM

[385] Ricardo Lent. Grid scheduling with makespan and energy-based goals. Journal of Grid Computing, 13(4):527– 546, December 2015. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s10723-015-
                                        9349-4.

Costa:2015:ISG

[386] Alessandro Costa, Pietro Massimino, and Marilena Bandieramonte. An innovative science gateway for the Cherenkov Telescope Array. Journal of Grid Computing, 13(4):547–559, December 2015. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL

                                        http://link.springer.com/article/
                                        10.1007/s10723-015-9330-2.

Mrozek:2015:SIP

[387] Dariusz Mrozek, Pawel Gosk, and Bozena Malysiak-Mrozek. Scaling ab initio predictions of 3D protein structures in Microsoft Azure Cloud. Journal of Grid Computing, 13(4): 561–585, December 2015. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s10723-015-9353-8;   http://
                                        link.springer.com/content/pdf/10.
                                        1007/s10723-015-9353-8.pdf.

Wang:2015:IMP

[388] Yaoguang Wang, Weiming Lu, Renjie Lou, and Baogang Wei. Improving MapReduce performance with partial speculative execution. Journal of Grid Computing, 13(4):587–604, December 2015. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL

http://link.springer.com/article/
10.1007/s10723-015-9350-y.

Khan:2015:MAD

[389] Atta ur Rehman Khan, Mazliza Othman, and Abdul Nasir Khan. MobiByte: An application development model for mobile cloud computing. Journal of Grid Computing, 13(4):605–628, December 2015. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL

http://link.springer.com/article/
10.1007/s10723-015-9335-x.

Distefano:2015:QAM

[390] Salvatore Distefano, Francesco Longo, and Marco Scarpa. QoS assessment of mobile crowdsensing services. Journal of Grid Computing, 13(4):629–650, December 2015. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL

http://link.springer.com/article/
10.1007/s10723-015-9338-7.

Khan:2015:CMB

[391] Abdul Nasir Khan, M. L. Mat Kiah, and Mazhar Ali. A cloud-manager-based re-encryption scheme for mobile users in cloud environment: a hybrid approach. Journal of Grid Computing, 13(4):651–675, December 2015. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL

http://link.springer.com/article/
10.1007/s10723-015-9352-9.

Dorronsoro:2016:E

[392] Bernabé Dorronsoro, Dzmitry Kliazovich, Pascal Bouvry, and Albert Y. Zomaya. Editorial. Journal of Grid Computing, 14(1):1–3, March 2016. CO-DEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL http:

//link.springer.com/article/10.
1007/s10723-016-9365-z;   http://
link.springer.com/content/pdf/10.
1007/s10723-016-9365-z.pdf.

Tchernykh:2016:OBO

[393] Andrei Tchernykh, Luz Lozano, Uwe Schwiegelshohn, Pascal Bouvry, Johnatan E. Pecero, Sergio Nesmachnow, and Alexander Yu. Drozdov. Online biobjective scheduling for IaaS clouds ensuring quality of service. Journal of Grid Computing, 14(1):5–22, March 2016. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL

http://link.springer.com/article/
10.1007/s10723-015-9340-0.

Kliazovich:2016:CDM

[394] Dzmitry Kliazovich, Johnatan E. Pecero, Andrei Tchernykh, Pascal Bouvry, Samee U. Khan, and Albert Y. Zomaya. CA-DAG: Modeling communication-aware applications for scheduling in cloud computing. Journal of Grid Computing, 14(1):23–39, March 2016. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL

http://link.springer.com/article/
10.1007/s10723-015-9337-8.

Kertesz:2016:PBV

[395] A. Kertesz, J. D. Dombi, and A. Benyi. A pliant-based virtual machine scheduling solution to improve the energy efficiency of IaaS clouds. Journal of Grid Computing, 14(1):41–53, March 2016. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL

http://link.springer.com/article/
10.1007/s10723-015-9336-9.

Tang:2016:EET

[396] Zhuo Tang, Ling Qi, Zhenzhen Cheng, Kenli Li, Samee U. Khan, and Keqin Li. An energy-efficient task scheduling algorithm in DVFS-enabled cloud environment. Journal of Grid Computing, 14(1):55–74, March 2016. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL

http://link.springer.com/article/
10.1007/s10723-015-9334-y.

daSilva:2016:TAV

[397] Rodrigo A. C. da Silva and Nelson L. S.

da Fonseca. Topology-aware virtual machine placement in data centers. Journal of Grid Computing, 14(1):75–90, March 2016. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL

http://link.springer.com/article/
10.1007/s10723-015-9343-x.

Borgetto:2016:HAE

[398] D. Borgetto, R. Chakode, B. Depardon, C. Eichler, J. M. Garcia, H. Hbaieb, T. Monteil, E. Pelorce, A. Rachdi, A. Al Sheikh, and P. Stolf. Hybrid approach for energy aware management of multi-cloud architecture integrating user machines. Journal of Grid Computing, 14(1):91–108, March 2016. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL

http://link.springer.com/article/
10.1007/s10723-015-9342-y.

Junior:2016:GCM

[399] Osvaldo Adilson de Carvalho Junior, Sarita Mazzini Bruschi, Regina Helena Carlucci Santana, and Marcos José Santana. Green cloud metascheduling. Journal of Grid Computing, 14(1):109–126, March 2016.

CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL

http://link.springer.com/article/
10.1007/s10723-015-9333-z.

Borylo:2016:GCP

[400] Piotr Borylo, Artur Lason, Jacek Rzasa, Andrzej Szymanski, and Andrzej Jajszczyk. Green cloud provisioning throughout cooperation of a WDM wide area network and a hybrid power IT infrastructure. Journal of Grid Computing, 14(1):127–151, March 2016. CO-DEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL http:

//link.springer.com/article/10.
1007/s10723-015-9354-7;   http://
link.springer.com/content/pdf/10.
1007/s10723-015-9354-7.pdf.

Marszalkowski:2016:TEP

[401] Jedrzej M. Marszalkowski, Maciej Drozdowski, and Jakub Marszalkowski. Time and energy performance of parallel systems with hierarchical memory. Journal of Grid Computing, 14(1):153–170, March 2016. CO-DEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL http:

//link.springer.com/article/10.
1007/s10723-015-9345-8;   http://
link.springer.com/content/pdf/10.
1007/s10723-015-9345-8.pdf.

Shi:2016:PPA

[402] Xiaoyu Shi, Jin Dong, Seddik M. Djouadi, Yong Feng, Xiao Ma, and Yefu Wang. PAPMSC: Power-aware performance management approach for virtualized Web servers via stochastic control. Journal of Grid Computing, 14(1):171–191, March 2016. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL

http://link.springer.com/article/
10.1007/s10723-015-9341-z.

Galante:2016:APC

[403] Guilherme Galante, Luis Carlos Erpen De Bona, Antonio Roberto Mury, Bruno Schulze, and Rodrigo da Rosa Righi. An analysis of public clouds elasticity in the execution of scientific applications: a survey. Journal of Grid Computing, 14(2):193–216, June 2016. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL

http://link.springer.com/article/
10.1007/s10723-016-9361-3.

Singh:2016:SRS

[404] Sukhpal Singh and Inderveer Chana.

A survey on resource scheduling in cloud computing: Issues and challenges. Journal of Grid Computing, 14(2):217–264, June 2016. CO-DEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL

http://link.springer.com/article/
10.1007/s10723-015-9359-2.

Peinl:2016:DCM

[405] René Peinl, Florian Holzschuher, and Florian Pfitzer. Docker cluster management for the cloud — survey results and own solution. Journal of Grid Computing, 14(2):265–282, June 2016. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL

http://link.springer.com/article/
10.1007/s10723-016-9366-y.

Gutierrez-Aguado:2016:IMA

[406] Juan Gutierrez-Aguado, Jose M. Alcaraz Calero, and Wladimiro Diaz Villanueva. IaaSMon: Monitoring architecture for public cloud computing data centers. Journal of Grid Computing, 14(2):283–297, June 2016. CO-DEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL http:

//link.springer.com/article/10.
1007/s10723-015-9357-4;   http://
link.springer.com/content/pdf/10.
1007/s10723-015-9357-4.pdf.

Sharifi:2016:EEC

[407] Leila Sharifi, Llorenç Cerdà-Alabern, Felix Freitag, and Luís Veiga. Energy efficient cloud service provisioning: Keeping data center granularity in perspective. Journal of Grid Computing, 14(2):299–325, June 2016. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL

http://link.springer.com/article/
10.1007/s10723-015-9358-3.

Kansal:2016:EAV

[408] Nidhi Jain Kansal and Inderveer Chana. Energy-aware virtual machine migration for cloud computing — a firefly optimization approach. Journal of Grid Computing, 14(2):327–345, June 2016. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL

http://link.springer.com/article/
10.1007/s10723-016-9364-0.

Bacso:2016:SSJ

[409] Gábor Bacsó, Tamás Kis, Ádám Visegrádi, Attila Kertész, and Zsolt Németh. A set of successive job allocation models in distributed computing infrastructures. Journal of Grid Computing, 14(2):347–358, June 2016. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL

http://link.springer.com/article/
10.1007/s10723-015-9347-6.

Bryk:2016:SAA

[410] Piotr Bryk, Maciej Malawski, Gideon Juve, and Ewa Deelman. Storage-aware algorithms for scheduling of workflow ensembles in clouds. Journal of Grid Computing, 14(2):359–378, June 2016. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL http:

//link.springer.com/article/10.
1007/s10723-015-9355-6;   http://
link.springer.com/content/pdf/10.
1007/s10723-015-9355-6.pdf.

Bajaber:2016:BDP

[411] Fuad Bajaber, Radwa Elshawi, Omar Batarfi, Abdulrahman Altalhi, Ahmed Barnawi, and Sherif Sakr. Big Data 2.0 processing systems: Taxonomy and open challenges. Journal of Grid Computing, 14(3):379–405, September 2016. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (elec-

tronic). URL http://link.springer.
com/accesspage/article/10.1007/
s10723-016-9371-1;   http://link.
springer.com/article/10.1007/s10723-
016-9371-1.

Zhang:2016:HDC

[412] Jing Zhang, Qianmu Li, and Wei Zhou.

HDCache: A distributed cache system for real-time cloud services. Journal of Grid Computing, 14(3):407– 428, September 2016. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (elec-

tronic). URL http://link.springer.
com/accesspage/article/10.1007/
s10723-015-9360-9;   http://link.
springer.com/article/10.1007/s10723-
015-9360-9.

Bazinet:2016:SLR

[413] Adam L. Bazinet and Michael P.

Cummings. Subdividing long-running, variable-length analyses into short, fixed-length BOINC workunits. Journal of Grid Computing, 14(3):429– 441, September 2016. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL http:

//link.springer.com/article/10.
1007/s10723-015-9348-5;   http://
link.springer.com/content/pdf/10.
1007/s10723-015-9348-5.pdf.

Coutinho:2016:DCD

[414] Rafaelli Coutinho, Yuri Frota, Kary Ocaña, Daniel de Oliveira, and Lúcia M. A. Drummond. A dynamic cloud dimensioning approach for parallel scientific workflows: a case study in the comparative genomics domain. Journal of Grid Computing, 14(3):443– 461, September 2016. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (elec-

tronic). URL http://link.springer.
com/accesspage/article/10.1007/
s10723-016-9367-x;   http://link.
springer.com/article/10.1007/s10723-
016-9367-x.

Yoo:2016:TSF

[415] Wucherl Yoo and Alex Sim. Time-series forecast modeling on high-bandwidth network measurements. Journal of Grid Computing, 14(3):463–476, September 2016. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (elec-

tronic). URL http://link.springer.
com/accesspage/article/10.1007/
s10723-016-9368-9;   http://link.
springer.com/article/10.1007/s10723-
016-9368-9.

Aziz:2016:MFG

[416] Benjamin Aziz. Modelling fine-grained access control policies in grids. Journal of Grid Computing, 14(3):477– 493, September 2016. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (elec-

tronic). URL http://link.springer.
com/accesspage/article/10.1007/
s10723-015-9351-x;   http://link.
springer.com/article/10.1007/s10723-
015-9351-x.

Gesing:2016:SGW

[417] Sandra Gesing, Nancy Wilkins-Diehr, Michelle Barker, and Gabriele Pierantoni. Science Gateway Workshops 2015 special issue conference publications. Journal of Grid Computing, 14(4):495–498, December 2016. CO-DEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL http:

//link.springer.com/article/10.
1007/s10723-016-9389-4;   http://
link.springer.com/content/pdf/10.
1007/s10723-016-9389-4.pdf.

Aguilera:2016:AGI

[418] Alvaro Aguilera, Richard Grunzke, Dirk Habich, Johannes Luong, Dirk Schollbach, Ulf Markwardt, and Jochen Garcke. Advancing a gateway infrastructure for wind turbine data analysis. Journal of Grid Computing, 14(4):499– 514, December 2016. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (elec-

tronic). URL http://link.springer.
com/accesspage/article/10.1007/
s10723-016-9376-9;   http://link.
springer.com/article/10.1007/s10723-
016-9376-9.

Sinnott:2016:BDR

[419] Richard O. Sinnott, Felix Beuschlein, Jemie Effendy, Graeme Eisenhofer, Stephan Gloeckner, and Anthony Stell. Beyond a disease registry: An integrated virtual environment for adrenal cancer research. Journal of Grid Computing, 14(4):515–532, December 2016. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (elec-

                                        tronic). URL http://link.springer.
                                        com/accesspage/article/10.1007/
                                        s10723-016-9375-x;   http://link.
                                        springer.com/article/10.1007/s10723-
                                        016-9375-x.

McGrath:2016:ECS

[420] Annette McGrath, Steve McMahon, Sean Li, Joel Ludbey, and Tim Ho. The essential components of a successful galaxy service. Journal of Grid Computing, 14(4):533–543, December 2016. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (elec-

                                        tronic). URL http://link.springer.
                                        com/accesspage/article/10.1007/
                                        s10723-016-9379-6;   http://link.
                                        springer.com/article/10.1007/s10723-
                                        016-9379-6.

Gesing:2016:USG

[421] Sandra Gesing, Jens Krüger, Richard Grunzke, Sonja Herres-Pawlis, and Alexander Hoffmann. Using science gateways for bridging the differences between research infrastructures. Journal of Grid Computing, 14(4):545– 557, December 2016. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (elec-

                                        tronic). URL http://link.springer.
                                        com/accesspage/article/10.1007/
                                        s10723-016-9385-8;   http://link.
                                        springer.com/article/10.1007/s10723-
                                        016-9385-8.

Piontek:2016:DSG

[422] T. Piontek, B. Bosak, M. Ciznicki, P. Grabowski, P. Kopta, M. Kulczewski, D. Szejnfeld, and K. Kurowski. Development of science gateways using QCG — lessons learned from the deployment on large scale distributed and HPC infrastructures. Journal of Grid Computing, 14(4):559–573, December 2016. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (elec-

tronic). URL http://link.springer.
com/accesspage/article/10.1007/
s10723-016-9384-9;   http://link.
springer.com/article/10.1007/s10723-
016-9384-9.

DAgostino:2016:LLR

[423] Daniele D’Agostino, Emanuele Danovaro, Andrea Clematis, Luca Roverelli, Gabriele Zereik, and Antonella Galizia. From lesson learned to the refactoring of the DRIHM science gateway for hydro-meteorological research. Journal of Grid Computing, 14(4):575– 588, December 2016. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (elec-

tronic). URL http://link.springer.
com/accesspage/article/10.1007/
s10723-016-9377-8;   http://link.
springer.com/article/10.1007/s10723-
016-9377-8.

Gugnani:2016:ESG

[424] Shashank Gugnani, Carlos Blanco, Tamas Kiss, and Gabor Terstyanszky. Extending science gateway frameworks to support big data applications in the cloud. Journal of Grid Computing, 14(4):589–601, December 2016. CO-DEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL http:

//link.springer.com/article/10.
1007/s10723-016-9369-8;   http://
link.springer.com/content/pdf/10.
1007/s10723-016-9369-8.pdf.

Sinnott:2016:PPG

[425] Richard O. Sinnott, Christopher Bayliss, Andrew Bromage, Gerson Galang, Yikai Gong, Philip Greenwood, Glenn Jayaputera, Davis Marques, Luca Morandini, Ghazal Nogoorani, Hossein Pursultani, Muhammad Sarwar, William Voorsluys, and Ivo Widjaja. Privacy preserving geo-linkage in the big urban data era. Journal of Grid Computing, 14(4):603– 618, December 2016. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (elec-

tronic). URL http://link.springer.
com/accesspage/article/10.1007/
s10723-016-9372-0;   http://link.
springer.com/article/10.1007/s10723-
016-9372-0.

Farkas:2016:EWO

[426] Zoltán Farkas, Péter Kacsuk, and Ákos Hajnal. Enabling workflow-oriented science gateways to access multi-cloud systems. Journal of Grid Computing, 14(4): 619–640, December 2016. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (elec-

tronic). URL http://link.springer.
com/accesspage/article/10.1007/
s10723-016-9388-5;   http://link.
springer.com/article/10.1007/s10723-
016-9388-5.

Kacsuk:2016:IAS

[427] Peter Kacsuk, Gabor Kecskemeti, Attila Kertesz, Zsolt Nemeth, József Kovács, and Zoltán Farkas. Infrastructure aware scientific workflows and infrastructure aware workflow managers in science gateways. Journal of Grid Computing, 14(4):641–654, December 2016. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (elec-

tronic). URL http://link.springer.
com/accesspage/article/10.1007/
s10723-016-9380-0;   http://link.
springer.com/article/10.1007/s10723-
016-9380-0.

Arshad:2016:FAS

[428] Junaid Arshad, Gabor Terstyanszky, Tamas Kiss, Noam Weingarten, and Giuliano Taffoni. A formal approach to support interoperability in scientific meta-workflows. Journal of Grid Computing, 14(4):655–671, December 2016. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (elec-

tronic). URL http://link.springer.
com/accesspage/article/10.1007/
s10723-016-9383-x;   http://link.
springer.com/article/10.1007/s10723-
016-9383-x.

Sanchez-Exposito:2016:WSB

[429] S. Sánchez-Expósito, P. Martín, J. E.

Ruiz, L. Verdes-Montenegro, J. Garrido, R. Sirvent, A. Ruiz Falcó, R. M. Badia, and D. Lezzi. Web services as building blocks for science gateways in astrophysics. Journal of Grid Computing, 14(4):673–685, December 2016. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (elec-

tronic). URL http://link.springer.
com/accesspage/article/10.1007/
s10723-016-9382-y;   http://link.
springer.com/article/10.1007/s10723-
016-9382-y.

Karoczkai:2016:MBF

[430] Krisztian Karoczkai, Attila Kertesz, and Peter Kacsuk. A meta-brokering framework for science gateways. Journal of Grid Computing, 14(4):687– 703, December 2016. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (elec-

tronic). URL http://link.springer.
com/accesspage/article/10.1007/
s10723-016-9378-7;   http://link.
springer.com/article/10.1007/s10723-
016-9378-7.

Matos:2017:REP

[431] Rubens Matos, Jamilson Dantas, Jean Araujo, Kishor S. Trivedi, and Paulo Maciel. Redundant Eucalyptus private clouds: Availability modeling and sensitivity analysis. Journal of Grid Computing, 15(1):1–22, March 2017. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (elec-

                                        tronic). URL http://link.springer.
                                        com/accesspage/article/10.1007/
                                        s10723-016-9381-z;   http://link.
                                        springer.com/article/10.1007/s10723-
                                        016-9381-z.

Wang:2017:RAT

[432] Shuli Wang, Kenli Li, Jing Mei, Guoqing Xiao, and Keqin Li. A reliabilityaware task scheduling algorithm based on replication on heterogeneous computing systems. Journal of Grid Computing, 15(1):23–39, March 2017. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (elec-

                                        tronic). URL http://link.springer.
                                        com/accesspage/article/10.1007/
                                        s10723-016-9386-7;   http://link.
                                        springer.com/article/10.1007/s10723-
                                        016-9386-7.

Grunzke:2017:MMM

[433] Richard Grunzke, Jens Krüger, René Jäkel, Wolfgang E. Nagel, Sonja Herres-Pawlis, and Alexander Hoffmann. Metadata management in the MoSGrid science gateway — evaluation and the expansion of quantum chemistry support. Journal of Grid Computing, 15 (1):41–53, March 2017. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (elec-

                                        tronic). URL http://link.springer.
                                        com/accesspage/article/10.1007/
                                        s10723-016-9362-2;   http://link.

springer.com/article/10.1007/s10723-
016-9362-2.

Hirsch:2017:TPE

[434] Matías Hirsch, Juan Manuel Rodríguez, Cristian Mateos, and Alejandro Zunino. A two-phase energy-aware scheduling approach for CPU-intensive jobs in mobile grids. Journal of Grid Computing, 15(1):55–80, March 2017. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (elec-

tronic). URL http://link.springer.
com/accesspage/article/10.1007/
s10723-016-9387-6;   http://link.
springer.com/article/10.1007/s10723-
016-9387-6.

Sidhu:2017:ITM

[435] Jagpreet Sidhu and Sarbjeet Singh. Improved TOPSIS method based trust evaluation framework for determining trustworthiness of cloud service providers. Journal of Grid Computing, 15(1):81–105, March 2017. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (elec-

tronic). URL http://link.springer.
com/accesspage/article/10.1007/
s10723-016-9363-1;   http://link.
springer.com/article/10.1007/s10723-
016-9363-1.

Kontodimas:2017:AES

[436] Konstantinos Kontodimas, Panagiotis Kokkinos, Yossi Kuperman, Athanasios Houbavlis, and Emmanouel Varvarigos. Analysis and evaluation of scheduling policies for consolidated I/O operations. Journal of Grid Computing, 15 (1):107–125, March 2017. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (elec-

tronic). URL http://link.springer.
com/accesspage/article/10.1007/
s10723-017-9392-4;   http://link.

                                        springer.com/article/10.1007/s10723-
                                        017-9392-4.

Field:2017:EIC

[437] Laurence Field and Rizos Sakellariou. An evaluation of information consistency in grid information systems. Journal of Grid Computing, 15(1):127–137, March 2017. CO-DEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s10723-016-9373-z;   http://
                                        link.springer.com/content/pdf/10.
                                        1007/s10723-016-9373-z.pdf.

Ferguson:2017:JGC

[438] Donald F. Ferguson and Víctor Méndez Muñoz. Journal of Grid Computing, special issue of cloud computing and services science. Journal of Grid Computing, 15(2):139–140, June 2017. CO-DEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s10723-017-9398-y;   https:
                                        //link.springer.com/content/pdf/
                                        10.1007/s10723-017-9398-y.pdf.

Moreno-Vozmediano:2017:IPF

[439] Rafael Moreno-Vozmediano, Rubén S. Montero, Eduardo Huedo, and Ignacio M. Llorente. Implementation and provisioning of federated networks in hybrid clouds. Journal of Grid Computing, 15(2):141–160, June 2017. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s10723-017-9395-
                                        1.

Lopez-Pires:2017:MOV

[440] Fabio López-Pires and Benjamín Barán.

Many-objective virtual machine placement. Journal of Grid Computing, 15 (2):161–176, June 2017. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (elec-

tronic). URL https://link.springer.
com/article/10.1007/s10723-017-
9399-x.

Chang:2017:FMP

[441] Victor Chang and Muthu Ramachandran. Financial modeling and prediction as a service. Journal of Grid Computing, 15(2):177–195, June 2017. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic).

URL https://link.springer.com/
article/10.1007/s10723-017-9393-
3.

Sidhu:2017:DCA

[442] Jagpreet Sidhu and Sarbjeet Singh.

Design and comparative analysis of MCDM-based multi-dimensional trust evaluation schemes for determining trustworthiness of cloud service providers. Journal of Grid Computing, 15(2):197–218, June 2017. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (elec-

tronic). URL https://link.springer.
com/article/10.1007/s10723-017-
9396-0.

Verginadis:2017:PHD

[443] Yiannis Verginadis, Antonis Michalas, Panagiotis Gouvas, Gunther Schiefer, Gerald Hübsch, and Iraklis Paraskakis. PaaSword: A holistic data privacy and security by design framework for cloud services. Journal of Grid Computing, 15(2):219–234, June 2017. CO-DEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL https:

//link.springer.com/article/10.
1007/s10723-017-9394-2;   https:
//link.springer.com/content/pdf/
10.1007/s10723-017-9394-2.pdf.

Peris:2017:DLB

[444] Antonio Delgado Peris, José M. Hernández, and Eduardo Huedo. Distributed late-binding scheduling and cooperative data caching. Journal of Grid Computing, 15(2):235–256, June 2017. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic).

URL https://link.springer.com/
article/10.1007/s10723-016-9374-
y.

Cai:2017:ETS

[445] Zhicheng Cai, Qianmu Li, and Xiaoping Li. ElasticSim: A toolkit for simulating workflows with cloud resource runtime auto-scaling and stochastic task execution times. Journal of Grid Computing, 15(2):257–272, June 2017. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic).

URL https://link.springer.com/
article/10.1007/s10723-016-9390-
y.

Tighe:2017:TAA

[446] Michael Tighe and Michael Bauer. Topology and application aware dynamic VM management in the cloud. Journal of Grid Computing, 15(2): 273–294, June 2017. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (elec-

tronic). URL https://link.springer.
com/article/10.1007/s10723-017-
9397-z;   https://link.springer.
com/content/pdf/10.1007/s10723-
017-9397-z.pdf.

Khezr:2017:MAC

[447] Seyed Nima Khezr and Nima Jafari Navimipour. MapReduce and its applications, challenges, and architecture: a comprehensive review and directions for future research. Journal of Grid Computing, 15(3):295–321, September 2017. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic).

URL https://link.springer.com/
article/10.1007/s10723-017-9408-
0.

Tang:2017:PCR

[448] Zhuo Tang, Zhongming Fu, Zherong Gong, Kenli Li, and Keqin Li. A parallel conditional random fields model based on Spark computing environment. Journal of Grid Computing, 15(3):323– 342, September 2017. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (elec-

tronic). URL https://link.springer.
com/article/10.1007/s10723-017-
9404-4.

Nawrocki:2017:ACB

[449] Piotr Nawrocki and Bartlomiej Sniezynski. Autonomous context-based service optimization in mobile cloud computing. Journal of Grid Computing, 15(3):343–356, September 2017. CO-DEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL https:

//link.springer.com/article/10.
1007/s10723-017-9406-2;   https:
//link.springer.com/content/pdf/
10.1007/s10723-017-9406-2.pdf.

Lordan:2017:CMP

[450] F. Lordan and Rosa M. Badia.

COMPSs-Mobile: Parallel programming for mobile cloud computing. Journal of Grid Computing, 15(3):357– 378, September 2017. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (elec-

                                        tronic). URL https://link.springer.
                                        com/article/10.1007/s10723-017-
                                        9409-z.

Rahmani:2017:FGD

[451] Amir Masoud Rahmani, Leila Azari, and Helder A. Daniel. A file group data replication algorithm for data grids. Journal of Grid Computing, 15(3):379– 393, September 2017. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (elec-

                                        tronic). URL https://link.springer.
                                        com/article/10.1007/s10723-017-
                                        9407-1.

Kananizadeh:2017:PMT

[452] Shahrzad Kananizadeh and Kirill Kononenko. Predictive mitigation of timing channels — threat defense for machine codes. Journal of Grid Computing, 15(3):395–414, September 2017. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s10723-017-9403-
                                        5.

Ramirez-Velarde:2017:ARA

[453] Raul Ramírez-Velarde, Andrei Tchernykh, Carlos Barba-Jimenez, Adán Hirales-Carbajal, and Juan Nolazco-Flores. Adaptive resource allocation with job runtime uncertainty. Journal of Grid Computing, 15(4):415–434, December 2017. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s10723-017-9410-

Jiang:2017:TEO

[454] Junqiang Jiang, Yaping Lin, Guoqi Xie, Li Fu, and Junfeng Yang. Time and energy optimization algorithms for the static scheduling of multiple workflows in heterogeneous computing system. Journal of Grid Computing, 15(4):435–456, December 2017. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic).

URL https://link.springer.com/
article/10.1007/s10723-017-9391-
5.

Li:2017:CRA

[455] Bo Li, Songtao Guo, Yan Wu, and Defang Liu. Construction and resource allocation of cost-efficient clustered virtual network in software defined networks. Journal of Grid Computing, 15(4):457–473, December 2017. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic).

URL https://link.springer.com/
article/10.1007/s10723-017-9402-
6.

Classe:2017:DIS

[456] Tadeu Classe, Regina Braga, José Maria N. David, Fernanda Campos, and Wagner Arbex. A distributed infrastructure to support scientific experiments. Journal of Grid Computing, 15(4):475–500, December 2017. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic).

URL https://link.springer.com/
article/10.1007/s10723-017-9401-
7.

Bruno:2017:FEM

[457] Rodrigo Bruno, Fernando Costa, and Paulo Ferreira. freeCycles — efficient multi-cloud computing platform. Journal of Grid Computing, 15(4):501– 526, December 2017. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (elec-

tronic). URL https://link.springer.
com/article/10.1007/s10723-017-
9414-2.

Tihanyi:2017:CEL

[458] Norbert Tihanyi, Attila Kovács, and József Kovács. Computing extremely large values of the Riemann zeta function. Journal of Grid Computing, 15(4): 527–534, December 2017. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (elec-

tronic). URL https://link.springer.
com/article/10.1007/s10723-017-
9416-0.

Rodrigues:2017:TEL

[459] Vinicius Facco Rodrigues, Rodrigo da Rosa Righi, Gustavo Rostirolla, Jorge Luis Victória Barbosa, Cristiano André da Costa, Antônio Marcos Alberti, and Victor Chang. Towards enabling live thresholding as utility to manage elastic master–slave applications in the cloud. Journal of Grid Computing, 15(4):535–556, December 2017. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic).

URL https://link.springer.com/
article/10.1007/s10723-017-9405-
3.

Mao:2017:CPM

[460] Hongyan Mao, Zhengwei Qi, Jiangang Duan, and Xinni Ge. Cost-performance modeling with automated benchmarking on elastic computing clouds. Journal of Grid Computing, 15(4):557–572, December 2017. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic).

URL https://link.springer.com/

article/10.1007/s10723-017-9412-
4.

Caballer:2018:GEI

[461] Miguel Caballer, Germán Moltó, and Ignacio Blanquer. Guest Editor’s introduction: Special issue on cloud computing orchestration. Journal of Grid Computing, 16(1):1–2, March 2018. CO-DEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL https:

//link.springer.com/article/10.
1007/s10723-018-9427-5;   https:
//link.springer.com/content/pdf/
10.1007/s10723-018-9427-5.pdf.

Caballer:2018:OCA

[462] Miguel Caballer, Sahdev Zala, Álvaro López García, Germán Moltó, Pablo Orviz Fernández, and Mathieu Velten. Orchestrating complex application architectures in heterogeneous clouds. Journal of Grid Computing, 16(1):3–18, March 2018. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic).

URL https://link.springer.com/
article/10.1007/s10723-017-9418-
y.

Kovacs:2018:OMC

[463] József Kovács and Péter Kacsuk. Occopus: a multi-cloud orchestrator to deploy and manage complex scientific infrastructures. Journal of Grid Computing, 16(1):19–37, March 2018. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic).

URL https://link.springer.com/
article/10.1007/s10723-017-9421-
3.

Moreno-Vozmediano:2018:ODH

[464] R. Moreno-Vozmediano, R. S. Montero, E. Huedo, and I. M. Llorente.

Orchestrating the deployment of high availability services on multi-zone and multi-cloud scenarios. Journal of Grid Computing, 16(1):39–53, March 2018. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic).

URL https://link.springer.com/
article/10.1007/s10723-017-9417-
z.

Kacsuk:2018:FCO

[465] Peter Kacsuk, József Kovács, and Zoltán Farkas. The Flowbster cloudoriented workflow system to process large scientific data sets. Journal of Grid Computing, 16(1):55–83, March 2018. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic).

URL https://link.springer.com/
article/10.1007/s10723-017-9420-
4.

Pascinski:2018:QAO

[466] Uros Pascinski, Jernej Trnkoczy, Vlado Stankovski, Matej Cigale, and Sandi Gec. QoS-Aware orchestration of network intensive software utilities within software defined data centres. Journal of Grid Computing, 16(1):85–112, March 2018. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic).

URL https://link.springer.com/
article/10.1007/s10723-017-9415-
1.

Guerrero:2018:GAM

[467] Carlos Guerrero, Isaac Lera, and Carlos Juiz. Genetic algorithm for multiobjective optimization of container allocation in cloud architecture. Journal of Grid Computing, 16(1):113–135, March 2018. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic).

URL https://link.springer.com/

article/10.1007/s10723-017-9419-
x.

Ramon-Cortes:2018:TOT

[468] Cristian Ramon-Cortes, Albert Serven, Jorge Ejarque, Daniele Lezzi, and Rosa M. Badia. Transparent orchestration of task-based parallel applications in containers platforms. Journal of Grid Computing, 16(1):137–160, March 2018. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic).

URL https://link.springer.com/
article/10.1007/s10723-017-9425-
z.

Stankovski:2018:GEI

[469] Vlado Stankovski and Radu Prodan. Guest Editors’ introduction: Special issue on storage for the big data era. Journal of Grid Computing, 16(2):161–163, June 2018. CO-DEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL https:

//link.springer.com/article/10.
1007/s10723-018-9439-1;   https:
//link.springer.com/content/pdf/
10.1007/s10723-018-9439-1.pdf.

Rafique:2018:PPB

[470] Ansar Rafique, Dimitri Van Landuyt, and Wouter Joosen. PERSIST: Policybased data management middleware for multi-tenant SaaS leveraging federated cloud storage. Journal of Grid Computing, 16(2):165–194, June 2018. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic).

URL https://link.springer.com/
article/10.1007/s10723-018-9434-

Zhang:2018:DSD

[471] Binqi Zhang, Chen Wang, Bing Bing Zhou, Dong Yuan, and Albert Y. Zomaya. DCDedupe: Selective deduplication and delta compression with effective routing for distributed storage. Journal of Grid Computing, 16 (2):195–209, June 2018. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (elec-

                                        tronic). URL https://link.springer.
                                        com/article/10.1007/s10723-018-
                                        9429-3.

Meoni:2018:DPP

[472] Marco Meoni, Raffaele Perego, and Nicola Tonellotto. Dataset popularity prediction for caching of CMS big data. Journal of Grid Computing, 16 (2):211–228, June 2018. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (elec-

                                        tronic). URL https://link.springer.
                                        com/article/10.1007/s10723-018-
                                        9436-4.

Sui:2018:PPC

[473] Peipei Sui and Xiaoyu Yang. A privacypreserving compression storage method for large trajectory data in road network. Journal of Grid Computing, 16 (2):229–245, June 2018. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (elec-

                                        tronic). URL https://link.springer.
                                        com/article/10.1007/s10723-018-
                                        9435-5.

Hajnal:2018:EVI

[474] Akos Hajnal, Gabor Kecskemeti, Attila Csaba Marosi, Jozsef Kovacs, Peter Kacsuk, and Robert Lovas. EN-TICE VM image analysis and optimised fragmentation. Journal of Grid Computing, 16(2):247–263, June 2018. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic).

URL https://link.springer.com/
article/10.1007/s10723-018-9430-
x.

Guerrero:2018:MAG

[475] Carlos Guerrero, Isaac Lera, and Carlos Juiz. Migration-aware genetic optimization for MapReduce scheduling and replica placement in Hadoop. Journal of Grid Computing, 16(2):265–284, June 2018. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic).

URL https://link.springer.com/
article/10.1007/s10723-018-9432-
8.

Tran:2018:NDL

[476] Xuan T. Tran, Tien Van Do, Csaba Rotter, and Dosam Hwang. A new data layout scheme for energy-efficient MapReduce processing tasks. Journal of Grid Computing, 16(2):285–298, June 2018. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic).

URL https://link.springer.com/
article/10.1007/s10723-018-9433-
7.

Liao:2018:BSS

[477] Jianwei Liao, Dong Yin, and Xiaoning Peng. Block I/O scheduling on storage servers of distributed file systems. Journal of Grid Computing, 16(2):299–316, June 2018. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic).

URL https://link.springer.com/
article/10.1007/s10723-017-9423-
1.

Djemaiel:2018:NGB

[478] Yacine Djemaiel, Sarra Berrahal, and Noureddine Boudriga. A novel graphbased approach for the management of health data on cloud-based WSANs. Journal of Grid Computing, 16(2): 317–344, June 2018. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (elec-

tronic). URL https://link.springer.
com/article/10.1007/s10723-018-
9438-2.

Sahal:2018:IIB

[479] Radhya Sahal, Marwah Nihad, Mohamed H. Khafagy, and Fatma A. Omara. iHOME: Index-based JOIN query optimization for limited big data storage. Journal of Grid Computing, 16 (2):345–380, June 2018. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (elec-

tronic). URL https://link.springer.
com/article/10.1007/s10723-018-
9431-9.

Salomoni:2018:IDP

[480] D. Salomoni, I. Campos, L. Gaido, J. Marco de Lucas, P. Solagna, J. Gomes, L. Matyska, P. Fuhrman, M. Hardt, G. Donvito, L. Dutka, M. Plociennik, R. Barbera, I. Blanquer, A. Ceccanti, and et al. INDIGO-DataCloud: a platform to facilitate seamless access to e-infrastructures. Journal of Grid Computing, 16(3): 381–408, September 2018. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL https:

//link.springer.com/article/10.
1007/s10723-018-9453-3;   https:
//link.springer.com/content/pdf/
10.1007/s10723-018-9453-3.pdf.

Song:2018:MBD

[481] Jie Song, HongYan He, Zhi Wang, Ge Yu, and Jean-Marc Pierson. Modulo based data placement algorithm for energy consumption optimization of MapReduce system. Journal of Grid Computing, 16(3):409–424, September 2018. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic).

URL https://link.springer.com/
article/10.1007/s10723-016-9370-
2.

Aazam:2018:TMI

[482] Mohammad Aazam, Eui-Nam Huh, and Marc St-Hilaire. Towards media intercloud standardization — evaluating impact of cloud storage heterogeneity. Journal of Grid Computing, 16(3):425– 443, September 2018. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (elec-

tronic). URL https://link.springer.
com/article/10.1007/s10723-015-
9356-5.

Shmueli:2018:FSF

[483] Edi Shmueli and Ilya Zaides. Framework for scalable file system metadata crawling and differencing. Journal of Grid Computing, 16(3):445–457, September 2018. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic).

URL https://link.springer.com/
article/10.1007/s10723-017-9400-
8.

Zhang:2018:EET

[484] Yi Zhang, Xiaohui Cheng, Liuhua Chen, and Haiying Shen. Energy-efficient tasks scheduling heuristics with multiconstraints in virtualized clouds. Journal of Grid Computing, 16(3):459– 475, September 2018. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (elec-

tronic). URL https://link.springer.
com/article/10.1007/s10723-018-
9426-6.

Aryania:2018:EAV

[485] Azra Aryania, Hadi S. Aghdasi, and Leyli Mohammad Khanli. Energyaware virtual machine consolidation algorithm based on ant colony system. Journal of Grid Computing, 16(3):477– 491, September 2018. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (elec-

                                        tronic). URL https://link.springer.
                                        com/article/10.1007/s10723-018-
                                        9428-4.

Hinz:2018:CMI

[486] Mauro Hinz, Guilherme Piegas Koslovski, Charles C. Miers, Laércio L. Pilla, and Maurício A. Pillon. A cost model for IaaS clouds based on virtual machine energy consumption. Journal of Grid Computing, 16(3):493–512, September 2018. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s10723-018-9440-
                                        8.

Pascual:2018:ERV

[487] Jose A. Pascual, Jose A. Lozano, and Jose Miguel-Alonso. Effects of reducing VMs management times on elastic applications. Journal of Grid Computing, 16(3):513–530, September 2018. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s10723-018-9441-
                                        7.

Pereira:2018:SIB

[488] José Luís Pereira, Orlando Belo, and Pascal Ravesteijn. Special issue on big data and digital transformation. Journal of Grid Computing, 16 (4):531–533, December 2018. CO-DEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL https:

//link.springer.com/article/10.
1007/s10723-018-9469-8;   https:
//link.springer.com/content/pdf/
10.1007/s10723-018-9469-8.pdf.

Yousefi:2018:TBG

[489] Mostafa Hadadian Nejad Yousefi and Maziar Goudarzi. A task-based greedy scheduling algorithm for minimizing energy of MapReduce jobs. Journal of Grid Computing, 16(4):535–551, December 2018. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic).

URL https://link.springer.com/
article/10.1007/s10723-018-9464-
0.

Gupta:2018:CTA

[490] Preeti Gupta, Rajni Jindal, and Arun Sharma. Community trolling: An active learning approach for topic based community detection in big data. Journal of Grid Computing, 16(4):553–567, December 2018. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic).

URL https://link.springer.com/
article/10.1007/s10723-018-9457-
z.

Fu:2018:IMR

[491] Xiaoyuan Fu, Jingyu Wang, Qi Qi, Jianxin Liao, and Tonghong Li. Incentive mechanisms for resource scalingout game of stream big data analytics. Journal of Grid Computing, 16(4): 569–585, December 2018. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (elec-

tronic). URL https://link.springer.
com/article/10.1007/s10723-018-
9458-y.

Enes:2018:BDO

[492] Jonatan Enes, Javier López Cacheiro, Roberto R. Expósito, and Juan Touriño. Big data-oriented PaaS architecture with disk-as-a-resource capability and container-based virtualization. Journal of Grid Computing, 16(4):587–605, December 2018. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s10723-018-9460-
                                        4.

Abdel-Hamid:2018:DSB

[493] Nahla B. Abdel-Hamid, Sally El-Ghamrawy, Ali El Desouky, and Hesham Arafat. A dynamic Sparkbased classification framework for imbalanced big data. Journal of Grid Computing, 16(4):607–626, December 2018. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s10723-018-9465-
                                        z.

Lordan:2018:TMC

[494] F. Lordan, J. Jensen, and R. M. Badia. Towards mobile cloud computing with single sign-on access. Journal of Grid Computing, 16(4):627–646, December 2018. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s10723-017-9413-
                                        3.

Fang:2018:PTS

[495] Wenjie Fang and Uwe Beckert. Parallel tree search in volunteer computing: a case study. Journal of Grid Computing, 16(4):647–662, December 2018. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL https:

//link.springer.com/article/10.
1007/s10723-017-9411-5;   https:
//link.springer.com/content/pdf/
10.1007/s10723-017-9411-5.pdf.

Memon:2018:TFS

[496] Shiraz Memon, Jensen Jens, Elbers Willem, Helmut Neukirchen, Matthias Book, and Morris Riedel. Towards federated service discovery and identity management in collaborative data and compute cloud infrastructures. Journal of Grid Computing, 16(4):663–681, December 2018. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic).

URL https://link.springer.com/
article/10.1007/s10723-018-9445-
3.

Fernandez:2018:UVA

[497] Pablo Orviz Fernández, João Pina, Álvaro López García, Isabel Campos Plasencia, Mário David, and Jorge Gomes. umd-verification: Automation of software validation for the EGI federated e-infrastructure. Journal of Grid Computing, 16(4):683–696, December 2018. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic).

URL https://link.springer.com/
article/10.1007/s10723-018-9454-
2.

Ricci:2019:E

[498] Laura Ricci, Alexandru Iosup, and Radu Prodan. Editorial. Journal of Grid Computing, 17(1):1–2, March 2019. CO-DEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL https:

//link.springer.com/article/10.
1007/s10723-019-09480-4;   https:
//link.springer.com/content/pdf/
10.1007/s10723-019-09480-4.pdf.

Bartoletti:2019:JBM

[499] Massimo Bartoletti, Bryn Bellomy, and Livio Pompianu. A journey into Bitcoin metadata. Journal of Grid Computing, 17(1):3–22, March 2019. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s10723-019-09473-
                                        3.

Guidi:2019:TDC

[500] Barbara Guidi, Andrea Michienzi, and Giulio Rossetti. Towards the dynamic community discovery in decentralized online social networks. Journal of Grid Computing, 17(1):23–44, March 2019. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s10723-018-9448-
                                        0.

Carlini:2019:AMF

[501] Emanuele Carlini and Alessandro Lulli. Analysis of movement features in multiplayer online battle arenas. Journal of Grid Computing, 17(1):45–57, March 2019. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s10723-018-9470-
                                        2.

Kavalionak:2019:DVS

[502] Hanna Kavalionak, Claudio Gennaro, Giuseppe Amato, Claudio Vairo, Costantino Perciante, Carlo Meghini, and Fabrizio Falchi. Distributed video surveillance using smart cameras. Journal of Grid Computing, 17(1):59–77, March 2019. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic).

URL https://link.springer.com/
article/10.1007/s10723-018-9467-
x.

Cai:2019:SSD

[503] Zhiming Cai, Ivan Lee, Shu-Chuan Chu, and Xuehong Huang. SimSim: A service discovery method preserving content similarity and spatial similarity in P2P mobile cloud. Journal of Grid Computing, 17(1):79–95, March 2019. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic).

URL https://link.springer.com/
article/10.1007/s10723-019-09475-
1.

Bistarelli:2019:EEV

[504] Stefano Bistarelli, Ivan Mercanti, Paolo Santancini, and Francesco Santini. Endto-end voting with non-permissioned and permissioned ledgers. Journal of Grid Computing, 17(1):97–118, March 2019. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic).

URL https://link.springer.com/
article/10.1007/s10723-019-09478-
y.

Krasovec:2019:EGC

[505] Barbara Krasovec and Andrej Filipcic.

Enhancing the grid with cloud computing. Journal of Grid Computing, 17(1): 119–135, March 2019. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (elec-

tronic). URL https://link.springer.
com/article/10.1007/s10723-018-
09472-w.

Park:2019:CRA

[506] Joonseok Park, Ungsoo Kim, Donggyu Yun, and Keunhyuk Yeom. C-RCE: an approach for constructing and managing a cloud service broker. Journal of Grid Computing, 17(1):137–168, March 2019. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic).

URL https://link.springer.com/
article/10.1007/s10723-017-9422-
2.

Selimi:2019:LSP

[507] Mennan Selimi, Llorenç Cerdà-Alabern, Felix Freitag, Luís Veiga, Arjuna Sathiaseelan, and Jon Crowcroft. A lightweight service placement approach for community network microclouds. Journal of Grid Computing, 17(1):169–189, March 2019. CO-DEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL https:

//link.springer.com/article/10.
1007/s10723-018-9437-3;   https:
//link.springer.com/content/pdf/
10.1007/s10723-018-9437-3.pdf.

deAlfonso:2019:MED

[508] Carlos de Alfonso, Miguel Caballer, Amanda Calatrava, Germán Moltó, and Ignacio Blanquer. Multi-elastic datacenters: Auto-scaled virtual clusters on energy-aware physical infrastructures. Journal of Grid Computing, 17 (1):191–204, March 2019. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (elec-

tronic). URL https://link.springer.
com/article/10.1007/s10723-018-
9449-z.

Anwar:2019:SIK

[509] Sajid Anwar and Álvaro Rocha. Special issue on knowledge discovery in big data (KDBD). Journal of Grid Computing, 17(2):205–208, June 2019. CO-DEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL https:

//link.springer.com/article/10.
1007/s10723-019-09484-0;   https:

//link.springer.com/content/pdf/
10.1007/s10723-019-09484-0.pdf.

Ullah:2019:DPQ

[510] Abrar Ullah, Hannan Xiao, and Trevor Barker. A dynamic profile questions approach to mitigate impersonation in online examinations. Journal of Grid Computing, 17(2):209–223, June 2019. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL https:

//link.springer.com/article/10.
1007/s10723-018-9442-6;   https:
//link.springer.com/content/pdf/
10.1007/s10723-018-9442-6.pdf.

Nauman:2019:BHT

[511] Mohammad Nauman, Hafeez Ur Rehman, Gianfranco Politano, and Alfredo Benso. Beyond homology transfer: Deep learning for automated annotation of proteins. Journal of Grid Computing, 17(2):225–237, June 2019. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic).

URL https://link.springer.com/
article/10.1007/s10723-018-9450-
6.

Khan:2019:OGF

[512] Salabat Khan, Amir Khan, Muazzam Maqsood, Farhan Aadil, and Mustansar Ali Ghazanfar. Optimized Gabor feature extraction for mass classification using cuckoo search for big data e-healthcare. Journal of Grid Computing, 17(2):239–254, June 2019. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic).

URL https://link.springer.com/
article/10.1007/s10723-018-9459-
x.

Badshah:2019:IMD

[513] Jan Badshah, Muhammad Kamran, Nadir Shah, and Shahbaz Akhtar Abid. An improved method to deploy cache servers in software defined network-based information centric networking for big data. Journal of Grid Computing, 17(2):255–277, June 2019. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s10723-019-09477-
                                        z.

Saeed:2019:WHA

[514] Zafar Saeed, Rabeeh Ayaz Abbasi, Onaiza Maqbool, Abida Sadaf, Imran Razzak, Ali Daud, Naif Radi Aljohani, and Guandong Xu. What’s happening around the world? A survey and framework on event detection techniques on Twitter. Journal of Grid Computing, 17(2):279–312, June 2019. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s10723-019-09482-
                                        2.

Majid:2019:APV

[515] Abdul Majid, Mukhtaj Khan, Nadeem Iqbal, Mian Ahmad Jan, Mushtaq Khan, and Salman. Application of parallel vector space model for large-scale DNA sequence analysis. Journal of Grid Computing, 17(2):313–324, June 2019. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s10723-018-9451-

Gupta:2019:ETU

[516] Akash Gupta, Harsh Sahu, Nihal Nanecha, Pradeep Kumar, Partha Pratim Roy, and Victor Chang. Enhancing text using emotion detected from EEG signals. Journal of Grid Computing, 17(2):325–340, June 2019. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic).

URL https://link.springer.com/
article/10.1007/s10723-018-9462-
2.

Kim:2019:ERT

[517] Beom-Su Kim, Monther Aldwairi, and Ki-Il Kim. An efficient real-time data dissemination multicast protocol for big data in wireless sensor networks. Journal of Grid Computing, 17(2):341–355, June 2019. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic).

URL https://link.springer.com/
article/10.1007/s10723-018-9447-
1.

Alam:2019:BDA

[518] Saqib Alam and Nianmin Yao. Big data analytics, text mining and modern English language. Journal of Grid Computing, 17(2):357–366, June 2019. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic).

URL https://link.springer.com/
article/10.1007/s10723-018-9452-
4.

Raza:2019:SOB

[519] Arslan Ali Raza, Asad Habib, Jawad Ashraf, and Muhammad Javed. Semantic orientation based decision making framework for big data analysis of sporadic news events. Journal of Grid Computing, 17(2):367–383, June 2019. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic).

URL https://link.springer.com/
article/10.1007/s10723-018-9466-
y.

Gill:2019:RPB

[520] Sukhpal Singh Gill and Rajkumar Buyya. Resource provisioning based scheduling framework for execution of heterogeneous and clustered workloads in clouds: from fundamental to autonomic offering. Journal of Grid Computing, 17(3):385–417, September 2019. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic).

URL https://link.springer.com/
article/10.1007/s10723-017-9424-
0.

Chunlin:2019:HCA

[521] Li Chunlin, Tang Jianhang, and Luo Youlong. Hybrid cloud adaptive scheduling strategy for heterogeneous workloads. Journal of Grid Computing, 17(3):419–446, September 2019. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic).

URL https://link.springer.com/
article/10.1007/s10723-019-09481-
3.

deSouza:2019:QST

[522] Felipe Rodrigo de Souza, Charles Christian Miers, Adriano Fiorese, Marcos Dias de Assunção, and Guilherme Piegas Koslovski. QVIA-SDN: Towards QoS-aware virtual infrastructure allocation on SDN-based clouds. Journal of Grid Computing, 17(3):447– 472, September 2019. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (elec-

tronic). URL https://link.springer.

com/article/10.1007/s10723-019-
09479-x.

Righi:2019:SGM

[523] Rodrigo da Rosa Righi, Matheus Lehmann, Marcio Miguel Gomes, Jeferson Campos Nobre, Cristiano André da Costa, Sandro José Rigo, Marcio Lena, Rodrigo Fraga Mohr, and Luiz Ricardo Bertoldi de Oliveira. A survey on global management view: Toward combining system monitoring, resource management, and load prediction. Journal of Grid Computing, 17(3):473–502, September 2019. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic).

URL https://link.springer.com/
article/10.1007/s10723-018-09471-
x.

Dragan:2019:SPM

[524] Ioan Dragan, Gabriel Iuhasz, and Dana Petcu. A scalable platform for monitoring data intensive applications. Journal of Grid Computing, 17(3):503– 528, September 2019. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (elec-

tronic). URL https://link.springer.
com/article/10.1007/s10723-019-
09483-1.

Kertesz:2019:MID

[525] A. Kertesz, T. Pflanzner, and T. Gyimothy. A mobile IoT device simulator for IoT–Fog-Cloud systems. Journal of Grid Computing, 17(3):529– 551, September 2019. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (elec-

tronic). URL https://link.springer.
com/article/10.1007/s10723-018-
9468-9.

Hsieh:2019:ILB

[526] Hui-Ching Hsieh and Mao-Lun Chiang. The incremental load balance cloud algorithm by using dynamic data deployment. Journal of Grid Computing, 17(3):553–575, September 2019. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s10723-019-09474-
                                        2.

Khan:2019:IDS

[527] Fakhri Alam Khan, Mujeeb ur Rehman, Afsheen Khalid, Muhammad Ali, Muhammad Imran, Muhammad Nawaz, and Attaur Rahman. An intelligent data service framework for heterogeneous data sources. Journal of Grid Computing, 17(3):577–589, September 2019. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s10723-018-9443-
                                        5.

Huancheng:2019:ART

[528] Liu Huancheng, Wu Tingting, and Álvaro Rocha. An analysis of research trends on data mining in Chinese academic libraries. Journal of Grid Computing, 17(3):591–601, September 2019. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s10723-018-9461-
                                        3.

Andrade:2019:PEC

[529] Ermeson Andrade and Bruno Nogueira. Performability evaluation of a cloudbased disaster recovery solution for IT environments. Journal of Grid Computing, 17(3):603–621, September 2019. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic).

URL https://link.springer.com/
article/10.1007/s10723-018-9446-
2.

Pahl:2019:SII

[530] Claus Pahl, Muthu Ramachandran, and Gary Wills. Special issue: Intelligent management of cloud, IoT and big data applications. Journal of Grid Computing, 17(4):623–624, December 2019. CO-DEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL https:

//link.springer.com/article/10.
1007/s10723-019-09496-w;   https:
//link.springer.com/content/pdf/
10.1007/s10723-019-09496-w.pdf.

Kotb:2019:CBM

[531] Yehia Kotb, Ismaeel Al Ridhawi, Moayad Aloqaily, Thar Baker, Yaser Jararweh, and Hissam Tawfik. Cloudbased multi-agent cooperation for IoT devices using workflow-nets. Journal of Grid Computing, 17(4):625–650, December 2019. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic).

URL https://link.springer.com/
article/10.1007/s10723-019-09485-
z.

Baraki:2019:OAM

[532] Harun Baraki, Alexander Jahl, Stefan Jakob, Corvin Schwarzbach, Malte Fax, and Kurt Geihs. Optimizing applications for mobile cloud computing through MOCCAA. Journal of Grid Computing, 17(4):651–676, December 2019. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic).

URL https://link.springer.com/

                                        article/10.1007/s10723-019-09492-
                                        0.

Aral:2019:AAL

[533] Atakan Aral, Ivona Brandic, Rafael Brundo Uriarte, Rocco De Nicola, and Vincenzo Scoca. Addressing application latency requirements through edge scheduling. Journal of Grid Computing, 17 (4):677–698, December 2019. CO-DEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s10723-019-09493-z;   https:
                                        //link.springer.com/content/pdf/
                                        10.1007/s10723-019-09493-z.pdf.

Lin:2019:SAH

[534] Weiwei Lin, Gaofeng Peng, Xinran Bian, Siyao Xu, Victor Chang, and Yin Li. Scheduling algorithms for heterogeneous cloud environment: Main resource load balancing algorithm and time balancing algorithm. Journal of Grid Computing, 17(4):699–726, December 2019. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s10723-019-09499-
                                        7.

Schomakers:2019:TOP

[535] Eva-Maria Schomakers, Chantal Lidynia, and Martina Ziefle. A typology of online privacy personalities. Journal of Grid Computing, 17(4):727–747, December 2019. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s10723-019-09500-
                                        3.

Oliveira:2019:PEO

[536] Danilo Oliveira, André Brinkmann, Nelson Rosa, and Paulo Maciel. Performability evaluation and optimization of workflow applications in cloud environments. Journal of Grid Computing, 17(4):749–770, December 2019. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic).

URL https://link.springer.com/
article/10.1007/s10723-019-09476-
0.

Salvadore:2019:LWB

[537] F. Salvadore and R. Ponzini. LincoSim: a Web based HPC–Cloud platform for automatic virtual towing tank analysis. Journal of Grid Computing, 17(4):771–795, December 2019. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic).

URL https://link.springer.com/
article/10.1007/s10723-019-09494-
y.

Sniezynski:2019:VRP

[538] Bartlomiej Sniezynski, Piotr Nawrocki, Michal Wilk, Marcin Jarzab, and Krzysztof Zielinski. VM reservation plan adaptation using machine learning in cloud computing. Journal of Grid Computing, 17(4):797–812, December 2019. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL https:

//link.springer.com/article/10.
1007/s10723-019-09487-x;   https:
//link.springer.com/content/pdf/
10.1007/s10723-019-09487-x.pdf.

Kovacs:2019:SPA

[539] József Kovács. Supporting programmable autoscaling rules for containers and virtual machines on clouds.

Journal of Grid Computing, 17(4): 813–829, December 2019. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL https:

//link.springer.com/article/10.
1007/s10723-019-09488-w;   https:
//link.springer.com/content/pdf/
10.1007/s10723-019-09488-w.pdf.

Shah:2019:EMF

[540] Ahsan Shah and Zahid Halim. On efficient mining of frequent itemsets from big uncertain databases. Journal of Grid Computing, 17(4):831–850, December 2019. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic).

URL https://link.springer.com/
article/10.1007/s10723-018-9456-
0.

Ghobaei-Arani:2020:RMA

[541] Mostafa Ghobaei-Arani, Alireza Souri, and Ali A. Rahmanian. Resource management approaches in fog computing: a comprehensive review. Journal of Grid Computing, 18(1):1–42, March 2020. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic).

URL https://link.springer.com/
article/10.1007/s10723-019-09491-
1.

Abbasi:2020:WAI

[542] Mahdi Abbasi, Ehsan Mohammadi Pasand, and Mohammad R. Khosravi. Workload allocation in IoT-Fog-Cloud architecture using a multiobjective genetic algorithm. Journal of Grid Computing, 18(1):43–56, March 2020. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic).

URL https://link.springer.com/
article/10.1007/s10723-020-09507-
1.

Javed:2020:IFE

[543] Asad Javed, Jérémy Robert, Keijo Heljanko, and Kary Främling. IoTEF: A federated edge-cloud architecture for fault-tolerant IoT applications. Journal of Grid Computing, 18(1):57–80, March 2020. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (elec-

tronic). URL https://link.springer.
com/article/10.1007/s10723-019-
09498-8;   https://link.springer.
com/content/pdf/10.1007/s10723-
019-09498-8.pdf.

Fernandez:2020:SQA

[544] Pablo Orviz Fernández, Mário David, Doina Cristina Duma, Elisabetta Ronchieri, Jorge Gomes, and Davide Salomoni. Software quality assurance in INDIGO–DataCloud Project: a converging evolution of software engineering practices to support European research e-infrastructures. Journal of Grid Computing, 18(1):81–98, March 2020. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic).

URL https://link.springer.com/
article/10.1007/s10723-020-09509-
z.

Anderson:2020:BPV

[545] David P. Anderson. BOINC: A platform for volunteer computing. Journal of Grid Computing, 18(1):99–122, March 2020. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (elec-

tronic). URL https://link.springer.
com/article/10.1007/s10723-019-
09497-9;   https://link.springer.
com/content/pdf/10.1007/s10723-
019-09497-9.pdf.

Kong:2020:HLB

[546] Lingfu Kong, Jean Pepe Buanga Mapetu, and Zhen Chen. Heuristic load balancing based zero imbalance mechanism in cloud computing. Journal of Grid Computing, 18(1):123–148, March 2020. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s10723-019-09486-
                                        y.

Zharikov:2020:AWF

[547] Eduard Zharikov, Sergii Telenyk, and Petro Bidyuk. Adaptive workload forecasting in cloud data centers. Journal of Grid Computing, 18(1):149–168, March 2020. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s10723-019-09501-
                                        2.

Zia:2020:RRH

[548] Tehseen Zia and Saad Razzaq. Residual recurrent highway networks for learning deep sequence prediction models. Journal of Grid Computing, 18(1):169–176, March 2020. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s10723-018-9444-
                                        4.

Xia:2020:RPD

[549] Kaijian Xia, Tao Hu, and Wen Si. Research on parallel deep learning for heterogeneous computing architecture. Journal of Grid Computing, 18(2):177–179, June 2020. CO-DEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL https:

                                        //link.springer.com/article/10.

1007/s10723-020-09520-4;   https:
//link.springer.com/content/pdf/
10.1007/s10723-020-09520-4.pdf.

Zhong:2020:MLB

[550] Shan Zhong, Jack Tan, Husheng Dong, Xuemei Chen, Shengrong Gong, and Zhenjiang Qian. Modeling-learning-based actor-critic algorithm with Gaussian process approximator. Journal of Grid Computing, 18(2):181–195, June 2020. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic).

URL https://link.springer.com/
article/10.1007/s10723-020-09512-
4.

Ni:2020:MTD

[551] Tongguang Ni, Xiaoqing Gu, Cong Zhang, Weibo Wang, and Yiqing Fan. Multi-task deep metric learning with boundary discriminative information for cross-age face verification. Journal of Grid Computing, 18(2):197–210, June 2020. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic).

URL https://link.springer.com/
article/10.1007/s10723-019-09495-
x.

Qian:2020:ECM

[552] Pengjiang Qian, Ke Xu, Tingyu Wang, Qiankun Zheng, Huan Yang, Atallah Baydoun, Junqing Zhu, Bryan Traughber, and Raymond F. Muzic, Jr. Estimating CT from MR abdominal images using novel generative adversarial networks. Journal of Grid Computing, 18(2):211–226, June 2020. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic).

URL https://link.springer.com/
article/10.1007/s10723-020-09513-

Hu:2020:DIA

[553] Yan Hu. Design and implementation of abnormal behavior detection based on deep intelligent analysis algorithms in massive video surveillance. Journal of Grid Computing, 18(2):227–237, June 2020. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s10723-020-09506-
                                        2.

Lu:2020:IMC

[554] Weijia Lu. Improved K-means clustering algorithm for big data mining under Hadoop parallel framework. Journal of Grid Computing, 18(2):239–250, June 2020. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s10723-019-09503-
                                        0.

Yuan:2020:ADM

[555] Jingzhen Yuan. An anomaly data mining method for mass sensor networks using improved PSO algorithm based on spark parallel framework. Journal of Grid Computing, 18(2):251–261, June 2020. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s10723-020-09505-
                                        3.

Xia:2020:RPA

[556] Dongliang Xia, Feifei Ning, and Weina He. Research on parallel adaptive canopy-k-means clustering algorithm for big data mining based on cloud platform. Journal of Grid Computing, 18 (2):263–273, June 2020. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (elec-

tronic). URL https://link.springer.
com/article/10.1007/s10723-019-
09504-z.

Ping:2020:LBA

[557] Yang Ping. Load balancing algorithms for big data flow classification based on heterogeneous computing in software definition networks. Journal of Grid Computing, 18(2):275–291, June 2020. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic).

URL https://link.springer.com/
article/10.1007/s10723-020-09511-
5.

Darabian:2020:DCM

[558] Hamid Darabian, Sajad Homayounoot, Ali Dehghantanha, Sattar Hashemi, Hadis Karimipour, Reza M. Parizi, and Kim-Kwang Raymond Choo. Detecting cryptomining malware: a deep learning approach for static and dynamic analysis. Journal of Grid Computing, 18(2):293–303, June 2020. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic).

URL https://link.springer.com/
article/10.1007/s10723-020-09510-
6.

Abbasi:2020:RAP

[559] Mahdi Abbasi and Mohammad R. Khosravi. A robust and accurate particle filter-based pupil detection method for big datasets of eye video. Journal of Grid Computing, 18(2):305–325, June 2020. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic).

URL https://link.springer.com/
article/10.1007/s10723-019-09502-
1.

Hosseinzadeh:2020:MOT

[560] Mehdi Hosseinzadeh, Marwan Yassin Ghafour, and Afsane Khoshnevis. Multiobjective task and workflow scheduling approaches in cloud computing: a comprehensive review. Journal of Grid Computing, 18(3):327–356, September 2020. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s10723-020-09533-
                                        z.

Singh:2020:EEA

[561] Vishakha Singh, Indrajeet Gupta, and Prasanta K. Jana. An energy efficient algorithm for workflow scheduling in IaaS cloud. Journal of Grid Computing, 18(3):357–376, September 2020. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s10723-019-09490-
                                        2.

Matani:2020:FTW

[562] Alemeh Matani, Hamid Reza Naji, and Hassan Motallebi. A fault-tolerant workflow scheduling algorithm for grid with near-optimal redundancy. Journal of Grid Computing, 18(3):377– 394, September 2020. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (elec-

                                        tronic). URL https://link.springer.
                                        com/article/10.1007/s10723-020-
                                        09522-2.

Dazzi:2020:SDI

[563] Patrizio Dazzi and Matteo Mordacchini. Scalable decentralized indexing and querying of multi-streams in the fog. Journal of Grid Computing, 18(3): 395–418, September 2020. CODEN ????

ISSN 1570-7873 (print), 1572-9184 (elec-

tronic). URL https://link.springer.
com/article/10.1007/s10723-020-
09521-3.

Kirthica:2020:ECI

[564] S. Kirthica, I. Saravanan, and Rajeswari Sridhar. Enhancing the Cloud Inter-operation Toolkit (CIT) to support multiple cloud service models. Journal of Grid Computing, 18(3):419– 439, September 2020. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (elec-

tronic). URL https://link.springer.
com/article/10.1007/s10723-020-
09516-0.

Park:2020:ASI

[565] Joonseok Park, Ungsoo Kim, and Keunhyuk Yeom. Approach for selecting and integrating cloud services to construct hybrid cloud. Journal of Grid Computing, 18(3):441–469, September 2020. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic).

URL https://link.springer.com/
article/10.1007/s10723-020-09519-
x.

Meng:2020:NPF

[566] Zeqian Meng, John Brooke, and Rizos Sakellariou. A negotiation protocol for fine-grained accountable resource provisioning and sharing in e-science. Journal of Grid Computing, 18(3):471– 490, September 2020. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (elec-

tronic). URL https://link.springer.
com/article/10.1007/s10723-020-
09515-1.

Do:2020:PIO

[567] Nam H. Do, Tien Van Do, and Csaba Rotter. Provisioning input and output data rates in data processing frameworks. Journal of Grid Computing, 18(3):491–506, September 2020. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic).

URL https://link.springer.com/
article/10.1007/s10723-020-09508-
0.

Ghorbani:2020:DFU

[568] M. Ghorbani, S. Swift, and A. M. Payne. Design of a flexible, user friendly feature matrix generation system and its application on biomedical datasets. Journal of Grid Computing, 18(3):507–527, September 2020. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic).

URL https://link.springer.com/
article/10.1007/s10723-020-09518-
y.

Temelkovski:2020:BSG

[569] Damjan Temelkovski, Tamas Kiss, and Pamela Greenwell. Building science gateways for analysing molecular docking results using a generic framework and methodology. Journal of Grid Computing, 18(3):529–546, September 2020. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic).

URL https://link.springer.com/
article/10.1007/s10723-020-09529-
9.

Ali:2020:APS

[570] Toqeer Ali, Yasar Khan, and Sajid Anwar. An automated permission selection framework for Android platform. Journal of Grid Computing, 18(3):547– 561, September 2020. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (elec-

tronic). URL https://link.springer.

com/article/10.1007/s10723-018-
9455-1.

Sousa:2020:LAM

[571] Maria José Sousa and Álvaro Rocha.

Learning analytics measuring impacts on organisational performance. Journal of Grid Computing, 18(3):563– 571, September 2020. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (elec-

tronic). URL https://link.springer.
com/article/10.1007/s10723-018-
9463-1.

Lin:2020:SIB

[572] Yimin Lin. Special issue: Blockchain theories and applications. Journal of Grid Computing, 18(4):573, December 2020. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic).

URL https://link.springer.com/
article/10.1007/s10723-020-09538-
8.

Strehle:2020:DOR

[573] Elias Strehle and Fred Steinmetz. Dominating OP returns: The impact of Omni and Veriblock on Bitcoin. Journal of Grid Computing, 18(4):575–592, December 2020. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic).

URL https://link.springer.com/
article/10.1007/s10723-020-09537-
9.

Lei:2020:BBC

[574] Kai Lei, Junjie Fang, and Kuai Xu.

Blockchain-based cache poisoning security protection and privacy-aware access control in NDN vehicular edge computing networks. Journal of Grid Computing, 18(4):593–613, December 2020. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic).

URL https://link.springer.com/
article/10.1007/s10723-020-09531-
1.

Sodhro:2020:TBE

[575] Ali Hassan Sodhro, Sandeep Pirbhulal, and Luo Zongwei. Towards blockchainenabled security technique for industrial Internet of Things based decentralized applications. Journal of Grid Computing, 18(4):615–628, December 2020. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic).

URL https://link.springer.com/
article/10.1007/s10723-020-09527-
x.

Le:2020:ADS

[576] Gejun Le, Qifeng Gu, and Jianping Fan. AirCargoChain: A distributed and scalable data sharing approach of blockchain for air cargo. Journal of Grid Computing, 18(4):629–638, December 2020. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic).

URL https://link.springer.com/
article/10.1007/s10723-020-09535-
x.

Shakarami:2020:SCO

[577] Ali Shakarami, Mostafa Ghobaei-Arani, and Mehdi Hosseinzadeh. A survey on the computation offloading approaches in mobile edge/cloud computing environment: a stochasticbased perspective. Journal of Grid Computing, 18(4):639–671, December 2020. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic).

URL https://link.springer.com/
article/10.1007/s10723-020-09530-
2.

Kochovski:2020:SCS

[578] Petar Kochovski, Vlado Stankovski, and Seungwoo Kum. Smart contracts for service-level agreements in edgeto-cloud computing. Journal of Grid Computing, 18(4):673–690, December 2020. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic).

URL https://link.springer.com/
article/10.1007/s10723-020-09534-
y.

Martins:2020:BSC

[579] Horácio Martins, Filipe Araujo, and Paulo Rupino da Cunha. Benchmarking serverless computing platforms. Journal of Grid Computing, 18(4):691–709, December 2020. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic).

URL https://link.springer.com/
article/10.1007/s10723-020-09523-
1.

Wood:2020:RMD

[580] Bradley Wood, Brock Watling, and Akramul Azim. Remote method delegation: a platform for grid computing. Journal of Grid Computing, 18(4): 711–725, December 2020. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (elec-

tronic). URL https://link.springer.
com/article/10.1007/s10723-020-
09525-z.

Masdari:2020:GCC

[581] Mohammad Masdari and Mehran Zangakani. Green cloud computing using proactive virtual machine placement: Challenges and issues. Journal of Grid Computing, 18(4):727–759, December 2020. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic).

URL https://link.springer.com/
article/10.1007/s10723-019-09489-
9.

Pierantoni:2020:DPT

[582] Gabriele Pierantoni, Tamas Kiss, and Hai-Van Dang. Describing and processing topology and quality of service parameters of applications in the cloud. Journal of Grid Computing, 18(4):761– 778, December 2020. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (elec-

tronic). URL https://link.springer.
com/article/10.1007/s10723-020-
09524-0.

Kosinska:2020:AMF

[583] Joanna Kosińska and Krzysztof Zieliński. Autonomic management framework for cloud-native applications. Journal of Grid Computing, 18(4):779–796, December 2020. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic).

URL https://link.springer.com/
article/10.1007/s10723-020-09532-
0.

Loyola-Gonzalez:2020:RSC

[584] Octavio Loyola-González, Miguel Angel Medina-Pérez, and Kim-Kwang Raymond Choo. A review of supervised classification based on contrast patterns: Applications, trends, and challenges. Journal of Grid Computing, 18(4):797–845, December 2020. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic).

URL https://link.springer.com/
article/10.1007/s10723-020-09526-
y.

Tavallali:2020:STP

[585] Pooya Tavallali, Mehran Yazdi, and Mohammad R. Khosravi. A systematic training procedure for Viola–Jones face detector in heterogeneous computing architecture. Journal of Grid Computing, 18(4):847–862, December 2020. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic).

URL https://link.springer.com/
article/10.1007/s10723-020-09517-
z.

Ur-Rehman:2020:VMH

[586] Attiq Ur-Rehman, Iqbal Gondal, and Alireza Jolfaei. Vulnerability modelling for hybrid industrial control system networks. Journal of Grid Computing, 18(4):863–878, December 2020. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic).

URL https://link.springer.com/
article/10.1007/s10723-020-09528-
w.

Farhang:2020:RMS

[587] Mandana Farhang and Faramarz Safi-Esfahani. Recognizing MapReduce straggler tasks in big data infrastructures using artificial neural networks. Journal of Grid Computing, 18(4):879– 901, December 2020. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (elec-

tronic). URL https://link.springer.
com/article/10.1007/s10723-020-
09514-2.

Wei:2021:DMT

[588] Hao Wei, Joaquin Salvachua Rodriguez, and Octavio Nieto-Taladriz Garcia. Deployment management and topology discovery of microservice applications in the multicloud environment. Journal of Grid Computing, 19(1):??, March 2021. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic).

URL https://link.springer.com/

                                        article/10.1007/s10723-021-09539-
                                        1.

Huedo:2021:ODD

[589] Eduardo Huedo, Rubén S. Montero, and Ignacio M. Llorente. Opportunistic deployment of distributed edge clouds for latency-critical applications. Journal of Grid Computing, 19(1):??, March 2021. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s10723-021-09545-
                                        3.

Demichev:2021:BPE

[590] Andrey Demichev, Alexander Kryukov, and Nikolai Prikhod’ko. Business process engineering for data storing and processing in a collaborative distributed environment based on provenance metadata, smart contracts and blockchain technology. Journal of Grid Computing, 19(1):??, March 2021. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s10723-021-09544-
                                        4.

Caballer:2021:DEV

[591] Miguel Caballer, Marica Antonacci, and Germán Moltó. Deployment of elastic virtual hybrid clusters across cloud sites. Journal of Grid Computing, 19(1):??, March 2021. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s10723-021-09543-
                                        5.

Tomarchio:2021:TTB

[592] Orazio Tomarchio, Domenico Calcaterra, and Pietro Mazzaglia. TORCH: a TOSCA-based orchestrator of multicloud containerised applications. Journal of Grid Computing, 19(1):??, March 2021. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic).

URL https://link.springer.com/
article/10.1007/s10723-021-09549-
z.

Zamani:2021:FDM

[593] Ghazal Zamani and Olivia Das. Faultdetection managers: More may not be the merrier. Journal of Grid Computing, 19(1):??, March 2021. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (elec-

tronic). URL https://link.springer.
com/article/10.1007/s10723-021-
09546-2.

Tabrizchi:2021:ERB

[594] Hamed Tabrizchi and Marjan Kuchaki Rafsanjani. Energy refining balance with ant colony system for cloud placement machines. Journal of Grid Computing, 19(1):??, March 2021. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic).

URL https://link.springer.com/
article/10.1007/s10723-021-09547-
1.

Xu:2021:ESD

[595] Jie Xu, Jingyu Wang, and Di Yang. Effective scheduler for distributed DNN training based on MapReduce and GPU cluster. Journal of Grid Computing, 19(1):??, March 2021. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (elec-

tronic). URL https://link.springer.
com/article/10.1007/s10723-021-
09550-6.

Bento:2021:AAD

[596] Andre Bento, Jaime Correia, and Jorge Cardoso. Automated analysis of distributed tracing: Challenges and research directions. Journal of Grid Computing, 19(1):??, March 2021. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s10723-021-09551-
                                        5.

Lin:2021:BTB

[597] Jenn-Wei Lin, Joseph M. Arul, and Jia-Ting Kao. A bottom-up tree based storage approach for efficient IoT data analytics in cloud systems. Journal of Grid Computing, 19(1):??, March 2021. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s10723-021-09553-
                                        3.

Sulaiman:2021:ECB

[598] Muhammad Sulaiman, Zahid Halim, and Shanshan Tu. An evolutionary computing-based efficient hybrid task scheduling approach for heterogeneous computing environment. Journal of Grid Computing, 19(1):??, March 2021. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s10723-021-09552-
                                        4.

Wei:2021:HCR

[599] Wei Wei, Qi Wang, and Yang Liu. Highly complex resource scheduling for stochastic demands in heterogeneous clouds. Journal of Grid Computing, 19(1):??, March 2021. CODEN ????

ISSN 1570-7873 (print), 1572-9184 (elec-

tronic). URL https://link.springer.
com/article/10.1007/s10723-021-
09555-1.

Hagras:2021:GDB

[600] Tarek Hagras, Asmaa Atef, and Yousef B. Mahdy. Greening duplicationbased dependent-tasks scheduling on heterogeneous large-scale computing platforms. Journal of Grid Computing, 19(1):??, March 2021. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (elec-

tronic). URL https://link.springer.
com/article/10.1007/s10723-021-
09554-2.

Hosseinzadeh:2021:IBO

[601] Mehdi Hosseinzadeh, Mohammad Masdari, and Sarkhel H. Taher Karim. Improved butterfly optimization algorithm for data placement and scheduling in edge computing environments. Journal of Grid Computing, 19(2):??, June 2021. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic).

URL https://link.springer.com/
article/10.1007/s10723-021-09556-

0. See correction [615].

Wang:2021:HRC

[602] Kun Wang, Xiaofeng Wang, and Xuan Liu. A high reliable computing offloading strategy using deep reinforcement learning for IoVs in edge computing. Journal of Grid Computing, 19(2):??, June 2021. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic).

URL https://link.springer.com/
article/10.1007/s10723-021-09542-

Shirmarz:2021:ASD

[603] Alireza Shirmarz and Ali Ghaffari. Automatic Software Defined Network (SDN) performance management using TOPSIS decision-making algorithm. Journal of Grid Computing, 19(2):??, June 2021. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s10723-021-09557-
                                        z.

Wang:2021:CRA

[604] Jianxi Wang and Liutao Wang. A computing resource allocation optimization strategy for massive Internet of Health Things devices considering privacy protection in cloud edge computing environment. Journal of Grid Computing, 19(2):??, June 2021. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s10723-021-09558-
                                        y.

Shahidinejad:2021:CAM

[605] Ali Shahidinejad, Fariba Farahbakhsh, and Toni Anwar. Context-aware multiuser offloading in mobile edge computing: a federated learning-based approach. Journal of Grid Computing, 19(2):??, June 2021. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (elec-

                                        tronic). URL https://link.springer.
                                        com/article/10.1007/s10723-021-
                                        09559-x.

Tarafdar:2021:EMA

[606] Anurina Tarafdar, Mukta Debnath, and Rajib K. Das. Energy and makespan aware scheduling of deadline sensitive tasks in the cloud environment. Journal of Grid Computing, 19(2):??, June 2021. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic).

URL https://link.springer.com/
article/10.1007/s10723-021-09548-
0.

Nawrocki:2021:CRD

[607] Piotr Nawrocki and Patryk Osypanka.

Cloud resource demand prediction using machine learning in the context of QoS parameters. Journal of Grid Computing, 19(2):??, June 2021. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic).

URL https://link.springer.com/
article/10.1007/s10723-021-09561-
3.

Princess:2021:HMH

[608] G. Annie Poornima Princess and A. S.

Radhamani. A hybrid meta-heuristic for optimal load balancing in cloud computing. Journal of Grid Computing, 19(2): ??, June 2021. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic).

URL https://link.springer.com/
article/10.1007/s10723-021-09560-
4.

Liu:2021:SEM

[609] Dong Liu, Longxi Chen, Zhiyong Wang, and Guangqiang Diao. Speech expression multimodal emotion recognition based on deep belief network. Journal of Grid Computing, 19(2):??, June 2021. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic).

URL https://link.springer.com/
article/10.1007/s10723-021-09564-

Besharati:2021:ICO

[610] Reza Besharati, Mohammad Hossein Rezvani, and Mohammad Mehdi Gilanian Sadeghi. An incentive-compatible offloading mechanism in fog-cloud environments using second-price sealed-bid auction. Journal of Grid Computing, 19 (3):??, September 2021. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (elec-

                                        tronic). URL https://link.springer.
                                        com/article/10.1007/s10723-021-
                                        09576-w.

Casani:2021:RLD

[611] Álvaro Fernández Casaní, Juan M. Orduña, and Santiago González de la Hoz. A reliable large distributed object store based platform for collecting event metadata. Journal of Grid Computing, 19(3):??, September 2021. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s10723-021-09580-
                                        0.

Chitanov:2021:PAP

[612] Vasiliy Chitanov, Michael Krieger, and Roumen Kakanakov. A pilot applied physics grid computing infrastructure for developing applications predicting the qualities of industrial coatings. Journal of Grid Computing, 19 (3):??, September 2021. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (elec-

                                        tronic). URL https://link.springer.
                                        com/article/10.1007/s10723-021-
                                        09566-y.

Foschini:2021:EEM

[613] Luca Foschini, Giuseppe Martuscelli, and Michele Solimando. Edge-enabled mobile crowdsensing to support effective rewarding for data collection in pandemic events. Journal of Grid Computing, 19(3):??, September 2021. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic).

URL https://link.springer.com/
article/10.1007/s10723-021-09569-
9.

Hedhli:2021:SSP

[614] Ameni Hedhli and Haithem Mezni.

A survey of service placement in cloud environments. Journal of Grid Computing, 19(3):??, September 2021. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic).

URL https://link.springer.com/
article/10.1007/s10723-021-09565-
z.

Hosseinzadeh:2021:CIB

[615] Mehdi Hosseinzadeh, Mohammad Masdari, and Sarkhel H. Taher Karim. Correction to: Improved Butterfly Optimization Algorithm for Data Placement and Scheduling in Edge Computing Environments. Journal of Grid Computing, 19(3):??, September 2021. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic).

URL https://link.springer.com/
article/10.1007/s10723-021-09562-

2. See [601].

Kumara:2021:SOA

[616] Indika Kumara, Paul Mundt, and Georgios Meditskos. SODALITE@RT: Orchestrating applications on cloudedge infrastructures. Journal of Grid Computing, 19(3):??, September 2021. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s10723-021-09572-
                                        0.

Li:2021:COR

[617] Xuezhu Li. A computing offloading resource allocation scheme using deep reinforcement learning in mobile edge computing systems. Journal of Grid Computing, 19(3):??, September 2021. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s10723-021-09568-
                                        w.

Muhammad:2021:BSB

[618] Asif Muhammad and Muhammad Aleem. BAN-Storm: a bandwidthaware scheduling mechanism for stream jobs. Journal of Grid Computing, 19 (3):??, September 2021. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (elec-

                                        tronic). URL https://link.springer.
                                        com/article/10.1007/s10723-021-
                                        09567-x.

Naas:2021:IDR

[619] Mohammed Islam Naas, Laurent Lemarchand, and Jalil Boukhobza. IoT data replication and consistency management in fog computing. Journal of Grid Computing, 19(3):??, September 2021. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s10723-021-09571-
                                        1.

Pires:2021:DDO

[620] André Pires, José Simão, and Luís Veiga. Distributed and decentralized orchestration of containers on edge clouds. Journal of Grid Computing, 19 (3):??, September 2021. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (elec-

tronic). URL https://link.springer.
com/article/10.1007/s10723-021-
09575-x.

Risco:2021:SWC

[621] Sebastián Risco, Germán Moltó, and Ignacio Blanquer. Serverless workflows for containerised applications in the cloud continuum. Journal of Grid Computing, 19(3):??, September 2021. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic).

URL https://link.springer.com/
article/10.1007/s10723-021-09570-
2.

Siar:2021:OCF

[622] Hajar Siar and Mohammad Izadi.

Offloading coalition formation for scheduling scientific workflow ensembles in fog environments. Journal of Grid Computing, 19(3):??, September 2021. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic).

URL https://link.springer.com/
article/10.1007/s10723-021-09574-
y.

Toka:2021:URL

[623] László Toka. Ultra-reliable and lowlatency computing in the edge with Kubernetes. Journal of Grid Computing, 19(3):??, September 2021. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (elec-

tronic). URL https://link.springer.
com/article/10.1007/s10723-021-
09573-z.

Tusa:2021:IET

[624] Francesco Tusa and Stuart Clayman.

The impact of encoding and transport for massive real-time IoT data on edge resource consumption. Journal of Grid Computing, 19(3):??, September 2021. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic).

URL https://link.springer.com/
article/10.1007/s10723-021-09577-
9.

Zhang:2021:SPM

[625] Jixian Zhang, Ning Xie, and Weidong Li. Strategy-proof mechanism for online time-varying resource allocation with restart. Journal of Grid Computing, 19(3):??, September 2021. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic).

URL https://link.springer.com/
article/10.1007/s10723-021-09563-
1.

Zhu:2021:COS

[626] Anqing Zhu and Youyun Wen. Computing offloading strategy using improved genetic algorithm in mobile edge computing system. Journal of Grid Computing, 19(3):??, September 2021. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic).

URL https://link.springer.com/
article/10.1007/s10723-021-09578-
8.

Ambrosio:2021:ERS

[627] Lenita Ambrósio, Heitor Linhares, and Rafael Capilla. Enhancing the reuse of scientific experiments for agricultural software ecosystems. Journal of Grid Computing, 19(4):??, December 2021. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic).

URL https://link.springer.com/
article/10.1007/s10723-021-09583-

Banerjee:2021:SAS

[628] Sounak Banerjee, Sarbani Roy, and Sunirmal Khatua. SLA-aware stochastic load balancing in dynamic cloud environment. Journal of Grid Computing, 19(4):??, December 2021. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (elec-

tronic). URL https://link.springer.
com/article/10.1007/s10723-021-
09592-w.

Brogi:2021:DAM

[629] Antonio Brogi, Stefano Forti, and Isaac Lera. Declarative application management in the fog. Journal of Grid Computing, 19(4):??, December 2021. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic).

URL https://link.springer.com/
article/10.1007/s10723-021-09582-
y.

Kaur:2021:FFC

[630] Mandeep Kaur and Rajni Aron.

FOCALB: Fog computing architecture of load balancing for scientific workflow applications. Journal of Grid Computing, 19(4):??, December 2021. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic).

URL https://link.springer.com/
article/10.1007/s10723-021-09584-
w.

Maciel:2021:CNS

[631] Douglas B. Maciel, Emidio P. Neto, and Silvio C. Sampaio. Cloud-network slicing MANO towards an efficient IoT-cloud continuum. Journal of Grid Computing, 19(4):??, December 2021. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic).

URL https://link.springer.com/
article/10.1007/s10723-021-09588-
6.

Mahdavisharif:2021:BDA

[632] Mahzad Mahdavisharif, Shahram Jamali, and Reza Fotohi. Big data-aware intrusion detection system in communication networks: a deep learning approach. Journal of Grid Computing, 19 (4):??, December 2021. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (elec-

tronic). URL https://link.springer.
com/article/10.1007/s10723-021-
09581-z.

Pandove:2021:CBR

[633] Divya Pandove and Avleen Malhi. A correlation based recommendation system for large data sets. Journal of Grid Computing, 19(4):??, December 2021. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic).

URL https://link.springer.com/
article/10.1007/s10723-021-09585-
9.

Saha:2021:BTS

[634] Aheli Saha, Yu-Dong Zhang, and Suresh Chandra Satapathy. Brain tumour segmentation with a mutipathway ResNet based UNet. Journal of Grid Computing, 19(4):??, December 2021. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic).

URL https://link.springer.com/
article/10.1007/s10723-021-09590-
y.

Sepulveda-Rodriguez:2021:SBS

[635] Luis E. Sepúlveda-Rodríguez, José Luis Garrido, and Gabriel Guerrero-Contreras. Study-based systematic mapping analysis of cloud technologies for leveraging IT resource and service management: The case study of the science gateway approach. Journal of Grid Computing, 19(4):??, December 2021. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic).

URL https://link.springer.com/
article/10.1007/s10723-021-09587-
7.

Ullah:2021:MET

[636] Amjad Ullah, Huseyin Dagdeviren, and James Bowden. MiCADO-Edge: Towards an application-level orchestrator for the cloud-to-edge computing continuum. Journal of Grid Computing, 19 (4):??, December 2021. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (elec-

tronic). URL https://link.springer.
com/article/10.1007/s10723-021-
09589-5.

Dubey:2022:SIA

[637] Kalka Dubey, S. C. Sharma, and Mohit Kumar. A secure IoT applications allocation framework for integrated fog-cloud environment. Journal of Grid Computing, 20(1):??, March 2022. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic).

URL https://link.springer.com/
article/10.1007/s10723-021-09591-
x.

Fourati:2022:EEP

[638] Mohamed Hedi Fourati, Soumaya Marzouk, and Mohamed Jmaiel. EPMA: Elastic platform for microservicesbased applications: Towards optimal resource elasticity. Journal of Grid Computing, 20(1):??, March 2022. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic).

URL https://link.springer.com/

                                        article/10.1007/s10723-021-09597-
                                        5.

Kumar:2022:SAB

[639] Dinesh Kumar, Gaurav Baranwal, and Deo Prakash Vidyarthi. A survey on auction based approaches for resource allocation and pricing in emerging edge technologies. Journal of Grid Computing, 20(1):??, March 2022. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s10723-021-09593-
                                        9.

Low:2022:RML

[640] Wan Shi Low, Chow Khuen Chan, and Khin Wee Lai. A review of machine learning network in human motion biomechanics. Journal of Grid Computing, 20(1):??, March 2022. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s10723-021-09595-
                                        7.

Ruggeri:2022:IBB

[641] Armando Ruggeri, Antonio Celesti, and Massimo Villari. An innovative blockchain-based orchestrator for osmotic computing. Journal of Grid Computing, 20(1):??, March 2022. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s10723-021-09579-
                                        7.

Wang:2022:SPT

[642] Shui-Hua Wang, Suresh Chandra Satapathy, and Yu-Dong Zhang. Secondary pulmonary tuberculosis identification via pseudo-Zernike moment and deep stacked sparse autoencoder. Journal of Grid Computing, 20(1):??, March 2022. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic).

URL https://link.springer.com/
article/10.1007/s10723-021-09596-
6.

Al-Sayed:2022:WTS

[643] Mustafa M. Al-Sayed. Workload time series cumulative prediction mechanism for cloud resources using neural machine translation technique. Journal of Grid Computing, 20(2):??, June 2022. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic).

URL https://link.springer.com/
article/10.1007/s10723-022-09607-
0.

Bidhendi:2022:ERM

[644] Zohreh Esmaeili Bidhendi and Ehsan Mousavi Khaneghah. ExaFlooding RD: a mathematical model to support unstructured resource discovery in distributed exascale computing environments. Journal of Grid Computing, 20(2):??, June 2022. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic).

URL https://link.springer.com/
article/10.1007/s10723-022-09608-
z.

Du:2022:CCE

[645] RuiZhong Du, Cui Liu, and ZiYuan Wang. Collaborative cloud-edge-end task offloading in NOMA-enabled mobile edge computing using deep learning. Journal of Grid Computing, 20(2): ??, June 2022. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic).

URL https://link.springer.com/

                                        article/10.1007/s10723-022-09605-
                                        2.

Helali:2022:SLC

[646] Leila Helali and Mohamed Nazih Omri. Software license consolidation and resource optimization in container-based virtualized data centers. Journal of Grid Computing, 20(2):??, June 2022. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s10723-022-09602-
                                        5.

Hummaida:2022:SVM

[647] Abdul Rahman Hummaida, Norman W. Paton, and Rizos Sakellariou. Scalable virtual machine migration using reinforcement learning. Journal of Grid Computing, 20(2):??, June 2022. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s10723-022-09603-
                                        4.

Liu:2022:MAC

[648] Yaqin Liu, Chubo Liu, and Keqin Li. Mobility-aware and code-oriented partitioning computation offloading in multi-access edge computing. Journal of Grid Computing, 20(2):??, June 2022. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s10723-022-09599-
                                        x.

Saini:2022:IFS

[649] Kanika Saini, Sheetal Kalra, and Sandeep K. Sood. An integrated framework for smart earthquake prediction:

IoT, fog, and cloud computing. Journal of Grid Computing, 20(2):??, June 2022. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic).

URL https://link.springer.com/
article/10.1007/s10723-022-09600-
7.

Zahoor:2022:FAI

[650] Ehtesham Zahoor, Asim Ikram, and Olivier Perrin. A formal approach for the identification of authorization policy conflicts within multi-cloud environments. Journal of Grid Computing, 20(2):??, June 2022. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (elec-

tronic). URL https://link.springer.
com/article/10.1007/s10723-022-
09606-1.

Zhao:2022:QAI

[651] Defu Zhao, Qunying Zou, and Milad Boshkani Zadeh. A QoS-aware IoT service placement mechanism in fog computing based on opensource development model. Journal of Grid Computing, 20(2):??, June 2022. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic).

URL https://link.springer.com/
article/10.1007/s10723-022-09604-
3.

Kumar:2022:RCN

[652] Abhinav Kumar, Jyoti Prakash Singh, and Amit Kumar Singh. Randomized convolutional neural network architecture for eyewitness tweet identification during disaster. Journal of Grid Computing, 20(3):??, September 2022. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic).

URL https://link.springer.com/

                                        article/10.1007/s10723-022-09609-
                                        y.

Tu:2022:EDL

[653] Jingzhi Tu, Gang Mei, and Francesco Piccialli. An efficient deep learning approach using improved generative adversarial networks for incomplete information completion of self-driving vehicles. Journal of Grid Computing, 20(3):??, September 2022. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s10723-022-09610-
                                        5.

Esmaili:2022:GSC

[654] Samaneh Esmaili and Kamran Kianfar. Grid scheduling considering energy consumption management and quality of service. Journal of Grid Computing, 20 (4):??, December 2022. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (elec-

                                        tronic). URL https://link.springer.
                                        com/article/10.1007/s10723-022-
                                        09620-3.

Ying:2022:CNC

[655] Qianjin Ying, Yulei Yu, and Changzhen Hu. CJSpector: A novel cryptojacking detection method using hardware trace and deep learning. Journal of Grid Computing, 20(4):??, December 2022. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s10723-022-09621-
                                        2.

Hong:2022:AEA

[656] Xiaobin Hong, Jiali Zhang, and Yeganeh Alizadeh. An autonomous evolutionary approach to planning the IoT services placement in the cloud–fog–IoT ecosystem. Journal of Grid Computing, 20(4):??, December 2022. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic).

URL https://link.springer.com/
article/10.1007/s10723-022-09622-
1.

Ahmed:2022:DDP

[657] Quazi Warisha Ahmed and Shruti Garg. Double Diagonal Puzzle Encryption Standard-512 for securing data over cloud environment. Journal of Grid Computing, 20(4):??, December 2022. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic).

URL https://link.springer.com/
article/10.1007/s10723-022-09612-
3.

Tian:2022:TEE

[658] Ping Tian, Huitao Shen, and Ahad Abolfathi. Towards efficient ensemble hierarchical clustering with MapReducebased clusters clustering technique and the innovative similarity criterion. Journal of Grid Computing, 20(4):??, December 2022. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic).

URL https://link.springer.com/
article/10.1007/s10723-022-09623-
0.

Pramanik:2023:MBR

[659] Pijush Kanti Dutta Pramanik, Tarun Biswas, and Prasenjit Choudhury. Multicriteria-based resource-aware scheduling in mobile crowd computing: a heuristic approach. Journal of Grid Computing, 21(1):??, March 2023. CO-DEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s10723-022-09633-y.

Al-Haboobi:2023:DWM

[660] Ali Al-Haboobi and Gabor Kecskemeti. Developing a workflow management system simulation for capturing internal IaaS behavioural knowledge. Journal of Grid Computing, 21(1):??, March 2023. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s10723-022-09638-
                                        7.

Ramezani:2023:MBF

[661] Faeze Ramezani, Saeid Abrishami, and Mehdi Feizi. A market-based framework for resource management in cloud federation. Journal of Grid Computing, 21(1):??, March 2023. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s10723-022-09635-
                                        w.

Chang:2023:SBM

[662] Victor Chang, Jagpreet Sidhu, Sarbjeet Singh, and Rajinder Sandhu. SLAbased multi-dimensional trust model for fog computing environments. Journal of Grid Computing, 21(1):??, March 2023. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s10723-022-09632-
                                        z.

Patel:2023:IDL

[663] Eva Patel and Dharmender Singh Kushwaha. An integrated deep learning prediction approach for efficient modelling of host load patterns in cloud computing. Journal of Grid Computing, 21(1):??, March 2023. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic).

URL https://link.springer.com/
article/10.1007/s10723-022-09639-
6.

Nawrocki:2023:DDA

[664] Piotr Nawrocki, Patryk Osypanka, and Beata Posluszny. Data-driven adaptive prediction of cloud resource usage. Journal of Grid Computing, 21(1):??, March 2023. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic).

URL https://link.springer.com/
article/10.1007/s10723-022-09641-
y.

Eng:2023:EBD

[665] KaiLun Eng, Abdullah Muhammed, Azizol Abdullah, Masnida Hussin, Sazlinah Hasan, and Mohamad Afendee Mohamed. An estimation-based dynamic load balancing algorithm for efficient load distribution and balancing in heterogeneous grid computing environment. Journal of Grid Computing, 21(1):??, March 2023. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic).

URL https://link.springer.com/
article/10.1007/s10723-022-09628-
9.

Lima:2023:ECA

[666] Stanley Lima, Filipe Araujo, Miguel de Oliveira Guerreiro, Jaime Correia, Andre Bento, and Raul Barbosa. Efficient causal access in georeplicated storage systems. Journal of Grid Computing, 21(1):??, March 2023. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s10723-022-09640-
                                        z.

Padulano:2023:LSA

[667] Vincenzo Eduardo Padulano, Ivan Donchev Kabadzhov, Enric Tejedor Saavedra, Enrico Guiraud, and Pedro Alonso-Jordá. Leveraging state-of-the-art engines for large-scale data analysis in high energy physics. Journal of Grid Computing, 21(1):??, March 2023. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s10723-023-09645-
                                        2.

Tarafdar:2023:PME

[668] Anurina Tarafdar, Soumi Sarkar, Rajib K. Das, and Sunirmal Khatua. Power modeling for energy-efficient resource management in a cloud data center. Journal of Grid Computing, 21(1):??, March 2023. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s10723-023-09642-
                                        5.

Long:2023:IFO

[669] Xinjian Long, Xiangyang Gong, Bo Zhang, and Huiyang Zhou. An intelligent framework for oversubscription management in CPU–GPU unified memory. Journal of Grid Computing, 21(1):??, March 2023. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s10723-023-09646-
                                        1.

Noorabad:2023:PDP

[670] Reyhaneh Noorabad, Nasrollah Moghadam Charkari, and Sadegh Dorri Nogoorani. PoMic: Dynamic power management of VM-microservices in overcommitted cloud. Journal of Grid Computing, 21(1):??, March 2023. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (elec-

tronic). URL https://link.springer.
com/article/10.1007/s10723-023-
09648-z.

Li:2023:IDP

[671] Na Li, XiaoLing Liu, Yu Wang, and Musa Mojarad. Improving dynamic placement of virtual machines in cloud data centers based on open-source development model algorithm. Journal of Grid Computing, 21(1):??, March 2023. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic).

URL https://link.springer.com/
article/10.1007/s10723-023-09651-
4.

Zheng:2023:CBC

[672] Yifeng Zheng, Lushan Zou, Wenjie Zhang, Jingmin Yang, Liwei Yang, and Ziqiong Lin. Contract-based cooperative computation and communication resources sharing in mobile edge computing. Journal of Grid Computing, 21(1):??, March 2023. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic).

URL https://link.springer.com/
article/10.1007/s10723-023-09649-
y.

Shan:2023:ISF

[673] Xiaofeng Shan, Chishe Wang, and Dongqin Zhou. Interfering spatiotemporal features and causes of bus bunching using empirical GPS trajectory data. Journal of Grid Computing, 21(1):??, March 2023. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic).

URL https://link.springer.com/
article/10.1007/s10723-023-09652-
3.

Dou:2023:PCB

[674] Jinfeng Dou, Fangzheng Yuan, Jiabao Cao, Xuejia Meng, Xiaoguang Ma, and Zhongwen Guo. Placement combination between heterogeneous services and heterogeneous capacitated servers in edge computing. Journal of Grid Computing, 21(1):??, March 2023. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic).

URL https://link.springer.com/
article/10.1007/s10723-023-09644-
3.

Tao:2023:PRO

[675] Limin Tao, Xikun Liang, Zhijing Wu, Lidong Han, and Jiangping Zhu. Plaintext related optical image hybrid encryption based on fractional Fourier transform and generalized chaos of multiple controlling parameters. Journal of Grid Computing, 21(1):??, March 2023. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic).

URL https://link.springer.com/
article/10.1007/s10723-023-09655-
0.

Loch:2023:STL

[676] Wilton Jaciel Loch and Guilherme Piêgas Koslovski. Sparbit: Towards to a logarithmic-cost and data locality-aware MPI Allgather algorithm. Journal of Grid Computing, 21(2):??, June 2023. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic).

URL https://link.springer.com/
article/10.1007/s10723-023-09650-
5.

Yi:2023:MLM

[677] Haibo Yi. Machine learning method with applications in hardware security of post-quantum cryptography. Journal of Grid Computing, 21(2):??, June 2023. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic).

URL https://link.springer.com/
article/10.1007/s10723-023-09643-
4.

Moina-Rivera:2023:EDS

[678] Wilmer Moina-Rivera, Miguel Garcia-Pineda, Jose M. Claver, and Juan Gutiérrez-Aguado. Event-driven serverless pipelines for video coding and quality metrics. Journal of Grid Computing, 21(2):??, June 2023. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (elec-

tronic). URL https://link.springer.
com/article/10.1007/s10723-023-
09647-0.

Hosny:2023:NIM

[679] Khalid M. Hosny, Ahmed I. Awad, Marwa M. Khashaba, and Ehab R. Mohamed. New improved multiobjective gorilla troops algorithm for dependent tasks offloading problem in multi-access edge computing. Journal of Grid Computing, 21(2):??, June 2023. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic).

URL https://link.springer.com/
article/10.1007/s10723-023-09656-
z.

Wu:2023:MSP

[680] Jimmy Ming-Tai Wu, Huiying Zhou, Jerry Chun-Wei Lin, Gautam Srivastava, and Mohamed Baza. Mining skyline patterns from big data environments based on a Spark framework. Journal of Grid Computing, 21(2):??, June 2023. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic).

URL https://link.springer.com/
article/10.1007/s10723-023-09653-
2.

Chetabi:2023:PAA

[681] Faeze Azimi Chetabi, Mehrdad Ashtiani, and Ehsan Saeedizade. A packageaware approach for function scheduling in serverless computing environments. Journal of Grid Computing, 21(2):??, June 2023. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic).

URL https://link.springer.com/
article/10.1007/s10723-023-09657-
y.

Han:2023:DAV

[682] Yibo Han, Pu Han, Bo Yuan, Zheng Zhang, Lu Liu, and John Panneerselvam. Design and application of vague set theory and adaptive grid particle swarm optimization algorithm in resource scheduling optimization. Journal of Grid Computing, 21(2):??, June 2023. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic).

URL https://link.springer.com/
article/10.1007/s10723-023-09660-
3.

Li:2023:EDO

[683] Mingye Li, Haiwei Lei, Huan Guo, Riza Sulaiman, Wejdan Deebani, and Meshal Shutaywi. Efficient data offloading using Markovian decision on state reward action in edge computing. Journal of Grid Computing, 21(2):??, June 2023. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic).

URL https://link.springer.com/
article/10.1007/s10723-023-09659-
w.

Chen:2023:EAM

[684] Liqiong Chen, Yingda Liu, Yijun Lu, and Huaiying Sun. Energyaware and mobility-driven computation offloading in MEC. Journal of Grid Computing, 21(2):??, June 2023. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic).

URL https://link.springer.com/
article/10.1007/s10723-023-09654-
1.

Ma:2023:CPS

[685] Xiang Ma, Laila Almutairi, Ahmed M.

Alwakeel, and Mohammed Hameed Alhameed. Cyber physical system for distributed network using DoS based hierarchical Bayesian network. Journal of Grid Computing, 21(2):??, June 2023. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic).

URL https://link.springer.com/
article/10.1007/s10723-023-09662-
1.

Kumar:2023:SIP

[686] Mantosh Kumar, Kumari Namrata, Nishant Kumar, and Gaurav Saini. Solar irradiance prediction using an optimized data driven machine learning models. Journal of Grid Computing, 21(2):??, June 2023. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic).

URL https://link.springer.com/
article/10.1007/s10723-023-09668-

Liu:2023:QRF

[687] Xiling Liu and Shuisheng Zhou. Qualityrelated fault detection based on approximate kernel partial least squares method. Journal of Grid Computing, 21(2):??, June 2023. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (elec-

                                        tronic). URL https://link.springer.
                                        com/article/10.1007/s10723-023-
                                        09670-1.

Wang:2023:RLF

[688] Zhechao Wang, Qiming Fu, Jianping Chen, Yunzhe Wang, You Lu, and Hongjie Wu. Reinforcement learning in few-shot scenarios: a survey. Journal of Grid Computing, 21(2):??, June 2023. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s10723-023-09663-
                                        0.

Hu:2023:TSM

[689] Qianxue Hu, Xiaofei Wu, and Shoubin Dong. A two-stage multi-objective task scheduling framework based on invasive tumor growth optimization algorithm for cloud computing. Journal of Grid Computing, 21(2):??, June 2023. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s10723-023-09665-
                                        y.

Sun:2023:AQP

[690] Changyuan Sun, Jingjing Li, Riza Sulaiman, Badr S. Alotaibi, Samia Elattar, and Mohammed Abuhussain. Air quality prediction and multi-task offloading based on deep learning methods in edge computing. Journal of Grid Computing, 21(2):??, June 2023. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic).

URL https://link.springer.com/
article/10.1007/s10723-023-09671-
0.

Li:2023:CAS

[691] Hongjian Li, Lisha Zhu, Shuaicheng Wang, and Lei Wang. Cost-aware scheduling and data skew alleviation for big data processing in heterogeneous cloud environment. Journal of Grid Computing, 21(3):??, September 2023. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic).

URL https://link.springer.com/
article/10.1007/s10723-023-09661-
2.

Kjorveziroski:2023:WEN

[692] Vojdan Kjorveziroski and Sonja Filiposka. WebAssembly as an enabler for next generation serverless computing. Journal of Grid Computing, 21 (3):??, September 2023. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (elec-

tronic). URL https://link.springer.
com/article/10.1007/s10723-023-
09669-8.

Chaudhary:2023:SAR

[693] Ajay Chaudhary, Sateesh K Peddoju, and Vikas Chouhan. Secure authentication and reliable cloud storage scheme for IoT–Edge-Cloud integration. Journal of Grid Computing, 21 (3):??, September 2023. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (elec-

tronic). URL https://link.springer.
com/article/10.1007/s10723-023-
09672-z.

Yang:2023:FFE

[694] Hailin Yang, Yanhong Huang, Jianqi Shi, and Yang Yang. A federated framework for edge computing devices with collaborative fairness and adversarial robustness. Journal of Grid Computing, 21(3):??, September 2023. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s10723-023-09658-
                                        x.

Dong:2023:HDQ

[695] YanRu Dong, Ahmed M. Alwakeel, Mohammed M. Alwakeel, Lubna A. Alharbi, and Sara A Althubiti. A heuristic deep Q learning for offloading in edge devices in 5 g networks. Journal of Grid Computing, 21(3):??, September 2023. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s10723-023-09667-
                                        w.

Xu:2023:NEI

[696] Kun Xu and Jichang Guo. A novel edgeinspired depth quality evaluation network for RGB-D salient object detection. Journal of Grid Computing, 21 (3):??, September 2023. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (elec-

                                        tronic). URL https://link.springer.
                                        com/article/10.1007/s10723-023-
                                        09674-x.

Rybinski:2023:VLC

[697] Kamil Rybiński, Michał Śmiałek, Agris Sostaks, Krzysztof Marek, Radosław Roszczyk, and Marek Wdowiak. Visual low-code language for orchestrating large-scale distributed computing.

Journal of Grid Computing, 21(3): ??, September 2023. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (elec-

tronic). URL https://link.springer.
com/article/10.1007/s10723-023-
09666-x.

Taghavi:2023:CEW

[698] Bahareh Taghavi, Behrooz Zolfaghari, and Saeid Abrishami. A cost-efficient workflow as a service broker using ondemand and spot instances. Journal of Grid Computing, 21(3):??, September 2023. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic).

URL https://link.springer.com/
article/10.1007/s10723-023-09676-
9.

Tong:2023:DSA

[699] Zhao Tong, Bilan Liu, Jing Mei, Jiake Wang, Xin Peng, and Keqin Li. Data security aware and effective task offloading strategy in mobile edge computing. Journal of Grid Computing, 21(3):??, September 2023. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic).

URL https://link.springer.com/
article/10.1007/s10723-023-09673-
y.

Tedeschi:2023:SCD

[700] Tommaso Tedeschi, Marco Baioletti, Diego Ciangottini, Valentina Poggioni, Daniele Spiga, Loriano Storchi, and Mirco Tracolli. Smart caching in a data lake for high energy physics analysis. Journal of Grid Computing, 21 (3):??, September 2023. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (elec-

tronic). URL https://link.springer.
com/article/10.1007/s10723-023-
09664-z.

Brenner:2023:DDS

[701] Garrett Brenner, Mohamed Baza, Amar Rasheed, Wassila Lalouani, Mahmoud Badr, and Hani Alshahrani. DPark: Decentralized smart private-parking system using blockchains. Journal of Grid Computing, 21(3):??, September 2023. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s10723-023-09677-
                                        8.

Kosinska:2023:ECN

[702] Joanna Kosińska and Krzysztof Zieliński. Enhancement of cloud-native applications with autonomic features. Journal of Grid Computing, 21(3):??, September 2023. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s10723-023-09675-
                                        w.

Raju:2023:BAC

[703] K. Raju, N. Ramshankar, J. Anvar Shathik, and R. Lavanya. Blockchain assisted cloud security and privacy preservation using hybridized encryption and deep learning mechanism in IoT–Healthcare application. Journal of Grid Computing, 21(3):??, September 2023. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s10723-023-09678-
                                        7.

Alelyani:2023:DDA

[704] Abdullah Alelyani, Amitava Datta, and Ghulam Mubashar Hassan. DAScheduler: Dependency-aware scheduling algorithm for containerized dependent jobs. Journal of Grid Computing, 21 (3):??, September 2023. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (elec-

tronic). URL https://link.springer.
com/article/10.1007/s10723-023-
09679-6.

Kshatriya:2023:EHS

[705] Divya Kshatriya and Vijayalakshmi A.

Lepakshi. An efficient hybrid scheduling framework for optimal workload execution in federated clouds to maintain performance SLAs. Journal of Grid Computing, 21(3):??, September 2023. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic).

URL https://link.springer.com/
article/10.1007/s10723-023-09682-
x.

Naqvi:2023:AAV

[706] Syed Muhammad Ali Naqvi, Mohammad Shabaz, Muhammad Attique Khan, and Syeda Iqra Hassan. Adversarial attacks on visual objects using the fast gradient sign method. Journal of Grid Computing, 21(4):??, December 2023. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic).

URL https://link.springer.com/
article/10.1007/s10723-023-09684-
9.

Chen:2023:MTO

[707] Dacheng Chen and Xinhua Liu. Mayfly Taylor optimization-based graph attention network for task scheduling in edge computing. Journal of Grid Computing, 21(4):??, December 2023. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic).

URL https://link.springer.com/
article/10.1007/s10723-023-09685-
8.

Cheng:2023:MPC

[708] Bei Cheng. Multi-population cooperative elite algorithm for efficient computation offloading in mobile edge computing. Journal of Grid Computing, 21(4):??, December 2023. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s10723-023-09688-
                                        5.

Brahmam:2023:IDI

[709] Madala Guru Brahmam and Vijay Anand R. ICSMPC: Design of an iterative-learning contextual side chaining model for improving security of priority-aware cloud resources. Journal of Grid Computing, 21(4):??, December 2023. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s10723-023-09687-
                                        6.

Luo:2023:IFE

[710] Tianye Luo, Shijun Li, Ji Li, Jie Guo, Ruilong Feng, Ye Mu, Tianli Hu, Yu Sun, Ying Guo, and He Gong. Image fuzzy edge information segmentation based on computer vision and machine learning. Journal of Grid Computing, 21(4):??, December 2023. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s10723-023-09697-
                                        4.

Ali:2023:IBB

[711] Jehad Ali, Gaoyang Shan, Noor Gul, and Byeong hee Roh. An intelligent blockchain-based secure link failure recovery framework for softwaredefined Internet-of-things. Journal of Grid Computing, 21(4):??, December 2023. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic).

URL https://link.springer.com/
article/10.1007/s10723-023-09693-
8.

Liu:2023:MJO

[712] ZhangRong Liu. A multi-joint optimisation method for distributed edge computing resources in IoT-based smart cities. Journal of Grid Computing, 21 (4):??, December 2023. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (elec-

tronic). URL https://link.springer.
com/article/10.1007/s10723-023-
09701-x.

Zhou:2023:HBI

[713] Xiangming Zhou, Laila Almutairi, Theyab R. Alsenani, and Mohammad Nazir Ahmad. Honeypot based industrial threat detection using game theory in cyber–physical system. Journal of Grid Computing, 21(4):??, December 2023. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic).

URL https://link.springer.com/
article/10.1007/s10723-023-09689-
4.

Lu:2023:AAI

[714] Jiazhong Lu, Jin Lan, Yuanyuan Huang, Maojia Song, and Xiaolei Liu. Anti-attack intrusion detection model based on MPNN and traffic spatiotemporal characteristics. Journal of Grid Computing, 21(4):??, December 2023. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic).

URL https://link.springer.com/

                                        article/10.1007/s10723-023-09703-
                                        9.

Ghandour:2023:CRS

[715] Oumaima Ghandour, Said El Kafhali, and Mohamed Hanini. Computing resources scalability performance analysis in cloud computing data center. Journal of Grid Computing, 21(4):??, December 2023. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s10723-023-09696-
                                        5.

Yang:2023:RTP

[716] Li Yang, Xiong Xiao, Xuedong Zhang, Zhechang Hu, and Zhuo Tang. A real-time partition generation mechanism for data skew mitigation in Spark computing environment. Journal of Grid Computing, 21(4):??, December 2023. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s10723-023-09700-
                                        y.

Golec:2023:BBE

[717] Muhammed Golec, Sukhpal Singh Gill, Mustafa Golec, Minxian Xu, Soumya K. Ghosh, Salil S. Kanhere, Omer Rana, and Steve Uhlig. BlockFaaS: Blockchain-enabled serverless computing framework for AI-driven IoT healthcare applications. Journal of Grid Computing, 21(4):??, December 2023. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s10723-023-09691-
                                        w.

Syed:2023:MDC

[718] Salman Azeez Syed, Deepak Kumar Sharma, and Gautam Srivastava. Modeling distributed and configurable hierarchical blockchain over SDN and fog-based networks for largescale Internet of Things. Journal of Grid Computing, 21(4):??, December 2023. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic).

URL https://link.springer.com/
article/10.1007/s10723-023-09698-
3.

Nasiri:2023:HSH

[719] Hamid Nasiri, Armin Darjani, Nima Kavand, and Maziar Goudarzi. H-Storm: a hybrid CPU–FPGA architecture to accelerate Apache Storm. Journal of Grid Computing, 21(4):??, December 2023. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic).

URL https://link.springer.com/
article/10.1007/s10723-023-09692-
9.

Meng:2023:NEC

[720] Lingbin Meng and Daofeng Li. Novel edge computing-based privacy-preserving approach for smart healthcare systems in the Internet of Medical Things. Journal of Grid Computing, 21(4):??, December 2023. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic).

URL https://link.springer.com/
article/10.1007/s10723-023-09695-
6.

Akhtar:2023:MDL

[721] Md. Mobin Akhtar, Abdallah Saleh Ali Shatat, Mukhtar Al-Hashimi, Abu Sarwar Zamani, Mohammed Rizwanullah, Sara Saadeldeen Ibrahim Mohamed, and Rashid Ayub. MapReduce with deep learning framework for student health monitoring system using IoT technology for big data. Journal of Grid Computing, 21(4):??, December 2023. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic).

URL https://link.springer.com/
article/10.1007/s10723-023-09690-
x.

Zhou:2023:NAC

[722] Hong Zhou. A novel approach to cloud resource management: Hybrid machine learning and task scheduling. Journal of Grid Computing, 21(4):??, December 2023. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic).

URL https://link.springer.com/
article/10.1007/s10723-023-09702-
w.

Awasthi:2023:PSD

[723] Charu Awasthi, Prashant Kumar Mishra, Pawan Kumar Pal, Surbhi Bhatia Khan, Ambuj Kumar Agarwal, Thippa Reddy Gadekallu, and Areej A. Malibari. Preservation of sensitive data using multi-level blockchain-based secured framework for edge network devices. Journal of Grid Computing, 21 (4):??, December 2023. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (elec-

tronic). URL https://link.springer.
com/article/10.1007/s10723-023-
09699-2.

Li:2023:HIW

[724] Jizhou Li, Qi Wang, Shuai Hu, and Ling Li. Hybrid immune whale differential evolution optimization (HI-WDEO) based computation offloading in MEC for IoT. Journal of Grid Computing, 21(4):??, December 2023. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic).

URL https://link.springer.com/
article/10.1007/s10723-023-09705-
7.

Zhang:2023:SSB

[725] Shuangshuang Zhang, Yue Tang, Dinghui Wang, Noorliza Karia, and Chenguang Wang. Secured SDN based task scheduling in edge computing for smart city health monitoring operation management system. Journal of Grid Computing, 21(4):??, December 2023. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic).

URL https://link.springer.com/
article/10.1007/s10723-023-09707-
5.

Zhong:2023:SED

[726] Yuting Zhong, Zesheng Qin, Abdulmajeed Alqhatani, Ahmed Sayed M. Metwally, Ashit Kumar Dutta, and Joel J. P. C. Rodrigues. Sustainable environmental design using green IOT with hybrid deep learning and building algorithm for smart city. Journal of Grid Computing, 21(4):??, December 2023. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic).

URL https://link.springer.com/
article/10.1007/s10723-023-09704-
8.

ZargarAzad:2023:ASA

[727] Matineh ZargarAzad and Mehrdad Ashtiani. An auto-scaling approach for microservices in cloud computing environments. Journal of Grid Computing, 21 (4):??, December 2023. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (elec-

tronic). URL https://link.springer.

                                        com/article/10.1007/s10723-023-
                                        09713-7.

Aknan:2023:ABA

[728] Mohammad Aknan, Maheshwari Prasad Singh, and Rajeev Arya. AI and blockchain assisted framework for offloading and resource allocation in fog computing. Journal of Grid Computing, 21(4):??, December 2023. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (elec-

                                        tronic). URL https://link.springer.
                                        com/article/10.1007/s10723-023-
                                        09694-7.

Zhang:2023:EPM

[729] Longxin Zhang, Minghui Ai, Runti Tan, Junfeng Man, Xiaojun Deng, and Keqin Li. Efficient prediction of makespan matrix workflow scheduling algorithm for heterogeneous cloud environments. Journal of Grid Computing, 21(4):??, December 2023. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s10723-023-09711-
                                        9.

Cheng:2023:SFI

[730] Caijun Cheng and Huazhen Huang. Smart financial Investor’s risk prediction system using mobile edge computing. Journal of Grid Computing, 21 (4):??, December 2023. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (elec-

                                        tronic). URL https://link.springer.
                                        com/article/10.1007/s10723-023-
                                        09710-w.

Tian:2023:DLB

[731] Yanjia Tian, Yan Dong, and Xiang Feng. Deep learning-based multidomain framework for end-to-end services in 5G networks. Journal of Grid Computing, 21(4):??, December 2023. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic).

URL https://link.springer.com/
article/10.1007/s10723-023-09714-
6.

Gao:2023:DRL

[732] Xiaohu Gao, Mei Choo Ang, and Sara A. Althubiti. Deep reinforcement learning and Markov decision problem for task offloading in mobile edge computing. Journal of Grid Computing, 21(4):??, December 2023. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic).

URL https://link.springer.com/
article/10.1007/s10723-023-09708-
4.

Sharma:2023:BAC

[733] Deepak Sharma, Rajeev Kumar, and Ki-Hyun Jung. A bibliometric analysis of convergence of artificial intelligence and blockchain for edge of things. Journal of Grid Computing, 21(4):??, December 2023. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic).

URL https://link.springer.com/
article/10.1007/s10723-023-09716-
4.

Bento:2023:CAA

[734] Andre Bento, Filipe Araujo, and Raul Barbosa. Cost-availability aware scaling: Towards optimal scaling of cloud services. Journal of Grid Computing, 21(4):??, December 2023. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (elec-

tronic). URL https://link.springer.
com/article/10.1007/s10723-023-
09718-2.

Jin:2023:ILC

[735] Hulin Jin, Zhiran Jin, Yong-Guk Kim, and Chunyang Fan. Integration of a lightweight customized 2D CNN model to an edge computing system for realtime multiple gesture recognition. Journal of Grid Computing, 21(4):??, December 2023. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s10723-023-09715-
                                        5.

Liu:2023:HFS

[736] Yang-Yang Liu, Ying Zhang, Yue Wu, and Man Feng. Healthcare and fitness services: a comprehensive assessment of blockchain, IoT, and edge computing in smart cities. Journal of Grid Computing, 21(4):??, December 2023. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s10723-023-09712-
                                        8.

Gai:2024:MFD

[737] Yuliang Gai, Yuxin Liu, Minghao Li, and Shengcheng Yang. Markovian with federated deep recurrent neural network for Edge–IoMT to improve healthcare in smart cities. Journal of Grid Computing, 22(1):??, March 2024. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s10723-023-09709-
                                        3.

Song:2024:DDL

[738] Yu Song, Xin He, Xiwang Tang, Bo Yin, Jie Du, Jiali Liu, Zhongbao Zhao, and Shigang Geng. DEEPBIN:

Deep learning based garbage classification for households using sustainable natural technologies. Journal of Grid Computing, 22(1):??, March 2024. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic).

URL https://link.springer.com/
article/10.1007/s10723-023-09722-
6.

Chai:2024:ECF

[739] Sheng Chai and Lantian Guo. Edge computing with fog-cloud for heart data processing using particle swarm optimized deep learning technique. Journal of Grid Computing, 22(1):??, March 2024. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic).

URL https://link.springer.com/
article/10.1007/s10723-023-09706-
6.

Zhang:2024:DAO

[740] Ming Zhang. Development of analytical offloading for innovative Internet of Vehicles based on mobile edge computing. Journal of Grid Computing, 22(1):??, March 2024. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic).

URL https://link.springer.com/
article/10.1007/s10723-023-09719-
1.

Zhen:2024:AIS

[741] Xi Zhen and Li Zhen. Accounting information systems and strategic performance: The interplay of digital technology and edge computing devices. Journal of Grid Computing, 22(1):??, March 2024. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic).

URL https://link.springer.com/
article/10.1007/s10723-023-09720-
8.

Mitropoulou:2024:ADC

[742] Katerina Mitropoulou, Panagiotis Kokkinos, Polyzois Soumplis, and Emmanouel Varvarigos. Anomaly detection in cloud computing using knowledge graph embedding and machine learning mechanisms. Journal of Grid Computing, 22(1):??, March 2024. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s10723-023-09727-
                                        1.

Han:2024:NTD

[743] Yibo Han, Pu Han, Bo Yuan, Zheng Zhang, Lu Liu, and John Panneerselvam. Novel transformation deep learning model for electrocardiogram classification and arrhythmia detection using edge computing. Journal of Grid Computing, 22(1):??, March 2024. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s10723-023-09717-
                                        3.

XuanWen:2024:PCB

[744] XuanWen and Hai Meng Sun. Parking cooperation-based mobile edge computing using task offloading strategy. Journal of Grid Computing, 22(1):??, March 2024. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s10723-023-09721-
                                        7.

Yin:2024:CEB

[745] Huayi Yin, Xindong Huang, and Erzhong Cao. A cloud–edge-based multi-objective task scheduling approach for smart manufacturing lines. Journal of Grid Computing, 22(1):??, March 2024. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic).

URL https://link.springer.com/
article/10.1007/s10723-023-09723-
5.

Liu:2024:HFN

[746] Bingtao Liu. Hybrid fuzzy neural network for joint task offloading in the Internet of Vehicles. Journal of Grid Computing, 22(1):??, March 2024. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic).

URL https://link.springer.com/
article/10.1007/s10723-023-09724-
4.

Nasr:2024:CNC

[747] Aida A. Nasr. CMSV: a new cloud multi-agents for self-driving vehicles as a services. Journal of Grid Computing, 22(1):??, March 2024. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (elec-

tronic). URL https://link.springer.
com/article/10.1007/s10723-023-
09734-2.

Yang:2024:OAI

[748] Xiaofeng Yang. Optimizing accounting informatization through simultaneous multi-tasking across edge and cloud devices using hybrid machine learning models. Journal of Grid Computing, 22(1):??, March 2024. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (elec-

tronic). URL https://link.springer.
com/article/10.1007/s10723-023-
09735-1.

Sinha:2024:ADT

[749] Avishek Sinha, Samayveer Singh, and Harsh K. Verma. AI-driven task scheduling strategy with blockchain integration for edge computing. Journal of Grid Computing, 22(1):??, March 2024. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic).

URL https://link.springer.com/
article/10.1007/s10723-024-09743-
9.

Li:2024:LTD

[750] Xuemei Li, Xuelian Liu, Da Xie, and Chong Chen. 3D lidar target detection method at the edge for the cloud continuum. Journal of Grid Computing, 22(1):??, March 2024. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic).

URL https://link.springer.com/
article/10.1007/s10723-023-09736-
0.

Song:2024:IDU

[751] Xiedong Song and Qinmin Ma. Intrusion detection using federated attention neural network for edge enabled Internet of Things. Journal of Grid Computing, 22(1):??, March 2024. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic).

URL https://link.springer.com/
article/10.1007/s10723-023-09725-
3.

Tsang:2024:CCD

[752] Y. P. Tsang, C. K. M. Lee, Kening Zhang, C. H. Wu, and W. H. Ip. On-chain and off-chain data management for blockchain–Internet of Things: a multi-agent deep reinforcement learning approach. Journal of Grid Computing, 22(1):??, March 2024. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic).

URL https://link.springer.com/

article/10.1007/s10723-023-09739-
x.

Entrialgo:2024:JAC

[753] Joaquín Entrialgo, Manuel García, Javier García, José María López, and José Luis Díaz. Joint autoscaling of containers and virtual machines for cost optimization in container clusters. Journal of Grid Computing, 22(1):??, March 2024. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic).

URL https://link.springer.com/
article/10.1007/s10723-023-09732-
4.

Zhang:2024:DBT

[754] Ziyang Zhang, Keyu Gu, and Zijie Xu. DRL-based task and computational offloading for Internet of Vehicles in decentralized computing. Journal of Grid Computing, 22(1):??, March 2024. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic).

URL https://link.springer.com/
article/10.1007/s10723-023-09729-
z.

Elma:2024:HBW

[755] K Johny Elma, Praveena Rachel Kamala S, and Saraswathi T. Hybridized black widow–honey badger optimization: Swarm intelligence strategy for node localization scheme in WSN. Journal of Grid Computing, 22(1):??, March 2024. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic).

URL https://link.springer.com/
article/10.1007/s10723-024-09740-
y.

Hu:2024:APR

[756] Daidi Hu. Automated pallet racking examination in edge platform based on MobileNetV2: Towards smart manufacturing. Journal of Grid Computing, 22(1):??, March 2024. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (elec-

tronic). URL https://link.springer.
com/article/10.1007/s10723-023-
09738-y.

Zhang:2024:FDR

[757] Xinyu Zhang, Zhigang Hu, Yang Liang, Hui Xiao, Aikun Xu, Meiguang Zheng, and Chuan Sun. A federated deep reinforcement learningbased low-power caching strategy for cloud-edge collaboration. Journal of Grid Computing, 22(1):??, March 2024. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic).

URL https://link.springer.com/
article/10.1007/s10723-023-09730-
6.

Ye:2024:CLS

[758] Wenxia Ye. E-commerce logistics and supply chain network optimization for cross-border. Journal of Grid Computing, 22(1):??, March 2024. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic).

URL https://link.springer.com/
article/10.1007/s10723-023-09737-
z.

Rai:2024:CAP

[759] Shanvendra Rai, Rituparna Paul, Subhasish Banerjee, Preetisudha Meher, and Gulab Sah. A combined approach of PUF and physiological data for mutual authentication and key agreement in WMSN. Journal of Grid Computing, 22(1):??, March 2024. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic).

URL https://link.springer.com/
article/10.1007/s10723-023-09731-
5.

Bajpai:2024:MGO

[760] Sanchika Abhay Bajpai and Archana B.

Patankar. Marine goal optimizer tuned deep BiLSTM-based self-configuring intrusion detection in cloud. Journal of Grid Computing, 22(1):??, March 2024. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic).

URL https://link.springer.com/
article/10.1007/s10723-023-09728-
0.

Wang:2024:SCT

[761] Mengqi Wang, Jiayuan Mao, Wei Zhao, Xinya Han, Mengya Li, Chuanjun Liao, Haomiao Sun, and Kexin Wang. Smart city transportation: a VANET edge computing model to minimize latency and delay utilizing 5G network. Journal of Grid Computing, 22(1):??, March 2024. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic).

URL https://link.springer.com/
article/10.1007/s10723-024-09747-
5.

Liu:2024:IBC

[762] Jianjia Liu, Xin Yang, Tiannan Liao, and Yong Hang. An IoT-based Covid-19 healthcare monitoring and prediction using deep learning methods. Journal of Grid Computing, 22(1):??, March 2024. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic).

URL https://link.springer.com/
article/10.1007/s10723-024-09742-
w.

Chai:2024:DTS

[763] Sheng Chai and Jimmy Huang. Dependent task scheduling using parallel deep neural networks in mobile edge computing. Journal of Grid Computing, 22(1):??, March 2024. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s10723-024-09744-
                                        8.

Zhao:2024:JTO

[764] Jie Zhao and Ahmed M. El-Sherbeeny. Joint task offloading and multi-task offloading based on NOMA enhanced Internet of Vehicles in edge computing. Journal of Grid Computing, 22(1):??, March 2024. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s10723-024-09748-
                                        4.

Wan:2024:DRM

[765] Xucheng Wan. Dynamic resource management in MEC powered by edge intelligence for smart city Internet of Things. Journal of Grid Computing, 22(1):??, March 2024. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s10723-024-09749-
                                        3.

Wang:2024:ECE

[766] Kemeng Wang, Shurui Kong, Xuezheng Chen, and Min Zhao. Edge computing empowered smart healthcare: Monitoring and diagnosis with deep learning methods. Journal of Grid Computing, 22(1):??, March 2024. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (elec-

tronic). URL https://link.springer.
com/article/10.1007/s10723-023-
09726-2.

Chauhan:2024:PDA

[767] Naveen Chauhan and Rajeev Agrawal.

A probabilistic deadline-aware application offloading in a multi-queueing fog system: a max entropy framework. Journal of Grid Computing, 22(1):??, March 2024. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic).

URL https://link.springer.com/
article/10.1007/s10723-024-09753-
7.

Tian:2024:ERP

[768] Hao Tian, Ruiheng Li, Yi Di, Qiankun Zuo, and Jinpeng Wang. Employing RNN and Petri nets to secure edge computing threats in smart cities. Journal of Grid Computing, 22(1):??, March 2024. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic).

URL https://link.springer.com/
article/10.1007/s10723-023-09733-
3.

Javadpour:2024:DAB

[769] Amir Javadpour, Arun Kumar Sangaiah, Weizhe Zhang, Ankit Vidyarthi, and HamidReza Ahmadi. Decentralized AI-based task distribution on blockchain for Cloud Industrial Internet of Things. Journal of Grid Computing, 22(1):??, March 2024. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic).

URL https://link.springer.com/
article/10.1007/s10723-024-09751-
9.

Jin:2024:JTO

[770] Hulin Jin, Yong-Guk Kim, Zhiran Jin, Chunyang Fan, and Yonglong Xu.

Joint task offloading based on distributed deep reinforcement learningbased genetic optimization algorithm for Internet of Vehicles. Journal of Grid Computing, 22(1):??, March 2024. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic).

URL https://link.springer.com/
article/10.1007/s10723-024-09741-
x.

Guo:2024:DMR

[771] Hao Guo and Weidong Li. Dynamic multi-resource fair allocation with elastic demands. Journal of Grid Computing, 22(1):??, March 2024. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic).

URL https://link.springer.com/
article/10.1007/s10723-024-09754-
6.

Xing:2024:WSC

[772] Yanli Xing. Work scheduling in cloud network based on deep Q-LSTM models for efficient resource utilization. Journal of Grid Computing, 22(1):??, March 2024. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic).

URL https://link.springer.com/
article/10.1007/s10723-024-09746-
6.

Pang:2024:DDD

[773] Haotian Pang and Zhanwei Wang. Dueling double deep Q network strategy in MEC for smart Internet of Vehicles edge computing networks. Journal of Grid Computing, 22(1):??, March 2024. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic).

URL https://link.springer.com/
article/10.1007/s10723-024-09752-
8.

Wang:2024:MAS

[774] Shangshang Wang, Yuqin Jing, Kezhu Wang, and Xue Wang. Multi-agent systems for collaborative inference based on deep policy q-inference network. Journal of Grid Computing, 22(1):??, March 2024. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic).

URL https://link.springer.com/
article/10.1007/s10723-024-09750-
w.

Li:2024:ASF

[775] Hongjian Li, Wei Luo, Wenbin Xie, Huaqing Ye, and Xiaolin Duan. Adaptive scheduling framework of streaming applications based on resource demand prediction with hybrid algorithms. Journal of Grid Computing, 22(1):??, March 2024. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic).

URL https://link.springer.com/
article/10.1007/s10723-024-09756-
4.

Karmakar:2024:CEW

[776] Kamalesh Karmakar, Anurina Tarafdar, Rajib K. Das, and Sunirmal Khatua. Cost-efficient workflow as a service using containers. Journal of Grid Computing, 22(1):??, March 2024. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic).

URL https://link.springer.com/
article/10.1007/s10723-024-09745-
7.

Qiu:2024:IPS

[777] Shujie Qiu. Improving performance of smart education systems by integrating machine learning on edge devices and cloud in educational institutions. Journal of Grid Computing, 22(1):??, March 2024. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s10723-024-09755-
                                        5.

Xu:2024:JDM

[778] Bo Xu, Jialu Guo, Fangling Ma, Menglan Hu, Wei Liu, and Kai Peng. On the joint design of microservice deployment and routing in cloud data centers. Journal of Grid Computing, 22(2): ??, June 2024. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s10723-024-09759-
                                        1.

Hai:2024:CSI

[779] Tao Hai, Muammer Aksoy, Celestine Iwendi, Ebuka Ibeke, and Senthilkumar Mohan. CIA security for Internet of Vehicles and Blockchain–AI integration. Journal of Grid Computing, 22(2):??, June 2024. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s10723-024-09757-
                                        3.

Kaushik:2024:MFB

[780] Neha Kaushik, Harish Kumar, and Vinay Raj. Micro frontend based performance improvement and prediction for microservices using machine learning. Journal of Grid Computing, 22(2): ??, June 2024. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s10723-024-09760-
                                        8.

Chen:2024:ESB

[781] Yu Chen, Yilun Qiu, Zhenyu Tang, Shuling Long, Lingfeng Zhao, and Zhong Tang. Exploring the synergy of blockchain, IoT, and edge computing in smart traffic management across urban landscapes. Journal of Grid Computing, 22(2):??, June 2024. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic).

URL https://link.springer.com/
article/10.1007/s10723-024-09762-
6.

Sus:2024:SBA

[782] Wiktor Sus and Piotr Nawrocki.

Signature-based adaptive cloud resource usage prediction using machine learning and anomaly detection. Journal of Grid Computing, 22(2):??, June 2024. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic).

URL https://link.springer.com/
article/10.1007/s10723-024-09764-
4.

Fu:2024:ESO

[783] Qiang Fu and Tao Yang. Enhancing service offloading for dense networks based on optimal stopping theory in virtual mobile edge computing. Journal of Grid Computing, 22(2):??, June 2024. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic).

URL https://link.springer.com/
article/10.1007/s10723-024-09765-
3.

Ayadi:2024:ECW

[784] Mouhamed Gaith Ayadi and Haithem Mezni. Enabling configurable workflows in smart environments with knowledgebased process fragment reuse. Journal of Grid Computing, 22(2):??, June 2024. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic).

URL https://link.springer.com/
article/10.1007/s10723-024-09763-
5.

Kosari:2024:HDG

[785] Saeed Kosari, Mirsaeid Hosseini Shirvani, Navid Khaledian, and Danial Javaheri. A hybrid discrete grey Wolf optimization algorithm imbalance-ness aware for solving two-dimensional binpacking problems. Journal of Grid Computing, 22(2):??, June 2024. CO-DEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL https:

//link.springer.com/article/10.
1007/s10723-024-09761-7.

Ayadi:2024:CEC

[786] Mouhamed Gaith Ayadi and Haithem Mezni. Correction: Enabling configurable workflows in smart environments with knowledge-based process fragment reuse. Journal of Grid Computing, 22(2):??, June 2024. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (elec-

tronic). URL https://link.springer.
com/article/10.1007/s10723-024-
09769-z.

Sun:2024:JTD

[787] Yuan Sun, Chen Zhang, and Tao Huang. Joint task dispatching and bandwidth allocation with hard deadlines in distributed serverless edge computing systems. Journal of Grid Computing, 22(2): ??, June 2024. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic).

URL https://link.springer.com/
article/10.1007/s10723-024-09770-

Yang:2024:HHR

[788] Zhe Yang, Zitong Ma, Wenbo Zhao, Lingzhi Li, and Fei Gu. HRNN: Hypergraph recurrent neural network for network intrusion detection. Journal of Grid Computing, 22(2):??, June 2024. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic).

URL https://link.springer.com/
article/10.1007/s10723-024-09767-
1.

Chen:2024:DLB

[789] Jiong Chen, Abdullah Alshammari, Mohammed Alonazi, Aisha M. Alqahtani, Sara A. Althubiti, and Romi Fadillah Rahmat. Deep learning based entropy controlled optimization for the detection of Covid-19. Journal of Grid Computing, 22(2):??, June 2024. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic).

URL https://link.springer.com/
article/10.1007/s10723-024-09766-
2.

Xu:2024:EPR

[790] Guangyu Xu and Mingde Xu. An effective prediction of resource using machine learning in edge environments for the smart healthcare industry. Journal of Grid Computing, 22(2):??, June 2024. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic).

URL https://link.springer.com/
article/10.1007/s10723-024-09768-
0.

Bakhshi:2024:ESP

[791] Zeinab Bakhshi, Guillermo Rodriguez-Navas, Hans Hansson, and Radu Prodan. Evaluation of storage placement in computing continuum for a robotic application. Journal of Grid Computing, 22(2):??, June 2024. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (elec-

                                        tronic). URL https://link.springer.
                                        com/article/10.1007/s10723-024-
                                        09758-2.

Li:2024:TEE

[792] Saihong Li, Yingying Ma, Yusha Zhang, and Yinghui Xie. Towards enhanced energy aware resource optimization for edge devices through multicluster communication systems. Journal of Grid Computing, 22(2):??, June 2024. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s10723-024-09773-
                                        3.